### v1

# Lennart Berg and Torben Wiedemann

Abstract
Using the well-known realisation of the root system $H_{3}$ as a folding of $D_{6}$, one can construct examples of $H_{3}$-graded groups from Chevalley groups of type $D_{6}$. Such Chevalley groups are defined over a commutative ring $\mathcal{R}$, and the root groups of the resulting $H_{3}$-grading are coordinatised by $\mathcal{R}\times\mathcal{R}$. We show that every $H_{3}$-graded group arises as the folding of a $D_{6}$-graded group, or in other words, that it is coordinatised by $\mathcal{R}\times\mathcal{R}$ for some commutative ring $\mathcal{R}$.
1 Introduction
In his seminal paper [Che55], Chevalley introduced the groups which are today known as Chevalley groups. By construction, any such group is equipped with a family of subgroups $(U_{\alpha})_{\alpha\in\Phi}$ indexed by a finite crystallographic reduced root system $\Phi$. These root groups are coordinatised by a commutative ring and the commutator of two root groups adheres to the celebrated Chevalley commutator formula.
Motivated by these observations, we say that a $\Phi$-grading of a group $G$ is a family of non-trivial subgroups $(U_{\alpha})_{\alpha\in\Phi}$ satisfying some (purely combinatorial) properties of the root groups in Chevalley groups. Here $\Phi$ is any finite root system which is not assumed to be crystallographic or reduced. Our definition of root graded groups was introduced in [Wie23] and encompasses several existing concepts in the literature as special cases. In particular, it generalises Tits’ notion of RGD-systems (see [Tit92, p. 258]) and Shi’s notion of groups graded by a root system (see [Shi93]). An overview of the historic background of root graded groups is given in the preface of [Wie23].
The main result of [Wie23] says that if $\Phi$ is irreducible, crystallographic and of rank at least $3$, then every $\Phi$-graded group is coordinatised by an algebraic structure (whose nature depends on $\Phi$). This algebraic structure is a division structure if and only if the corresponding $\Phi$-graded group is an RGD-system.
In this paper, we prove a similar result for the non-crystallographic root systems $H_{3}$ and $H_{4}$, thereby extending the coordinatisation results of [Wie23] to the class of all irreducible finite root systems of rank at least 3. Specifically, we show in 8.31 that any root graded group $G$ of type $H_{3}$ or $H_{4}$ is coordinatised by $\mathcal{R}\times\mathcal{R}$ where $\mathcal{R}$ is a commutative ring (which depends on $G$). This is in line with Tits’ classical result from [Tit77, 3.5] that there exist no thick buildings of types $H_{3}$ and $H_{4}$: The ring $\mathcal{R}\times\mathcal{R}$ cannot be a division ring, and hence there do not exist RGD-systems of types $H_{3}$ and $H_{4}$. Most of the time, we can restrict our attention to the root system $H_{3}$, the transfer to $H_{4}$ being immediate.
Since $H_{3}$ is not crystallographic, there are no Chevalley groups of type $H_{3}$, so it is not evident a priori that there actually exist $H_{3}$-graded groups. However, using the well-known realisation of the root system $H_{3}$ as a folding of $D_{6}$, examples of $H_{3}$-graded groups can be constructed from Chevalley groups of type $D_{6}$ (or, in fact, from any $D_{6}$-graded group). We will give a detailed description of this construction, which generalises the construction of the twisted Chevalley groups (see, for example, [Car72, Chapter 13]). In a similar way, one obtains examples of $H_{4}$-gradings as foldings of $E_{8}$-gradings. It is a consequence of our main result that, in fact, all root graded groups of types $H_{3}$ and $H_{4}$ arise in this way (see 8.40).
The proof of our main result relies on the two main tools from [Wie23]: Firstly, the parametrisation theorem, which yields canonical isomorphisms between the root groups of an arbitrary $H_{3}$-graded groups $G$ and hence provides us with a parametrisation of the root groups by a single abelian group $(\mathcal{S},+)$. Secondly, the blueprint technique, which is a computational tool that allows us to equip $(\mathcal{S},+)$ with the structure of a commutative ring decomposing as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}$ for isomorphic ideals $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. The blueprint technique is inspired by the work of Ronan-Tits on the construction of buildings in [RT87].
In [MW23, 9.16], it is shown that an analogue of our coordinatisation result for $H_{3}$-graded groups holds for $H_{2}$-graded groups which arise from a Tits pentagon that satisfies a certain thickness condition. By [MW19, Section 5], the $H_{2}$-graded groups arising from a Tits pentagon are precisely those which satisfy a certain stability condition. Our main result is true regardless of any stability assumptions, but in return, it uses in a crucial way that the root systems $H_{3}$ and $H_{4}$ are of rank higher than 2.
This paper is organised as follows. In Sections 2 and 3, we assemble some basic facts about root systems and introduce the notion of root graded groups. In Section 4, we construct examples of $H_{3}$-graded groups via foldings and explicitly compute the commutator formulas in these groups. The groundwork for the coordinatisation of $H_{3}$-graded groups is laid in Section 5, where we investigate the commutator relations and Weyl elements in such groups. The main result of this section is 5.20, which provides formulas for the action of squares of Weyl elements on all root groups. These formulas allow us to apply the parametrisation theorem, which is proven in [Wie23] for arbitrary root systems. In Section 6, we will formulate the parametrisation theorem for the root system $H_{3}$ and, for the reader’s convenience, sketch a simplified version of its proof in this special case. In Section 7, we will describe the general idea of the blueprint technique and then apply it to the special case of $H_{3}$-graded groups, which results in our coordinatisation theorem (Theorem 8.31). As a corollary, we obtain that every $H_{3}$-grading is the folding of a $D_{6}$-grading (Theorem 8.40).
Acknowledgements
We are grateful to Bernhard Mühlherr for numerous insightful discussions and for his encouragement. The work of the first author was supported by the Studienstiftung des deutschen Volkes and the work of the second author was supported by the DFG Grant MU1281/7-1.
2 Preliminaries
Convention 2.1.
Let $G$ be a group. For any subset $U$ of $G$ we denote by $\langle U\rangle$ the subgroup of $G$ generated by $U$. Furthermore, for all $g,h\in G$, we put $g^{h}\coloneqq h^{-1}gh$ and $[g,h]\coloneqq g^{-1}h^{-1}gh$.
Remark 2.2.
With the conventions in 2.1, we have the following relations for all $g,g_{1},g_{2},h,h_{1},h_{2}$ in any group $G$:
  1. (i)
    $g^{h}=[h,g^{-1}]g=g[g,h]$.
  2. (ii)
    $gh=hg^{h}$.
  3. (iii)
    $gh=hg[g,h]=[h^{-1},g^{-1}]hg$.
  4. (iv)
    $[g,h]^{-1}=[h,g]$.
  5. (v)
    $[g_{1}g_{2},h]=[g_{1},h]^{g_{2}}[g_{2},h]$.
  6. (vi)
    $[g,h_{1}h_{2}]=[g,h_{2}][g,h_{1}]^{h_{2}}$.
Convention 2.3.
We only consider unital rings in this paper, and homomorphism of rings preserve the identity. Further, all rings are assumed to be associative except when we specifically refer to “nonassociative rings”.
Throughout this paper, we will mostly use the standard terminology for finite root systems that are not necessarily crystallographic. However, though $H_{3}$ is a reduced root system, it will be practical to consider the non-reduced root system $GH_{3}$ (which we will introduce in 4.5) at some points. Hence we use the following definition of root systems.
Definition 2.4.
A root system is a finite subset $\Phi$ of a Euclidean space $(V,\cdot)$ such that $\Phi^{\sigma(\alpha)}=\Phi$ for all $\alpha\in\Phi$ where $\sigma(\alpha)\coloneqq\sigma_{\alpha}\colon V\rightarrow V,v\mapsto v-2\frac{\alpha\cdot v}{\alpha\cdot\alpha}\alpha$ denotes the reflection along $\alpha^{\perp}$. It is called reduced if $\Phi\cap\mathbb{R}\alpha=\{\alpha,-\alpha\}$ for all $\alpha\in\Phi$ and its Weyl group is $\operatorname{Weyl}(\Phi)\coloneqq\langle\sigma_{\alpha}\mid\alpha\in\Phi\rangle$.
Notation 2.5.
Let $\Phi$ be a (fixed) root system and let $A\subseteq\Phi$. Then $\Psi_{A}\coloneqq\langle A\rangle_{\mathbb{R}}\cap\Phi$ is called the root subsystem (of $\Phi$) spanned by $A$. In general, a root subsystem of $\Phi$ is the root subsystem spanned by some subset of $\Phi$. If $A$ has exactly two elements $\alpha$, $\beta$, we will write $\Psi_{\alpha,\beta}$ in place of $\Psi_{S}$. If $X_{n}$ is the type of a root system, we will often write $X_{n}$-subsystem” in place of “root subsystem of type $X_{n}$.
We briefly recall the notion of root intervals, which plays an important role in the theory of root gradings.
Definition 2.6.
Let $\Phi$ be a root system and let $A,B\subseteq\Phi$. We denote by $\mathord{]A,B[}$ (respectively, $\mathord{]A,B[}_{\text{cry}}$) the set of all roots $\gamma\in\Phi$ which can be written as $\gamma=\sum_{i=1}^{n}\lambda_{i}\alpha_{i}+\sum_{j=1}^{m}\mu_{j}\beta_{j}$ for $n,m\in\mathbb{Z}_{>0}$, $\alpha_{1},\ldots,\alpha_{n}\in A$, $\beta_{1},\ldots,\beta_{m}\in B$ and $\lambda_{1},\ldots,\lambda_{n},\mu_{1},\ldots,\mu_{m}$ positive real (respectively, integral) numbers. For non-proportional roots $\alpha,\beta\in\Phi$, the root interval between $\alpha$ and $\beta$ and the crystallographic root interval between $\alpha$ and $\beta$ are $\mathord{]\alpha,\beta[}\coloneqq\mathord{]\{\alpha\},\{\beta\}[}$ and $\mathord{]\alpha,\beta[}_{\text{cry}}\coloneqq\mathord{]\{\alpha\},\{\beta\}[}_{\text{cry}}$, and $\alpha,\beta$ are called (crystallographically) adjacent if $\mathord{]\alpha,\beta[}$ (respectively, $\mathord{]\alpha,\beta[}_{\text{cry}}$) is empty.
Lemma 2.7.
Let $\Phi$ be a reduced root system, let $\beta,\gamma\in\Phi$ be non-proportional and denote by $k$ the cardinality of $\mathord{]\beta,\gamma[}$. Then there exists a unique way to label the roots in $\mathord{]\beta,\gamma[}$ by $\alpha_{1},\ldots,\alpha_{k}$ such that $\mathord{]\alpha_{i},\alpha_{j}[}=\{\alpha_{p}\mid i<p<j\}$ for all $i<j\in\{0,\ldots,k+1\}$ where $\alpha_{0}\coloneqq\beta$ and $\alpha_{k+1}\coloneqq\gamma$. The tuple $(\alpha_{1},\ldots,\alpha_{k})$ is called the interval ordering of $\mathord{]\beta,\gamma[}$ starting from $\beta$.
We end this section with the definition of $H_{2}$, $H_{3}$ and $H_{4}$ and some of their basic properties.
Notation 2.8.
[…]
Lemma 2.9.
The Weyl group of $H_{3}$ acts transitively on $H_{3}$.
Remark 2.10.
In the literature, the root system $H_{2}$ is sometimes referred to as $I_{2}(5)$. Note that it is not crystallographic because $2\frac{\rho_{2}\cdot\rho_{3}}{\rho_{3}\cdot\rho_{3}}=2\cos(\frac{\pi}{5})=\frac{1+\sqrt{5}}{2}$ where $(\rho_{2},\rho_{3})$ denotes a root base of $H_{2}$. This number is known as the golden ratio, and we will denote it by $\tau$. Similarly, $H_{3}$ and $H_{4}$ are not crystallographic.
Remark 2.11.
Throughout this paper, we will make several assertions which follow from a straightforward but lengthy computation. The interested reader may find GAP [GAP21] code which verifies these assertions in [BW24].
Proposition 2.12.
Let $\alpha$ be any root in $H_{3}$. Then the following hold:
  1. (i)
    $\alpha$ is contained in exactly $2$ subsystems of type $H_{2}$.
  2. (ii)
    $\alpha$ is contained in exactly $2$ subsystems of type $A_{2}$.
  3. (iii)
    $\alpha$ is contained in exactly $2$ subsystems of type $A_{1}\times A_{1}$.
Proof.
It follows from a straightforward inspection of $H_{3}$ that the assertion holds for $\alpha=\rho_{1}$ (see 2.11 for details). Then the general assertion follows from 2.9. ∎
Definition 2.13.
Let $\Phi$, $\Phi^{\prime}$ be root systems such that $\Phi^{\prime}$ is of rank 2. A pair $(\alpha,\beta)\in\Phi^{2}$ is called a $\Phi^{\prime}$-pair (in $\Phi$) if the root subsystem $\Psi_{\alpha,\beta}$ of $\Phi$ is of type $\Phi^{\prime}$ and $(\alpha,\beta)$ is a root base of this subsystem. If $(\alpha,\varepsilon)$ is an $H_{2}$-pair in $\Phi$, then the $H_{2}$-quintuple associated to $(\alpha,\varepsilon)$ is the tuple $(\alpha,\beta,\gamma,\delta,\varepsilon)$ where $\beta\coloneqq\tau\alpha+\varepsilon$, $\gamma\coloneqq\tau\alpha+\tau\varepsilon$ and $\delta\coloneqq\alpha+\tau\varepsilon$. Note that $(\beta,\gamma,\delta)$ is the interval ordering of $\mathord{]\alpha,\epsilon[}$ starting from $\alpha$. Similarly, the $A_{2}$-triple associated to an $A_{2}$-pair $(\alpha,\gamma)$ is $(\alpha,\alpha+\gamma,\gamma)$.
3 Root graded groups
Notation 3.1.
Throughout this section, we denote by $\Phi$ an arbitrary root system with root base $\Delta$. From after 3.4 on, we denote by $(G,(U_{\alpha})_{\alpha\in\Phi})$ a $\Phi$-graded group.
In this section, we introduce root graded groups in the sense of [Wie23].
Notation 3.2.
If $G$ is a group, $(U_{\alpha})_{\alpha\in\Phi}$ is a family of subgroups and $A$ is a subset of $\Phi$, we put $U_{A}\coloneqq\langle U_{\alpha}\mid\alpha\in A\rangle$.
Definition 3.3.
Let $G$ be a group and let $(U_{\alpha})_{\alpha\in\Phi}$ be a family of non-trivial subgroups. We say that $G$ has $\Phi$-commutator relations with root groups $(U_{\alpha})_{\alpha\in\Phi}$ if it satisfies the following conditions:
  1. (i)
    $[U_{\alpha},U_{\beta}]\subseteq U_{\mathord{]\alpha,\beta[}}$ for all non-proportional $\alpha,\beta\in\Phi$.
  2. (ii)
    Let $\alpha\in\Phi$ and $\beta\in\mathbb{R}_{>0}\alpha\cap\Phi$. Then $[U_{\alpha},U_{\beta}]\subseteq U_{\alpha+\beta}$ if $\alpha+\beta$ is a root and $[U_{\alpha},U_{\beta}]=\{1_{G}\}$ otherwise. Further, $U_{\beta}\subseteq U_{\alpha}$ if $\beta\in\mathbb{Z}_{>0}\alpha\cap\Phi$.
If, moreover, we have $[U_{\alpha},U_{\beta}]\subseteq U_{\mathord{]\alpha,\beta[}_{\text{cry}}}$ for all non-proportional $\alpha,\beta\in\Phi$, we say that $G$ has crystallographic $\Phi$-commutator relations with root groups $(U_{\alpha})_{\alpha\in\Phi}$.
Definition 3.4.
Let $G$ be a group. A (crystallographic) $\Phi$-grading is a family $(U_{\alpha})_{\alpha\in\Phi}$ of non-trivial subgroups satisfying the following conditions.
  1. (i)
    The group $G$ is generated by $(U_{\alpha})_{\alpha\in\Phi}$.
  2. (ii)
    The group $G$ has (crystallographic) $\Phi$-commutator relations with root groups $(U_{\alpha})_{\alpha\in\Phi}$.
  3. (iii)
    For all $\alpha\in\Phi$, the set
    $M_{\alpha}\coloneqq\{w_{\alpha}\in U_{-\alpha}U_{\alpha}U_{-\alpha}\mid U_{\beta}^{w_{\alpha}}=U_{\beta^{\sigma(\alpha)}}\text{ for all }\beta\in\Phi\}$
    of so-called $\alpha$-Weyl elements is non-empty.
  4. (iv)
    $U_{\Pi}\cap U_{-\alpha}=\{1_{G}\}$ for all positive systems of roots $\Pi$ and all $\alpha\in\Pi$.
We will also refer to $(G,(U_{\alpha})_{\alpha\in\Phi})$ or $G$ as a (crystallographic) $\Phi$-graded group if these conditions are satisfied. Further, we define
$U_{\alpha}^{\sharp}\coloneqq\{u_{\alpha}\in U_{\alpha}\mid\text{There exist }u_{-\alpha},u_{-\alpha}^{\prime}\in U_{-\alpha}\text{ such that }u_{-\alpha}u_{\alpha}u_{-\alpha}^{\prime}\in M_{\alpha}\}$
for all $\alpha\in\Phi$.
For the rest of this section, we denote by $(G,(U_{\alpha})_{\alpha\in\Phi})$ an arbitrary $\Phi$-graded group.
Remark 3.5.
Axiom 3.3 (ii) differs from the axiom (RGG-Div) used in [Wie23, 2.1.4, 2.5.2]. The reason for this is, as explained in [Wie23, 2.1.9], that (RGG-Div) is specifically tailored for root systems which are either reduced or crystallographic. The root system $GH_{3}$, which we will encounter in 4.5, satisfies neither assumption, so an adjustment of this axiom is necessary. If $\Phi$ is reduced, then a $\Phi$-grading in the sense of [Wie23, 2.5.2] is a $\Phi$-grading in the sense of 3.4 if and only if all root groups are abelian. If $\Phi=H_{3}$, we will see in 5.3 (i) that the latter assumption is automatically satisfied without requiring Axiom 3.3 (ii), so the two notions of $H_{3}$-graded groups coincide.
Remark 3.6.
Since the Weyl group acts transitively on the set of positive systems in $\Phi$, the existence of Weyl elements implies that it suffices to require that Axiom 3.4 (iv) holds for some positive system $\Pi$.
We now collect some basic properties of root graded groups.
Proposition 3.7 ([Wie23, 2.5.11]).
Let $\Psi$ be a root subsystem of $\Phi$ and put $G_{\Psi}\coloneqq\langle U_{\gamma}\mid\gamma\in\Psi\rangle$. Then $(U_{\gamma})_{\gamma\in\Psi}$ is a $\Psi$-grading of $G_{\Psi}$.
Lemma 3.8 ([Wie23, 2.1.13]).
Let $\alpha,\gamma\in\Phi$ such that the root interval $\mathord{]\alpha,\gamma[}$ contains exactly one element and let $x_{\alpha},x_{\alpha}^{\prime}\in U_{\alpha}$, $y_{\gamma},y_{\gamma}^{\prime}\in U_{\gamma}$. Then
$[x_{\alpha}x_{\alpha}^{\prime},y_{\gamma}]=[x_{\alpha},y_{\gamma}][x_{\alpha}^{\prime},y_{\gamma}]\hskip 11.38092pt\text{and}\hskip 11.38092pt[x_{\alpha},y_{\gamma}y_{\gamma}^{\prime}]=[x_{\alpha},y_{\gamma}][x_{\alpha},y_{\gamma}^{\prime}].$
In particular, $[x_{\alpha}^{-1},y_{\gamma}]=[x_{\alpha},y_{\gamma}]^{-1}=[x_{\alpha},y_{\gamma}^{-1}]$.
Proof.
Denote by $\beta$ the unique root in $\mathord{]\alpha,\gamma[}$. By 2.2,
$[x_{\alpha}x_{\alpha}^{\prime},y_{\gamma}]=[x_{\alpha},y_{\gamma}]^{x_{\alpha}^{\prime}}[x_{\alpha}^{\prime},y_{\gamma}]\quad\text{and}\quad[x_{\alpha},y_{\gamma}y_{\gamma}^{\prime}]=[x_{\alpha},y_{\gamma}][x_{\alpha},y_{\gamma}^{\prime}]^{y_{\gamma}^{\prime}}.$
Since $U_{\beta}$ commutes with $U_{\alpha}$ and $U_{\gamma}$, the assertion follows. ∎
Proposition 3.9 ([Wie23, 2.2.6]).
Let $\alpha\in\Phi$. Then the following hold:
  1. (i)
    $M_{\alpha}=M_{\alpha}^{-1}\coloneqq\{w_{\alpha}^{-1}\mid w_{\alpha}\in M_{\alpha}\}$ and $U_{\alpha}^{\sharp}=(U_{\alpha}^{\sharp})^{-1}$.
  2. (ii)
    $M_{\alpha}^{w_{\beta}}=M_{\beta^{\sigma(\alpha)}}$ and $(U_{\alpha}^{\sharp})^{w_{\beta}}=U_{\beta^{\sigma(\alpha)}}^{\sharp}$ for all $\beta\in\Phi$ and $w_{\beta}\in M_{\beta}$.
  3. (iii)
    Let $w_{\alpha}$ be an $\alpha$-Weyl element with factorization $w_{\alpha}=a_{-\alpha}b_{\alpha}c_{-\alpha}$ for some $b_{\alpha}\in U_{\alpha}$, $a_{-\alpha},c_{-\alpha}\in U_{-\alpha}$. Then $w_{\alpha}=c_{-\alpha}^{w_{\alpha}^{-1}}a_{-\alpha}b_{\alpha}$ and $w_{\alpha}=b_{\alpha}c_{-\alpha}a_{-\alpha}^{w_{\alpha}}$. In particular, $w_{\alpha}$ is also a ($-\alpha$)-Weyl element and $M_{\alpha}=M_{-\alpha}$.
  4. (iv)
    Let $b_{\alpha}\in U_{\alpha}^{\sharp}$ and choose $a_{-\alpha},c_{-\alpha}\in U_{-\alpha}$ such that $a_{-\alpha}b_{\alpha}c_{-\alpha}$ is an $\alpha$-Weyl element. Then $a_{-\alpha},c_{-\alpha}\in U_{-\alpha}^{\sharp}$.
The following assertion is a consequence of Axiom 3.4 (iv).
Proposition 3.10 ([Wie23, 2.5.9, 2.1.19]).
Assume that $\Phi$ is reduced, let $\beta,\gamma\in\Phi$ be non-proportional and denote by $(\alpha_{1},\ldots,\alpha_{k})$ the interval ordering of $\mathord{]\beta,\gamma[}$ starting from $\beta$ (as in 2.7). Then the product map
$U_{\alpha_{1}}\times\cdots\times U_{\alpha_{k}}\rightarrow U_{\mathord{]\beta,\gamma[}},(x_{1},\ldots,x_{k})\mapsto x_{1}\cdots x_{k}$
is bijective.
Using 3.10, we can introduce the following notation, which will be used frequently.
Notation 3.11.
Assume that $\Phi$ is reduced, let $\beta,\gamma\in\Phi$ be non-proportional and denote by $(\alpha_{1},\ldots,\alpha_{k})$ the interval ordering of $\mathord{]\beta,\gamma[}$ starting from $\beta$ (as in 2.7). Let $x_{\beta}\in U_{\beta}$ and $x_{\gamma}\in U_{\gamma}$. By 3.10, there exist unique elements $x_{i}\in U_{\alpha_{i}}$ for all $i\in\{1,\ldots,k\}$ such that $[x_{\beta},x_{\gamma}]=x_{1}\cdots x_{k}$. We put $[x_{\beta},x_{\gamma}]_{\alpha_{i}}\coloneqq x_{i}$ for all $i\in\{1,\ldots,n\}$. When we write $[x_{\beta},x_{\gamma}]_{\alpha_{i}}^{g}$ for some $g\in G$ or $[x_{\beta},x_{\gamma}]_{\alpha_{i}}^{-1}$, we always mean $([x_{\beta},x_{\gamma}]_{\alpha_{i}})^{g}$ or $([x_{\beta},x_{\gamma}]_{\alpha_{i}})^{-1}$, respectively.
Remark 3.12.
In general, $[x_{\beta},x_{\gamma}]_{\alpha_{i}}$ depends on the (ordered) tuple $(x_{\beta},x_{\gamma})$ and not only on the element $[x_{\beta},x_{\gamma}]$ because the interval ordering $(\alpha_{1},\ldots,\alpha_{k})$ depends on the starting root $\beta$.
Lemma 3.13 ([Wie23, 2.1.18, 2.2.9]).
Assume that $\Phi$ is reduced, let $\alpha,\gamma\in\Phi$ be non-proportional, let $\beta\in\mathord{]\alpha,\gamma[}$ and let $x_{\alpha}\in U_{\alpha}$, $x_{\gamma}\in U_{\gamma}$. Then for any root $\delta$ and any $\delta$-Weyl element $w_{\delta}$, we have
$[x_{\alpha},x_{\gamma}]_{\beta}^{w_{\delta}}=[x_{\alpha}^{w_{\delta}},x_{\gamma}^{w_{\delta}}]_{\beta^{\sigma(\delta)}}\quad\text{and}\quad[x_{\alpha},x_{\gamma}]_{\beta}^{w_{\delta}^{2}}=[x_{\alpha}^{w_{\delta}^{2}},x_{\gamma}^{w_{\delta}^{2}}]_{\beta}.$
Further, $[x_{\alpha},x_{\gamma}]_{\beta}^{-1}=[x_{\gamma},x_{\alpha}]_{\beta}$.
Remark 3.14.
The assertion that $[x_{\alpha},x_{\gamma}]_{\beta}^{-1}=[x_{\gamma},x_{\alpha}]_{\beta}$ is stated in [Wie23, 2.1.18] only for the case that $\mathord{]\alpha,\gamma[}$ has at most two elements, but the same proof remains valid in general. This fact is related to 3.12.
The well-known braid relations in Coxeter groups also hold for Weyl elements in root graded groups. This result goes back to [TW02, (6.9)], where it is stated and proven in a less general context. However, essentially the same proof remains valid in our setting, as is verified in [Wie23, 2.2.34].
Proposition 3.15 ([TW02, (6.9)]).
Assume that $\Phi$ is reduced, let $\alpha,\beta\in\Delta$ and denote by $m$ the order of $\sigma_{\alpha}\sigma_{\beta}$ in the Weyl group. Then for all Weyl elements $w_{\alpha}\in M_{\alpha}$, $w_{\beta}\in M_{\beta}$, we have
$w_{\alpha}w_{\beta}w_{\alpha}\cdots=w_{\beta}w_{\alpha}w_{\beta}\cdots$
where the products on both sides have length $m$. In other words, $(w_{\alpha}w_{\beta})^{m/2}=(w_{\beta}w_{\alpha})^{m/2}$ if $m$ is even and $(w_{\alpha}w_{\beta})^{(m-1)/2}w_{\alpha}=(w_{\beta}w_{\alpha})^{(m-1)/2}w_{\beta}$ if $m$ is odd.
4 Foldings
Notation 4.1.
Let $n\geq 4$. In this section, we use the standard representations of the root systems $A_{n-1}$ and $D_{n}$ from [Hum78, 2.10]:
$A_{n-1}$ =$=\{e_{i}-e_{j}\mid i\neq j\in\{1,\ldots,n\}\},\quad D_{n}=\{\pm e_{i}\pm e_{j}\mid i<j\in\{1,\ldots,n\}\}.$
Here $(e_{1},\ldots,e_{n})$ denotes the standard basis of $\mathbb{R}^{n}$ (with the standard inner product). Further, we fix a commutative ring $\mathcal{R}$.
Given a root system $\Phi$ and an automorphism $\rho$ of its Dynkin diagram, one can construct a new root system $\Phi^{\prime}$ by projecting $\Phi$ orthogonally to the fixed space of an isometry $\tau$ which is related to $\rho$ (with $\tau=\rho$ if $\Phi$ is reduced, which is the only case we will be interested in). In particular, we have a map $\Phi\rightarrow\Phi^{\prime}$ which we call a folding, and we will also refer to $\Phi^{\prime}$ as a folding of $\Phi$. Further, any Chevalley group of type $\Phi$ gives rise to a so-called twisted Chevalley group of type $\Phi^{\prime}$. This procedure is described in [Car72, Chapter 13]. In a similar way, every $\Phi$-grading $(U_{\alpha})_{\alpha\in\Phi}$ of a group $G$ gives rise to a $\Phi^{\prime}$-grading $(U_{\alpha^{\prime}}^{\prime})_{\alpha^{\prime}\in\Phi^{\prime}}$ of (the same group) $G$.
It is well-known that there exists a more general class of foldings which do not arise from diagram automorphisms. See [Mü93] for a description of this phenomenon. The goal of this section is to construct an $H_{3}$-graded group $G$ from a Chevalley group of type $D_{6}$ by way of the folding $D_{6}\rightarrow H_{3}$. Further, we will compute the commutator relations in $G$ (Figure 4). The main result of this paper is that every $H_{3}$-graded group has a coordinatisation by $\mathcal{R}^{\prime}\times\mathcal{R}^{\prime}$ for a commutative ring $\mathcal{R}^{\prime}$ such that the same commutator relations as in $G$ are satisfied (Theorem 8.31) and that every $H_{3}$-graded group arises from a $D_{6}$-graded group by the folding procedure described in this section (Theorem 8.40).
To illustrate the general folding procedure, we begin with the easier example of the folding $A_{4}\rightarrow H_{2}$.
Notation 4.2.
In order to emphasise that we consider specific examples of root graded groups (and not arbitrary root graded groups) in this section, we place a tilde on the root groups. Further, for the purpose of distinguishing different gradings of the same surrounding group $G$, we will use superscripts to indicate the root system. For example, we will write $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$ for an $A_{4}$-grading.
Example 4.3.
For $1\leq i\neq j\leq 5$ denote by $e_{ij}$ the $(5\times 5)$-matrix with $1_{\mathcal{R}}$ at position $(i,j)$ and $0_{\mathcal{R}}$ at every other position. For all roots $\alpha=e_{i}-e_{j}\in A_{4}$, we define a root group
$\tilde{U}_{\alpha}^{A}\coloneqq\tilde{U}_{ij}^{A}\coloneqq\{\mathord{\mathord{\mathrm{id}}}+ae_{ij}\mid a\in\mathcal{R}\}\leq\operatorname{SL}_{5}(\mathcal{R}).$
We denote by $\operatorname{E}_{5}(\mathcal{R})$ the group generated by these root groups, which is a Chevalley group of type $A_{4}$. Then $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$ is an $A_{4}$-grading of $\operatorname{E}_{5}(\mathcal{R})$.
Denote the ten roots in $H_{2}$ by $\chi_{1},\ldots,\chi_{10}$ where $(\chi_{1},\chi_{2},\chi_{3},\chi_{4},\chi_{5})$ is an (arbitrarily chosen) $H_{2}$-quintuple and $\chi_{i+5}\coloneqq-\chi_{i}$ for all $i\in\{1,\ldots,5\}$. We define root groups $\tilde{U}_{\chi_{i}}^{H}\coloneqq\tilde{U}_{i}^{H}$ for all $i\in\{1,\ldots,10\}$ by
$\tilde{U}_{1}^{H}\coloneqq\tilde{U}_{12}^{A}\tilde{U}_{34}^{A},\>\tilde{U}_{2}^{H}\coloneqq\tilde{U}_{35}^{A}\tilde{U}_{14}^{A},\>\tilde{U}_{3}^{H}\coloneqq\tilde{U}_{24}^{A}\tilde{U}_{15}^{A},\>\tilde{U}_{4}^{H}\coloneqq\tilde{U}_{13}^{A}\tilde{U}_{25}^{A},\>\tilde{U}_{5}^{H}\coloneqq\tilde{U}_{45}^{A}\tilde{U}_{23}^{A},$
$\tilde{U}_{6}^{H}\coloneqq\tilde{U}_{21}^{A}\tilde{U}_{43}^{A},\>\tilde{U}_{7}^{H}\coloneqq\tilde{U}_{53}^{A}\tilde{U}_{41}^{A},\>\tilde{U}_{8}^{H}\coloneqq\tilde{U}_{42}^{A}\tilde{U}_{51}^{A},\>\tilde{U}_{9}^{H}\coloneqq\tilde{U}_{31}^{A}\tilde{U}_{52}^{A},\>\tilde{U}_{10}^{H}\coloneqq\tilde{U}_{54}^{A}\tilde{U}_{32}^{A}.$
By a straightforward computation, which we leave out because this example serves only motivational purposes, $(\tilde{U}_{\chi}^{H})_{\chi\in H_{2}}$ is an $H_{2}$-grading of $\operatorname{E}_{5}(\mathcal{R})$. We call $(\tilde{U}_{\chi}^{H})_{\chi\in H_{2}}$ a folding of $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$. Note that the surrounding group $\operatorname{E}_{5}(\mathcal{R})$ is the same for both gradings: only the family of root groups changes.
Observe that the root group of the simple root $\chi_{1}$ is obtained by “merging” the root groups of the simple roots $\delta_{1}\coloneqq e_{1}-e_{2}$ and $\delta_{3}\coloneqq e_{3}-e_{4}$ while $\tilde{U}_{\chi_{5}}^{H}$ is obtained from the simple roots $\delta_{2}\coloneqq e_{2}-e_{3}$ and $\delta_{4}\coloneqq e_{4}-e_{5}$. Hence our folding corresponds to the partition $\{\{\delta_{1},\delta_{3}\},\{\delta_{2},\delta_{4}\}\}$ of the root base $\{\delta_{1},\ldots,\delta_{4}\}$ of $A_{4}$, which we encode by the diagram in Figure 1(a). Note further that for each $i\in\{1,\ldots,10\}$, the two root groups in the definition of $\tilde{U}_{i}^{H}$ commute, so that $\tilde{U}_{i}^{H}$ is isomorphic to $\mathcal{R}\times\mathcal{R}$.
$\delta_{1}$$\delta_{2}$$\delta_{3}$$\delta_{4}$
(a) The folding $A_{4}\rightarrow H_{2}$.
$\delta_{1}$$\delta_{2}$$\delta_{3}$$\delta_{6}$$\delta_{4}$$\delta_{5}$
(b) The folding $D_{6}\rightarrow H_{3}$.
$\delta_{1}$$\delta_{2}$$\delta_{3}$$\delta_{4}$$\delta_{5}$$\delta_{6}$$\delta_{7}$$\delta_{8}$
(c) The folding $E_{8}\rightarrow H_{4}$.
Figure 1: Three foldings.
Remark 4.4.
On the level of Coxeter groups, $(\langle s_{1},s_{2}\rangle,\{s_{1},s_{2}\})$ is a Coxeter system of type $H_{2}$ where $s_{1}\coloneqq\sigma_{e_{1}-e_{2}}\sigma_{e_{3}-e_{4}}$ and $s_{2}\coloneqq\sigma_{e_{2}-e_{3}}\sigma_{e_{4}-e_{5}}$. This realisation of $\operatorname{Weyl}(H_{2})$ as a subgroup of $\operatorname{Weyl}(A_{4})$ is a special case of the construction in [Mü93]: Given a Coxeter system $(W,S)$ and a partition $I=\{I_{1},\ldots,I_{k}\}$ of $S$ which has the so-called property of being admissible, the pair $(\langle R\rangle,R)$ is a Coxeter system where $R=\{r_{1},\ldots,r_{k}\}$ and $r_{j}$ is the longest element in the Coxeter system $(\langle I_{j}\rangle,I_{j})$ for all $j\in\{1,\ldots,k\}$. In all cases we consider, each set $I_{j}$ has cardinality $2$ and its elements $s_{j}$, $t_{j}$ commute (that is, $(\langle I_{j}\rangle,I_{j})$ is of type $A_{1}\times A_{1}$), so that $r_{j}=s_{j}t_{j}$.
We now turn to the realisation of $H_{3}$ as a folding of $D_{6}$. We will find the non-reduced root system $GH_{3}$, the “golden version of $H_{3}$, as an intermediate step between $D_{6}$ and $H_{3}$.
Construction 4.5.
We consider the ordered root base
$\Delta_{D}=(\delta_{1},\ldots,\delta_{6})=(e_{1}-e_{2},e_{2}-e_{3},e_{3}-e_{4},e_{4}-e_{5},e_{5}-e_{6},e_{5}+e_{6})$
of $D_{6}$ and the partition $\{\{\delta_{1},\delta_{6}\},\{\delta_{2},\delta_{4}\},\{\delta_{3},\delta_{5}\}\}$ of $\Delta_{D}$ given by the diagram in Figure 1(b). By [Mü93, 1.1, 5.4], $(\langle s_{1},s_{2},s_{3}\rangle,R)$ is a Coxeter system of type $H_{3}$ where $s_{1}\coloneqq\sigma_{\delta_{1}}\sigma_{\delta_{6}}$, $s_{2}\coloneqq\sigma_{\delta_{2}}\sigma_{\delta_{4}}$, $s_{3}\coloneqq\sigma_{\delta_{3}}\sigma_{\delta_{5}}$ and $R\coloneqq\{s_{1},s_{2},s_{3}\}$.
We briefly sketch how to realise the embedding of $\operatorname{Weyl}(H_{3})$ into $\operatorname{Weyl}(D_{6})$ on the level of root systems. Put $W\coloneqq\langle\delta_{1},\ldots,\delta_{6}\rangle_{\mathbb{R}}$ and define subspaces
$W_{1}$ $\coloneqq\langle\delta_{3}-\tau\delta_{5},\delta_{4}-\tau\delta_{2},\delta_{1}-\tau\delta_{6}\rangle_{\mathbb{R}},\quad W_{2}\coloneqq\langle\delta_{5}+\tau\delta_{3},\delta_{2}+\tau\delta_{4},\delta_{6}+\tau\delta_{1}\rangle_{\mathbb{R}}$
where $\tau\coloneqq\frac{1+\sqrt{5}}{2}$ denotes the golden ratio. A straightforward computation (see 2.11) shows that $W=W_{1}\oplus W_{2}$ is an orthogonal, $R$-invariant decomposition of $W$ and that $\pi(D_{6})$ is a non-reduced root system in $W_{2}$ with $60$ roots where $\pi:W\rightarrow W_{2}$ denotes the orthogonal projection onto $W_{2}$. More precisely, we have $\pi(D_{6})=X\cup\tau X$ where $X$ denotes the set of indivisible roots in $\pi(D_{6})$. The only root systems with 30 roots are $A_{5}$ and $H_{3}$, and an inspection shows that $X$ is of type $H_{3}$. Hence $\pi(D_{6})$ is a root system of type $GH_{3}\coloneqq H_{3}\cup\tau H_{3}$, which we call the golden version of $H_{3}$. Note that the scaling map $\phi\colon v\mapsto\frac{1}{\lVert v\rVert}v$ maps $\pi(D_{6})$ onto $H_{3}$ and that $\pi\colon D_{6}\rightarrow\pi(D_{6})$ is a bijection.
Notation 4.6.
For the rest of this section, we identify $H_{3}$ with the image of $\pi(D_{6})$ under the scaling map $\phi$ in 4.5. Then $\Delta_{H}\coloneqq(\rho_{1},\rho_{2},\rho_{3})$ is a root base of $H_{3}$ in standard order where $\rho_{1}\coloneqq\phi\pi(e_{1}-e_{2})=\phi\pi(e_{5}+e_{6})$, $\rho_{2}\coloneqq\phi\pi(e_{2}-e_{3})=\phi\pi(e_{4}-e_{5})$ and $\rho_{3}\coloneqq\phi\pi(e_{3}-e_{4})=\phi\pi(e_{5}-e_{6})$. We will also denote roots in $H_{3}$ by $[a_{1},a_{2},a_{3}]\coloneqq\sum_{i=1}^{3}a_{i}\rho_{i}$ for $a_{1},a_{2},a_{3}\in\mathbb{R}$. For all $\alpha\in H_{3}$, we denote by $\kappa^{\text{sh}}(\alpha)$ and $\kappa^{\text{lo}}(\alpha)$ the two roots in $(\phi\circ\pi)^{-1}(\alpha)\cap D_{6}$, ordered so that $\pi(\kappa^{\text{lo}}(\alpha))=\tau\cdot\pi(\kappa^{\text{sh}}(\alpha))$ (that is, $\pi(\kappa^{\text{lo}}(\alpha))$ is long in $GH_{3}$ and $\pi(\kappa^{\text{sh}}(\alpha))$ is short). Further, we denote by $\Pi_{D}$, $\Pi_{H}$ and $\Pi_{GH}$ the positive systems in $D_{6}$, $H_{3}$ and $GH_{3}$ corresponding to $\Delta_{D}$, $\Delta_{H}$ and $\pi(\Delta_{D})$, respectively.
$\alpha$ $\kappa^{\text{sh}}(\alpha)$ $\kappa^{\text{lo}}(\alpha)$
$[1,0,0]$ $e_{1}-e_{2}$ $e_{5}+e_{6}$
$[0,1,0]$ $e_{2}-e_{3}$ $e_{4}-e_{5}$
$[0,0,1]$ $e_{5}-e_{6}$ $e_{3}-e_{4}$
$[1,1,0]$ $e_{1}-e_{3}$ $e_{4}+e_{6}$
$[1,1,\tau]$ $e_{1}-e_{4}$ $e_{3}+e_{5}$
$[1,\tau^{2},\tau]$ $e_{1}-e_{5}$ $e_{2}+e_{4}$
$[1,\tau^{2},\tau^{2}]$ $e_{1}-e_{6}$ $e_{2}+e_{3}$
$[0,1,\tau]$ $e_{2}-e_{4}$ $e_{3}-e_{6}$
$\alpha$ $\kappa^{\text{sh}}(\alpha)$ $\kappa^{\text{lo}}(\alpha)$
$[\tau,\tau^{2},\tau]$ $e_{2}+e_{6}$ $e_{1}+e_{4}$
$[\tau,2\tau,\tau^{2}]$ $e_{3}+e_{4}$ $e_{1}+e_{2}$
$[0,\tau,\tau]$ $e_{3}-e_{5}$ $e_{2}-e_{6}$
$[\tau,\tau,\tau]$ $e_{3}+e_{6}$ $e_{1}+e_{5}$
$[\tau,\tau,1]$ $e_{4}+e_{5}$ $e_{1}+e_{6}$
$[0,\tau,1]$ $e_{4}-e_{6}$ $e_{2}-e_{5}$
$[\tau,\tau^{2},\tau^{2}]$ $e_{2}+e_{5}$ $e_{1}+e_{3}$
Figure 2: The preimages of the positive system in $H_{3}$ under $\phi\circ\pi$ in $D_{6}$. See 4.5 and 4.6.
Observation 4.7.
For all $\alpha\in H_{3}$, the root $\kappa^{\text{lo}}(\alpha)$ is adjacent to $\kappa^{\text{sh}}(\alpha)$ and $-\kappa^{\text{sh}}(\alpha)$ (in $D_{6}$). This follows from an inspection of Figure 2.
Observation 4.8.
We have $\Pi_{D}=\{\kappa^{\text{lo}}(\alpha),\kappa^{\text{sh}}(\alpha)\mid\alpha\in\Pi_{H}\}$, which follows from an inspection of Figure 2. Hence $\Pi_{H}=\phi\pi(\Pi_{D})$.
Remark 4.9.
In a similar way as in 4.5, the orthogonal projection of $A_{4}$ on a suitable subspace yields the non-reduced root system $GH_{2}=H_{2}\cup\tau H_{2}$. See Figure 3 for a graphic depiction of $GH_{2}$ and [MW23] for more details on $GH_{2}$, which inspired our construction of $GH_{3}$.
$\tau\alpha$$\alpha$$\tau\beta$$\beta$$\tau\gamma$$\gamma$$\tau\delta$$\delta$$\tau\varepsilon$$\varepsilon$
Figure 3: The root system $GH_{2}$ with an $H_{2}$-quintuple $(\alpha,\beta,\gamma,\delta,\varepsilon)$.
Now that we have defined maps $D_{6}\rightarrow GH_{3}\rightarrow H_{3}$, we can give an example of an $H_{3}$-graded group by imitating the construction in 4.3. Along the we, we will also obtain a group with crystallographic $GH_{3}$-commutator relations.
Construction 4.10.
Denote by $e_{ij}\in\mathcal{R}^{6\times 6}$ the matrix with $1_{\mathcal{R}}$ at position $(i,j)$ and $0_{\mathcal{R}}$ at every other position. For each root $\alpha\in D_{6}$, we define a root homomorphism $\tilde{\theta}_{\alpha}^{D}\colon(\mathcal{R},+)\rightarrow\operatorname{GL}_{12}(\mathcal{R})$ by the formulas
$\tilde{\theta}_{e_{i}-e_{j}}^{D}(\lambda)\coloneqq\mathord{\mathrm{id}}+\lambda\begin{pmatrix}e_{ij}&0\\ 0&-e_{ji}\end{pmatrix}$
for all $i\neq j\in\{1,\ldots,6\}$ and
$\tilde{\theta}_{e_{i}+e_{j}}^{D}(\lambda)\coloneqq\mathord{\mathrm{id}}+\lambda\begin{pmatrix}0&e_{ij}-e_{ji}\\ 0&0\end{pmatrix},\quad\tilde{\theta}_{-e_{i}-e_{j}}^{D}(\lambda)\coloneqq\mathord{\mathrm{id}}+\lambda\begin{pmatrix}0&0\\ e_{ji}-e_{ij}&0\end{pmatrix}$
for all $i<j\in\{1,\ldots,6\}$, where $\lambda\in\mathcal{R}$. For each $\beta\in D_{6}$, we define a root group $\tilde{U}_{\beta}^{D}\coloneqq\tilde{\theta}_{\beta}^{D}(\mathcal{R})$ and we denote by $\operatorname{Ch}(D_{6},\mathcal{R})$ the group generated by all root groups. This is a Chevalley group of type $D_{6}$. Since $GH_{3}=\pi(D_{6})$ up to scaling and since the map $\pi\colon D_{6}\rightarrow\pi(D_{6})$ is a bijection (see 4.5), we may also define $\tilde{U}_{\beta}^{GH}\coloneqq\tilde{U}_{\beta^{\prime}}^{D}$ for all $\beta\in GH_{3}$ where $\beta^{\prime}$ is the unique root in $D_{6}$ with $\pi(\beta^{\prime})=\beta$. In other words, $\tilde{U}_{\alpha}^{D}=\tilde{U}_{\pi(\alpha)}^{GH}$ for all $\alpha\in D_{6}$.
Now let $\alpha\in H_{3}$ be arbitrary. We define a corresponding root group by
$\tilde{U}_{\alpha}^{H}\coloneqq\langle\tilde{U}_{\beta}^{D}\mid\beta\in D_{6}\cap(\phi\circ\pi)^{-1}(\alpha)\rangle=\langle\tilde{U}_{\beta}^{GH}\mid\beta\in GH_{3}\cap\phi^{-1}(\alpha)\rangle.$
Using the notation of 4.6, we have $D_{6}\cap(\phi\circ\pi)^{-1}(\alpha)=\{\beta_{1},\beta_{2}\}$ where $\beta_{1}\coloneqq\kappa^{\text{sh}}(\alpha)$ and $\beta_{2}\coloneqq\kappa^{\text{lo}}(\alpha)$. By 4.7, the groups $\tilde{U}_{\beta_{1}}^{D}$ and $\tilde{U}_{\beta_{2}}^{D}$ commute, so $\tilde{U}_{\alpha}^{H}=\tilde{U}_{\beta_{1}}^{D}\tilde{U}_{\beta_{2}}^{D}=\tilde{U}_{\beta_{2}}^{D}\tilde{U}_{\beta_{1}}^{D}$. Hence the root homomorphism
$\tilde{\theta}_{\alpha}^{H}\colon\mathcal{R}\times\mathcal{R}\rightarrow\tilde{U}_{\alpha}^{H},(x,y)\mapsto\begin{cases}\tilde{\theta}_{\beta_{1}}^{D}(x)\tilde{\theta}_{\beta_{2}}^{D}(y)&\text{if }\alpha\notin\{\rho_{1},-\rho_{1}\},\\ \tilde{\theta}_{\beta_{1}}^{D}(x)\tilde{\theta}_{\beta_{2}}^{D}(-y)&\text{if }\alpha\in\{\rho_{1},-\rho_{1}\}\end{cases}$
is bijective. The reason to define $\tilde{\theta}_{\alpha_{13}}^{H}$ in a “twisted” way is of aesthetic nature: It ensures that the $A_{2}$-pair $(\rho_{1},\rho_{2})=(\alpha_{13},\alpha_{7})$ of simple roots satisfies the commutator relation
$\left[\tilde{\theta}_{\rho_{1}}^{H}(a,b),\tilde{\theta}_{\rho_{2}}^{H}(c,d)\right]=\tilde{\theta}_{\rho_{1}+\rho_{2}}^{H}(ac,bd)$
for $(a,b),(c,d)\in\mathcal{R}\times\mathcal{R}$ (without an additional sign in this relation).
Remark 4.11.
The construction of the root groups $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ in 4.10 applies to any $D_{6}$-graded group $(G,(\tilde{U}_{\alpha}^{D})_{\alpha\in D_{6}})$ in place of $\operatorname{Ch}(D_{6},\mathcal{R})$. By [Wie23, 5.7.14], any such group $G$ has a parametrisation $(\tilde{\theta}_{\alpha}^{D}\colon(\mathcal{R}^{\prime},+)\rightarrow\tilde{U}_{\alpha}^{D})_{\alpha\in D_{6}}$ by a commutative ring $\mathcal{R}^{\prime}$, a result which was first proven in [Shi93]. This allows us to define root isomorphisms $(\tilde{\theta}_{\alpha}^{H}\colon(\mathcal{R}^{\prime}\times\mathcal{R}^{\prime},+)\rightarrow\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ as in 4.10. The main result 4.18 of this section and its proof remain true in this generality except that, in order for $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ to satisfy Axiom 3.4 (iv), we have to require that $\tilde{U}_{-\alpha}^{D}\tilde{U}_{-\alpha^{\prime}}^{D}\cap\tilde{U}_{\Pi_{D}}^{D}=\{1_{G}\}$ for all $\alpha\in\Pi_{D}$ where $\alpha^{\prime}$ is the unique root in $\Pi_{D}\setminus\{\alpha\}$ with $\phi\pi(\alpha^{\prime})=\phi\pi(\alpha)$. This requirement is only slightly stronger than Axiom 3.4 (iv) for $(\tilde{U}_{\alpha}^{D})_{\alpha\in D_{6}}$, and always satisfied if $\tilde{U}_{\Pi_{D}}^{D}\cap\tilde{U}_{-\Pi_{D}}^{D}=\{1\}$.
Notation 4.12.
For the rest of this section, we fix the notation from 4.10.
Before we can prove that $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ is an $H_{3}$-grading of $\operatorname{Ch}(D_{6},R)$, we have to verify some properties of the chain of maps $D_{6}\rightarrow GH_{3}\rightarrow H_{3}$ and state (the mild generalisation of) a general result from [Wie23].
Lemma 4.13.
Let $(i,j)\in\{(1,6),(3,5),(2,4)\}$, so that $r\coloneqq\sigma_{\delta_{i}}\sigma_{\delta_{j}}$ is an element of $R$ in 4.5. Then $\pi(\alpha^{r})=\pi(\alpha)^{r}=\pi(\alpha)^{\sigma(\pi(\delta_{i}))}=\pi(\alpha)^{\sigma(\pi(\delta_{j}))}$ for all $\alpha\in D_{6}$. Further, $\kappa^{\text{sh}}(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}=\kappa^{\text{sh}}(\gamma^{\sigma(\pi(\delta_{i}))})$ and $\kappa^{\text{lo}}(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}=\kappa^{\text{lo}}(\gamma^{\sigma(\pi(\delta_{i}))})$ for all $\gamma\in H_{3}$.
Proof.
Let $\alpha\in D_{6}$. Since $\pi$ is an orthogonal projection, we have $\pi(\alpha^{r})=\pi(\alpha)^{r}$. The assertion that $\pi(\alpha)^{r}=\pi(\alpha)^{\sigma(\pi(\delta_{i}))}=\pi(\alpha)^{\sigma(\pi(\delta_{j}))}$ is a straightforward computation, see 2.11 for details. Now let $\gamma\in H_{3}$. Then for $\beta\in\{\kappa^{\text{lo}},\kappa^{\text{sh}}\}$,
$\phi\circ\pi\big{(}\beta(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}\big{)}$ $=\phi\big{(}\pi(\beta(\gamma))^{\sigma(\pi(\delta_{i}))}\big{)}=\big{(}\phi\circ\pi(\beta(\gamma)))^{\sigma(\pi(\delta_{i}))}=\gamma^{\sigma(\pi(\delta_{i}))}.$
Further, for all $\gamma\in H_{3}$,
$\pi\big{(}\kappa^{\text{lo}}(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}\big{)}$ $=\pi\big{(}\kappa^{\text{lo}}(\gamma)\big{)}^{\sigma(\pi(\delta_{i}))}=\tau\cdot\pi\big{(}\kappa^{\text{sh}}(\gamma)\big{)}^{\sigma(\pi(\delta_{i}))}$
$=\tau\cdot\pi\big{(}\kappa^{\text{sh}}(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}\big{)}.$
The remaining assertion follows. ∎
Remark 4.14.
For all invertible $r\in\mathcal{R}$ and all $\beta\in D_{6}$, it is well-known that $w^{D}_{\beta}(r)\coloneqq\tilde{\theta}_{-\beta}^{D}(-r^{-1})\tilde{\theta}_{\alpha}^{D}(r)\tilde{\theta}_{-\beta}^{D}(-r^{-1})$ is a $\beta$-Weyl element in the $D_{6}$-grading $(\tilde{U}_{\gamma}^{D})_{\gamma\in D_{6}}$ (see [Ste16, Lemma 20 (b)] and [Wie23, 3.3.6]). Now let $\alpha\in H_{3}$ and let $\beta_{1}\coloneqq\kappa^{\text{sh}}(\alpha),\beta_{2}\coloneqq\kappa^{\text{lo}}(\alpha)\in D_{6}$ be as in 4.6. For invertible $r,s\in\mathcal{R}$, we put
$w^{H}_{\alpha}(r,s)\coloneqq w^{D}_{\beta_{1}}(r)w^{D}_{\beta_{2}}(s)=\tilde{\theta}_{-\alpha}^{H}(-r^{-1},-s^{-1})\tilde{\theta}_{\alpha}^{H}(r,s)\tilde{\theta}_{-\alpha}^{H}(-r^{-1},-s^{-1}).$
The proof of [Wie23, 2.6.6] applies verbatim in this situation, except that references to [Wie23, 1.4.10] must be replaced by 4.13. We conclude that $w^{H}_{\alpha}(r,s)$ is an $\alpha$-Weyl element with respect to $(\tilde{U}_{\gamma}^{H})_{\gamma\in H_{3}}$. Alternatively, it is a straightforward computation to verify this fact.
Lemma 4.15.
Let $\alpha,\beta\in GH_{3}$ (or $\alpha,\beta\in H_{3}$). Then there exists a positive system $\Pi$ in $D_{6}$ which contains $D_{6}\cap\pi^{-1}(\{\alpha,\beta\})$ (or $D_{6}\cap(\phi\pi)^{-1}(\{\alpha,\beta\})$).
Proof.
Let $\alpha,\beta\in GH_{3}$. Then $\alpha,\beta$ lie in a common positive system in $GH_{3}$, which must be conjugate to the system $\Pi_{GH}$ from 4.8. Hence there exist $\zeta_{1},\ldots,\zeta_{k}\in\{\rho_{1},\rho_{2},\rho_{3}\}=\Delta_{H}$ such that
$\alpha,\beta\in(\Pi_{GH})^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})}=(\pi(\Pi_{D}))^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})}.$
By 4.13, there exists $w\in\operatorname{Weyl}(D_{6})$ such that $(\pi(\Pi_{D}))^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})}=\pi((\Pi_{D})^{w})$. Then $\pi(-(\Pi_{D})^{w})=-(\pi(\Pi_{D}))^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})}$ as well, so $\pi^{-1}((\Pi_{GH})^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})})=(\Pi_{D})^{w}$. Thus $\Pi\coloneqq(\Pi_{D})^{w}$ is a positive system in $D_{6}$ with the desired property. The same choice of $\Pi$ works for $\alpha^{\prime}\coloneqq\phi(\alpha),\beta^{\prime}\coloneqq\phi(\beta)\in H_{3}$. ∎
Lemma 4.16.
Let $\bar{\pi}\colon\Phi\rightarrow\Phi^{\prime}$ be a surjective map between root systems $\Phi,\Phi^{\prime}$ such that $\Phi$ is reduced. Assume that for all $\alpha^{\prime},\beta^{\prime}\in\Phi^{\prime}$, there exists a positive system $\Pi$ in $\Phi$ which contains $\bar{\pi}^{-1}(\{\alpha^{\prime},\beta^{\prime}\})$. Let $G$ be a group which has (crystallographic) $\Phi$-commutator relations with root groups $(\bar{V}_{\alpha})_{\alpha\in\Phi}$ and put $\bar{U}_{\alpha^{\prime}}\coloneqq\bar{V}_{\bar{\pi}^{-1}(\alpha^{\prime})}$ for all $\alpha^{\prime}\in\Phi^{\prime}$. If, in the notation of 2.6,
$\bar{\pi}(\mathord{]A,B[})\subseteq\mathord{]\bar{\pi}(A),\bar{\pi}(B)[}\qquad\text{for all }A,B\subseteq\Phi,$
then $(\bar{U}_{\alpha^{\prime}})_{\alpha^{\prime}\in\Phi^{\prime}}$ satisfies Axiom 3.3 (i). If the commutator relations of $G$ are crystallographic and
$\bar{\pi}(\mathord{]A,B[}_{\text{cry}})\subseteq\mathord{]\bar{\pi}(A),\bar{\pi}(B)[}_{\text{cry}}\qquad\text{for all }A,B\subseteq\Phi,$
then $(\bar{U}_{\alpha^{\prime}})_{\alpha^{\prime}\in\Phi^{\prime}}$ satisfies the crystallographic version of Axiom 3.3 (i).
Proof.
This follows from the same arguments as in [Wie23, 2.1.29]. ∎
Remark 4.17.
If $\bar{\pi}$ is linear or (induced by) the scaling map $\phi\colon v\mapsto\frac{1}{\lVert v\rVert}v$, then the condition in 4.16 that $\bar{\pi}(\mathord{]A,B[})\subseteq\mathord{]\bar{\pi}(A),\bar{\pi}(B)[}$ (or $\bar{\pi}(\mathord{]A,B[}_{\text{cry}})\subseteq\mathord{]\bar{\pi}(A),\bar{\pi}(B)[}_{\text{cry}}$) is satisfied.
Proposition 4.18.
The group $\operatorname{Ch}(D_{6},\mathcal{R})$ has crystallographic $GH_{3}$-commutator relations with root groups $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$, and $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ is a (non-crystallographic) $H_{3}$-grading of $\operatorname{Ch}(D_{6},\mathcal{R})$.
Proof.
The groups $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ and $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ satisfy Axiom 3.3 (ii) by 4.7 and because all groups $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ are abelian. Further, by 4.16 and 4.17, $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ and $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ also satisfy Axiom 3.3 (i) and $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ even satisfies the crystallographic version of this axiom.
By 4.14, Weyl elements in $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ exist. Lastly, note that the positive systems $\Pi_{D}$ and $\Pi_{H}$ in 4.8 satisfy $\tilde{U}_{\Pi_{H}}^{H}\cap\tilde{U}_{-\Pi_{H}}^{H}=\tilde{U}_{\Pi_{D}}^{D}\cap\tilde{U}_{-\Pi_{D}}^{D}=\{1\}$ because $(\tilde{U}_{\beta}^{D})_{\beta\in D_{6}}$ is the grading of a Chevalley group (see [Ste16, Corollary 3 of Lemma 18]). By 3.6, this finishes the proof. ∎
Remark 4.19.
The groups $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ satisfy $\tilde{U}_{\Pi_{GH}}^{GH}\cap\tilde{U}_{-\Pi_{GH}}^{GH}=\{1\}$ as well, but one can show that they do not have Weyl elements.
Now that we have an explicit example of an $H_{3}$-graded group at hand, we can compute the commutator formulas in this group.
Notation 4.20.
For all non-proportional roots $\alpha,\gamma\in H_{3}$ and all $\beta\in\mathord{]\alpha,\gamma[}$, we define a map $\tilde{\psi}_{\alpha,\gamma}^{\beta}\colon(\mathcal{R}\times\mathcal{R})^{2}\rightarrow\mathcal{R}\times\mathcal{R}$ by the property that
$[\tilde{\theta}_{\alpha}^{H}(x_{1},y_{1}),\tilde{\theta}_{\gamma}^{H}(x_{2},y_{2})]_{\beta}=\tilde{\theta}_{\beta}^{H}\big{(}\tilde{\psi}_{\alpha,\gamma}^{\beta}\big{(}(x_{1},y_{1}),(x_{2},y_{2})\big{)}\big{)}$
for all $x_{1},x_{2},y_{1},y_{2}\in\mathcal{R}$. If $\mathord{]\alpha,\gamma[}$ has only one element $\beta$, we will write $\tilde{\psi}_{\alpha,\gamma}$ in place of $\tilde{\psi}_{\alpha,\gamma}^{\beta}$.
Remark 4.21.
Let $\alpha\coloneqq\rho_{2}$, $\varepsilon\coloneqq\rho_{3}\in\Delta_{H}$ and denote the corresponding $H_{2}$-quintuple by $(\alpha,\beta,\gamma,\delta,\varepsilon)$. We can determine the maps $\tilde{\psi}_{\zeta,\xi}^{\rho}$ for $\zeta,\xi\in\{\alpha,\beta,\gamma,\delta,\varepsilon\}$ and $\rho\in\mathord{]\zeta,\xi[}$ by straightforward matrix computations (see 2.11). The results are given in Figure 4.
[…]
Figure 4: The commutator relations in $\operatorname{Ch}(D_{6},R)$ for the $H_{3}$-grading $(\tilde{U}_{\zeta}^{H})_{\zeta\in H_{3}}$. See also 4.21.
Remark 4.22.
The family $(\tilde{U}_{\xi}^{GH})_{\xi\in GH_{3}}$ should be regarded as a refinement of the $H_{3}$-grading $(\tilde{U}_{\zeta}^{H})_{\zeta\in H_{3}}$ because its commutator relations are crystallographic. For example, let $(\alpha,\beta,\gamma,\delta,\varepsilon)$ be an $H_{2}$-quintuple in $GH_{3}$ and put $\xi^{\prime}\coloneqq\phi(\xi)$ for all $\xi\in GH_{3}$. Then $\tilde{U}_{\alpha^{\prime}}^{H}=\tilde{U}_{\alpha}^{GH}\tilde{U}_{\tau\alpha}^{GH}$ and $\tilde{U}_{\delta^{\prime}}^{H}=\tilde{U}_{\delta}^{GH}\tilde{U}_{\tau\delta}^{GH}$, so arbitrary $y_{\alpha^{\prime}}\in\tilde{U}_{\alpha^{\prime}}^{H}$ and $y_{\delta^{\prime}}\in\tilde{U}_{\delta^{\prime}}^{H}$ can be written as $y_{\alpha^{\prime}}=x_{\alpha}x_{\tau\alpha}$ and $y_{\delta^{\prime}}=x_{\delta}x_{\tau\delta}$ where $x_{\xi}\in\tilde{U}_{\xi}^{GH}$ for all $\xi\in\{\alpha,\tau\alpha,\delta,\tau\delta\}$. By relations 2.2 (v) and 2.2 (vi),
$[y_{\alpha^{\prime}},y_{\delta^{\prime}}]$ $=[x_{\alpha}x_{\tau\alpha},x_{\delta}x_{\tau\delta}]=[x_{\alpha},x_{\delta}x_{\tau\delta}]^{x_{\tau\delta}}[x_{\tau\alpha},x_{\delta}x_{\tau\delta}]$
$=[x_{\alpha},x_{\tau\delta}]^{x_{\tau\alpha}}[x_{\alpha},x_{\delta}]^{x_{\tau\delta}x_{\tau\alpha}}[x_{\tau\alpha},x_{\tau\delta}][x_{\tau\alpha},x_{\delta}]^{x_{\tau\delta}}.$
Using the crystallographic $GH_{3}$-commutator relations and the crystallographic adjacency relations that we can read off from Figure 3, we infer that
$[y_{\alpha^{\prime}},y_{\delta^{\prime}}]=[x_{\alpha},x_{\tau\delta}][x_{\tau\alpha},x_{\delta}]\in\tilde{U}_{\tau\gamma}^{GH}\tilde{U}_{\tau\beta}^{GH}.$
We conclude that $[\tilde{U}_{\alpha^{\prime}}^{H},\tilde{U}_{\delta^{\prime}}^{H}]\subseteq\tilde{U}_{\tau\gamma}^{GH}\tilde{U}_{\tau\beta}^{GH}$, which is a refinement of the non-crystallographic commutator relation
$[\tilde{U}_{\alpha^{\prime}}^{H},\tilde{U}_{\delta^{\prime}}^{H}]\subseteq\tilde{U}_{\beta}^{H}\tilde{U}_{\gamma}^{H}=\tilde{U}_{\beta}^{GH}\tilde{U}_{\tau\beta}^{GH}\tilde{U}_{\gamma}^{GH}\tilde{U}_{\tau\gamma}^{GH}.$
Similar computations can be performed for all other pairs of non-proportional roots. See also the commutator relations in Figure 4.
Remark 4.23.
In a similar way as in this section, one can fold a Chevalley group of type $E_{8}$ along the diagram in Figure 1(c) to obtain an $H_{4}$-graded group.
5 The action of Weyl elements in $H_{3}$-graded groups
Notation 5.1.
In this section, $(G,(U_{\alpha})_{\alpha\in H_{3}})$ denotes an $H_{3}$-graded group, $\Psi$ is a fixed $H_{2}$-subsystem of $H_{3}$ and $(\alpha,\beta,\gamma,\delta,\varepsilon)$ is a fixed $H_{2}$-quintuple in $\Psi$.
The goal of this section is 5.20, which provides formulas for the actions of squares of Weyl elements on all root groups in $G$. Namely, we will see that these actions “factor through” an action of the group $A\coloneqq\{\pm 1\}^{2}$, the so-called twisting group of $G$ (see 6.11). The existence of a twisting group is the main prerequisite that is necessary to apply the parametrisation theorem in Section 6.
Remark 5.2.
Throughout this section, we will often implicitly use 3.7 to reduce statements about $H_{3}$-graded groups to statements about its rank-2 subgroups. Similarly, we will see in 5.22 that all results in this section transfer to the case of $H_{4}$-graded groups without difficulty. Note that all rank-2 subsystem of $H_{3}$ are of type $H_{2}$, $A_{2}$ or $A_{1}\times A_{1}$. In particular, pairs of orthogonal roots are necessarily adjacent. Further, we will frequently apply many statements in this section to the $H_{2}$-quintuple $(\epsilon,\delta,\gamma,\beta,\alpha)$ in place of $(\alpha,\beta,\gamma,\delta,\epsilon)$, which is legitimate because $(\alpha,\beta,\gamma,\delta,\epsilon)$ is chosen arbitrarily.
The following results summarise what is known from the case of $A_{2}$-gradings.
Proposition 5.3.
Let $\zeta\in H_{3}$ be any root. Then the following hold:
  1. (i)
    The root group $U_{\zeta}$ is abelian even if we do not explicitly require Axiom 3.3 (ii).
  2. (ii)
    For each $b_{\zeta}\in U_{\zeta}^{\sharp}$, there exists a unique $a_{-\zeta}\in U_{-\zeta}$ such that $a_{-\zeta}b_{\zeta}a_{-\zeta}$ is an $\zeta$-Weyl element.
  3. (iii)
    If $w_{\zeta}$ is an $\zeta$-Weyl element, then there exist unique $a_{-\zeta}\in U_{-\zeta}$ and $b_{\zeta}\in U_{\zeta}$ such that $w_{\zeta}=a_{-\zeta}b_{\zeta}a_{-\zeta}$.
  4. (iv)
    Let $w_{\zeta}=a_{-\zeta}b_{\zeta}a_{-\zeta}$ be a $\zeta$-Weyl element with $a_{-\zeta}\in U_{-\zeta}$ and $b_{\zeta}\in U_{\zeta}$. Let $\xi\in H_{3}$ such that $(\zeta,\xi)$ is an $A_{2}$-pair and let $x_{\xi}\in U_{\xi}$. Then $x_{\xi}^{w_{\zeta}}=[x_{\xi},b_{\zeta}]$.
  5. (v)
    Let $w_{\zeta}=a_{-\zeta}b_{\zeta}a_{-\zeta}$ be a $\zeta$-Weyl element with $a_{-\zeta}\in U_{-\zeta}$ and $b_{\zeta}\in U_{\zeta}$. Then $(a_{-\zeta},b_{\zeta},a_{-\zeta})$ is balanced in the sense of [Wie23, 2.2.14]. That is, we have $a_{-\zeta}^{w_{\zeta}}=b_{\zeta}$, $b_{\zeta}^{w_{\zeta}}=a_{-\zeta}$ and $w_{\zeta}=b_{\zeta}a_{-\zeta}b_{\zeta}$.
Proof.
Since $\alpha$ is contained in an $A_{2}$-subsystem by 2.12, this follows from the corresponding assertions about $A_{2}$-graded groups in [Wie23, 5.4.3, 5.4.9, 5.4.10]. For the first assertion, see also 3.5. ∎
Proposition 5.4.
Let $\zeta,\xi\in H_{3}$, let $w_{\zeta}$ be an $\zeta$-Weyl element and put $w\coloneqq w_{\zeta}^{2}$. Then the following hold:
  1. (i)
    If $\Psi_{\zeta,\xi}$ is of type $A_{2}$, then $w$ acts on $U_{\xi}$ by inversion. That is, $x_{\xi}^{w}=x_{\xi}^{-1}$ for all $x_{\xi}\in U_{\xi}$.
  2. (ii)
    If $\Psi_{\zeta,\xi}$ is of type $A_{1}$ or $A_{1}\times A_{1}$, then $w$ acts trivially on $U_{\xi}$. That is, $x_{\xi}^{w}=x_{\xi}$ for all $x_{\xi}\in U_{\xi}$.
Proof.
If $\Psi_{\zeta,\xi}$ is of type $A_{1}\times A_{1}$, then $U_{\xi}$ is centralised by $U_{\zeta}$ and $U_{-\zeta}$ and hence by $w$. The remaining cases are covered by [Wie23, 5.4.13, 5.4.15]. ∎
It remains to study the more complicated case of $\zeta$-Weyl elements acting on $U_{\xi}$ where $\Psi_{\zeta,\xi}$ is of type $H_{2}$. We begin with collecting some “rank-2 results”, that is, results which hold in arbitrary $H_{2}$-graded groups (and not only those which are embeddable into an $H_{3}$-graded group).
Lemma 5.5.
Let $x_{\alpha},x^{\prime}_{\alpha}\in U_{\alpha}$ and $y_{\varepsilon},y^{\prime}_{\varepsilon}\in U_{\varepsilon}$. Then the following hold:
  1. (i)
    $[x_{\alpha},y_{\varepsilon}y^{\prime}_{\varepsilon}]_{\beta}=[x_{\alpha},y^{\prime}_{\varepsilon}]_{\beta}[x_{\alpha},y_{\varepsilon}]_{\beta}=[x_{\alpha},y_{\varepsilon}]_{\beta}[x_{\alpha},y^{\prime}_{\varepsilon}]_{\beta}$.
  2. (ii)
    $[x_{\alpha}x^{\prime}_{\alpha},y_{\varepsilon}]_{\delta}=[x_{\alpha},y_{\varepsilon}]_{\delta}[x^{\prime}_{\alpha},y_{\varepsilon}]_{\delta}$.
In particular, $[x_{\alpha},y_{\varepsilon}^{-1}]_{\beta}=[x_{\alpha},y_{\varepsilon}]_{\beta}^{-1}$ and $[x_{\alpha}^{-1},y_{\varepsilon}]_{\delta}=[x_{\alpha},y_{\varepsilon}]_{\delta}^{-1}$.
Proof.
Using 2.2 (vi), we compute
$[x_{\alpha},y_{\varepsilon}y^{\prime}_{\varepsilon}]$ $=[x_{\alpha},y^{\prime}_{\varepsilon}][x_{\alpha},y_{\varepsilon}]^{y^{\prime}_{\varepsilon}}$
$=[x_{\alpha},y^{\prime}_{\varepsilon}]_{\beta}[x_{\alpha},y^{\prime}_{\varepsilon}]_{\gamma}[x_{\alpha},y^{\prime}_{\varepsilon}]_{\delta}\big{(}[x_{\alpha},y_{\varepsilon}]_{\beta}[x_{\alpha},y_{\varepsilon}]_{\gamma}[x_{\alpha},y_{\varepsilon}]_{\delta}\big{)}^{y^{\prime}_{\varepsilon}}.$
By 2.2 (i), $([x_{\alpha},y_{\varepsilon}]_{\beta}[x_{\alpha},y_{\varepsilon}]_{\gamma}[x_{\alpha},y_{\varepsilon}]_{\delta})^{y^{\prime}_{\varepsilon}}$ is equal to
$\left[y^{\prime}_{\varepsilon},[x_{\alpha},y_{\varepsilon}]_{\beta}^{-1}\right]_{\delta}\left[y^{\prime}_{\varepsilon},[x_{\alpha},y_{\varepsilon}]_{\beta}^{-1}\right]_{\gamma}[x_{\alpha},y_{\varepsilon}]_{\beta}\left[y^{\prime}_{\varepsilon},[x_{\alpha},y_{\varepsilon}]_{\gamma}^{-1}\right][x_{\alpha},y_{\varepsilon}]_{\gamma}[x_{\alpha},y_{\varepsilon}]_{\delta}.$
Note that the only factor in this product that lies in $U_{\beta}$ is $[x_{\alpha},y_{\varepsilon}]_{\beta}$. Using relation 2.2 (iii), we can reorder the factors to obtain that
$[x_{\alpha},y_{\varepsilon}y^{\prime}_{\varepsilon}]=[x_{\alpha},y^{\prime}_{\varepsilon}]_{\beta}[x_{\alpha},y_{\varepsilon}]_{\beta}z_{\gamma}z_{\delta}$
for some $z_{\gamma}\in U_{\gamma}$ and $z_{\delta}\in U_{\delta}$. By 3.10, the first equation in i follows. Since $U_{\beta}$ is abelian, the second equation in i holds as well. Finally, assertion ii follows from i (applied to the $H_{2}$-quintuple $(\epsilon,\delta,\gamma,\beta,\alpha)$) and 3.13:
$[x_{\alpha}x_{\alpha}^{\prime},y_{\epsilon}]_{\delta}$ $=[y_{\epsilon},x_{\alpha}x_{\alpha}^{\prime}]_{\delta}^{-1}=\big{(}[y_{\epsilon},x_{\alpha}^{\prime}]_{\delta}[y_{\epsilon},x_{\alpha}]_{\delta}\big{)}^{-1}=[x_{\alpha},y_{\epsilon}]_{\delta}[x_{\alpha}^{\prime},y_{\epsilon}]_{\delta},$
as desired. ∎
Lemma 5.6.
Let $x_{\alpha},x_{\alpha}^{\prime}\in U_{\alpha}$ and $y_{\delta},y_{\delta}^{\prime}\in U_{\delta}$. Then the following hold:
  1. (i)
    $[x_{\alpha},y_{\delta}y_{\delta}^{\prime}]_{\beta}=[x_{\alpha},y_{\delta}]_{\beta}[x_{\alpha},y_{\delta}^{\prime}]_{\beta}$.
  2. (ii)
    $[x_{\alpha}x_{\alpha}^{\prime},y_{\delta}]_{\beta}=[x_{\alpha},y_{\delta}]_{\beta}[x_{\alpha}^{\prime},y_{\delta}]_{\beta}\big{[}[x_{\alpha},y_{\delta}]_{\gamma},x_{\alpha}^{\prime}]$.
  3. (iii)
    $[x_{\alpha},y_{\delta}y_{\delta}]_{\beta}=[x_{\alpha},y_{\delta}]_{\beta}[x_{\alpha},y_{\delta}^{\prime}]_{\beta}\big{[}[x_{\alpha},y_{\delta}]_{\beta},y_{\delta}^{\prime}]$.
  4. (iv)
    $[x_{\alpha}x_{\alpha}^{\prime},y_{\delta}]_{\gamma}=[x_{\alpha},y_{\delta}]_{\gamma}[x_{\alpha}^{\prime},y_{\delta}]_{\gamma}$.
Proof.
This is a similar computation as in 5.5. See [Wie23, 7.4.4] for details. ∎
Lemma 5.7.
Let $x_{\varepsilon}\in U_{\varepsilon}$ and let $w_{\alpha}=a_{-\alpha}b_{\alpha}a_{-\alpha}$ be an $\alpha$-Weyl element for $b_{\alpha}\in U_{\alpha}^{\sharp}$ and $a_{-\alpha}\in U_{-\alpha}^{\sharp}$. Then the following hold:
  1. (i)
    $x_{\varepsilon}^{w_{\alpha}}=[b_{\alpha},x_{\varepsilon}^{-1}]_{\beta}$.
  2. (ii)
    $x_{\varepsilon}^{w_{\alpha}^{2}}=\left[a_{-\alpha},[b_{\alpha},x_{\varepsilon}^{-1}]_{\beta}^{-1}\right]_{\varepsilon}=\left[a_{-\alpha},[b_{\alpha},x_{\varepsilon}]_{\beta}\right]_{\varepsilon}$.
  3. (iii)
    $x_{\varepsilon}^{w_{\alpha}^{-1}}=[b_{\alpha}^{-1},x_{\varepsilon}^{-1}]_{\beta}$.
  4. (iv)
    $x_{\varepsilon}=\left[a_{-\alpha}^{-1},[b_{\alpha},x_{\varepsilon}]_{\beta}\right]_{\varepsilon}$.
Proof.
Let $b_{\alpha}\in U_{\alpha}^{\sharp}$ and $a_{-\alpha}\in U_{-\alpha}$ such that $w_{\alpha}=a_{-\alpha}b_{\alpha}a_{-\alpha}$ is an $\alpha$-Weyl element and let $x_{\varepsilon}\in U_{\varepsilon}$. Using 2.2 (i), we compute
$x_{\varepsilon}^{w_{\alpha}}$ $=x_{\varepsilon}^{a_{-\alpha}b_{\alpha}a_{-\alpha}}=x_{\varepsilon}^{b_{\alpha}a_{-\alpha}}=([b_{\alpha},x_{\varepsilon}^{-1}]x_{\varepsilon})^{a_{-\alpha}}$
$=[b_{\alpha},x_{\varepsilon}^{-1}]_{\beta}^{a_{-\alpha}}[b_{\alpha},x_{\varepsilon}^{-1}]_{\gamma}^{a_{-\alpha}}[b_{\alpha},x_{\varepsilon}^{-1}]_{\delta}^{a_{-\alpha}}x_{\varepsilon}$
$=\left[a_{-\alpha},[b_{\alpha},x_{\varepsilon}^{-1}]_{\beta}^{-1}\right][b_{\alpha},x_{\varepsilon}^{-1}]_{\beta}\left[a_{-\alpha},[b_{\alpha},x_{\varepsilon}^{-1}]_{\gamma}^{-1}\right][b_{\alpha},x_{\varepsilon}^{-1}]_{\gamma}$
$\qquad\mathord{}\cdot\big{[}a_{-\alpha},[b_{\alpha},x_{\varepsilon}^{-1}]_{\delta}^{-1}\big{]}[b_{\alpha},x_{\varepsilon}^{-1}]_{\delta}x_{\varepsilon}.$
Just as in the proof of 5.5, reordering the elements in the above product does not produce any new factors in $U_{\beta}$ other than $[b_{\alpha},x_{\varepsilon}^{-1}]_{\beta}$. Since $x_{\varepsilon}^{w_{\alpha}}\in U_{\varepsilon^{\sigma(\alpha)}}=U_{\beta}$, we conclude that (i) holds. By applying (i) to the $\alpha$-Weyl element $w_{\alpha}^{-1}=a_{-\alpha}^{-1}b_{\alpha}^{-1}a_{-\alpha}^{-1}$, we obtain (iii).
For (ii), recall from 3.9 that $w_{\alpha}$ is also a ($-\alpha$)-Weyl element with factorisation $w_{\alpha}=b_{\alpha}a_{-\alpha}a_{-\alpha}^{w_{\alpha}}$. Hence by two applications of (i),
$x_{\varepsilon}^{w_{\alpha}^{2}}$ $=[b_{\alpha},x_{\varepsilon}^{-1}]_{\beta}^{w_{\alpha}}=\left[a_{-\alpha},[b_{\alpha},x_{\varepsilon}^{-1}]_{\beta}^{-1}\right]_{\varepsilon}$
The second equality in (ii) now follows by 5.5.
Finally, we have
$x_{\varepsilon}=x_{\varepsilon}^{w_{\alpha}w_{\alpha}^{-1}}=[b_{\alpha},x_{\varepsilon}^{-1}]_{\beta}^{w_{\alpha}^{-1}}.$
Using the factorisation $w_{\alpha}^{-1}=(a_{-\alpha}^{-1})^{w_{\alpha}}a_{-\alpha}^{-1}b_{\alpha}^{-1}$ of $w_{\alpha}^{-1}$ as a $(-\alpha)$-Weyl element (see 3.9), an application of (i) to the $H_{2}$-quintuple $(-\alpha,\varepsilon,\delta,\gamma,\beta)$ yields $x_{\varepsilon}=\left[a_{-\alpha}^{-1},[b_{\alpha},x_{\varepsilon}]_{\beta}\right]_{\varepsilon}$. This finishes the proof. ∎
Proposition 5.8.
Let $b_{\alpha}\in U_{\alpha}^{\sharp}$. Then the map $U_{\varepsilon}\rightarrow U_{\beta},~{}x_{\varepsilon}\mapsto[b_{\alpha},x_{\varepsilon}]_{\beta}$ is an isomorphism of groups. In particular, for each $x_{\beta}\in U_{\beta}$ there exist $x_{\alpha}\in U_{\alpha}$ and $x_{\varepsilon}\in U_{\varepsilon}$ such that $x_{\beta}=[x_{\alpha},x_{\varepsilon}]_{\beta}$.
Proof.
Let $w_{\alpha}$ be an $\alpha$-Weyl element in $U_{-\alpha}b_{\alpha}U_{-\alpha}$. By 5.5 and 5.7, the map in the assertion is $x_{\varepsilon}\mapsto(x_{\varepsilon}^{w_{\alpha}})^{-1}$, which is an isomorphism by the definition of Weyl elements and because $U_{\beta}$ is abelian. ∎
From now on, we will use in a crucial way that the grading of $G$ has rank (at least) 3. At first, we show that the action of the square of a $\zeta$-Weyl element (for a fixed root $\zeta$) does not depend on the choice of the Weyl element.
Proposition 5.9.
Let $w_{\alpha}$, $v_{\alpha}$ be two $\alpha$-Weyl elements and let $\zeta\in\Psi\setminus\{\alpha,-\alpha\}$. Then $w_{\alpha}^{2}$ and $v_{\alpha}^{2}$ act identically on $U_{\zeta}$ and $w_{\alpha}^{4}$ acts trivially on $U_{\zeta}$.
Proof.
Since $w_{\alpha}$ is also a $(-\alpha)$-Weyl element by 3.9 (iii), we may assume that $\zeta\in\{\beta,\gamma,\delta,\varepsilon\}$. In each case, we will derive a formula for the action of $w_{\alpha}^{2}$ on $U_{\zeta}$ which does not depend on the choice of $w_{\alpha}$, which proves the first assertion. It then follows that $w_{\alpha}^{2}$ and $(w_{\alpha}^{-1})^{2}$ act identically on $U_{\zeta}$, which says precisely that $w_{\alpha}^{4}$ acts trivially on $U_{\zeta}$.
Recall from 2.12 that $\zeta$ lies in a unique $H_{2}$-subsystem $\tilde{\Psi}$ different from $\Psi$. We choose an $H_{2}$-quintuple $(\tilde{\alpha},\tilde{\beta},\tilde{\gamma},\tilde{\delta},\tilde{\varepsilon})$ in $\tilde{\Psi}$ such that $\zeta=\tilde{\beta}$. Since any rank-2 subsystem is uniquely determined by two non-proportional roots in this subsystem, we have $\Psi\cap\tilde{\Psi}=\{\zeta,-\zeta\}=\{\tilde{\beta},-\tilde{\beta}\}$ and
$\Psi,\Psi_{\alpha,\tilde{\alpha}},\Psi_{\alpha,\tilde{\varepsilon}},\Psi_{\alpha,\tilde{\gamma}},\Psi_{\alpha,-\tilde{\delta}}\text{ are pairwise distinct.}$. (5.10)
Now let $x_{\zeta}\in U_{\zeta}$ and put $w\coloneqq w_{\alpha}^{2}$. By 5.8 we can write $x_{\zeta}=[x_{\tilde{\alpha}},x_{\tilde{\varepsilon}}]_{\zeta}$ for some $x_{\tilde{\alpha}}\in U_{\tilde{\alpha}}$ and $x_{\tilde{\varepsilon}}\in U_{\tilde{\varepsilon}}$. Hence by 3.13,
$x_{\zeta}^{w}=[x_{\tilde{\alpha}},x_{\tilde{\varepsilon}}]_{\zeta}^{w}=[x_{\tilde{\alpha}}^{w},x_{\tilde{\varepsilon}}^{w}]_{\zeta}.$
If $\Psi_{\alpha,\tilde{\alpha}}$ and $\Psi_{\alpha,\tilde{\varepsilon}}$ are of type $A_{1}\times A_{1}$ or $A_{2}$, then 5.4 provides formulas for the action of $w$ on $U_{\tilde{\alpha}}$ and $U_{\tilde{\varepsilon}}$ which do not depend on the choice of $w_{\alpha}$. Thus the first assertion holds in this case. Further, since $\alpha$ lies in exactly two $H_{2}$-subsystems by 2.12, it follows from (5.10) that at most one of the subsystems $\Psi_{\alpha,\tilde{\alpha}}$ and $\Psi_{\alpha,\tilde{\varepsilon}}$ can be of type $H_{2}$.
We remain with the case that one of the subsystems $\Psi_{\alpha,\tilde{\alpha}}$ and $\Psi_{\alpha,\tilde{\varepsilon}}$ is of type $H_{2}$. Again by applying 5.8, but to the $H_{2}$-pair $(\tilde{\gamma},-\tilde{\delta})$, we can write $x_{\zeta}=[x_{\tilde{\gamma}},x_{-\tilde{\delta}}]_{\zeta}$ for some $x_{\tilde{\gamma}}\in U_{\tilde{\gamma}}$ and $x_{-\tilde{\delta}}\in U_{-\tilde{\delta}}$. As above, we then have $x_{\zeta}^{w}=[x_{\tilde{\gamma}}^{w},x_{-\tilde{\delta}}^{w}]_{\zeta}$. By our case assumption, 2.12 and (5.10), both $\Psi_{\alpha,\tilde{\gamma}}$ and $\Psi_{\alpha,-\tilde{\delta}}$ are of type $A_{1}\times A_{1}$ or $A_{2}$. By 5.4, we conclude that the first assertion holds in this case as well, which finishes the proof. ∎
We may already suspect from the proof of 5.9 that the action of $w_{\alpha}^{2}$ on $U_{\zeta}$ depends on the choice of $\zeta\in\Psi\setminus\{\alpha,-\alpha\}$. However, we will see in 5.15 that only the two cases in the following definition are distinct, and in 5.17 that the actions in these two cases differ by inversion.
Definition 5.11.
Let $\zeta,\xi\in H_{3}$. We say that $\xi$ is in involution position with respect to $\zeta$ if $(\zeta,\xi)$ or $(\zeta,-\xi)$ is an $H_{2}$-pair. We say that $\xi$ is in inverted involution position with respect to $\zeta$ if $\Psi_{\xi,\zeta}$ is of type $H_{2}$ but $\xi$ is not in involution position with respect to $\zeta$. For any $\rho\in H_{3}$, we say that $\xi$, $\zeta$ are in the same position with respect to $\rho$ if both are in involution position with respect to $\rho$ or both are in inverted involution position with respect to $\rho$.
Remark 5.12.
In the notation of this section, $\beta,-\beta,\varepsilon$ and $-\varepsilon$ are in involution position with respect to $\alpha$ while $\gamma,-\gamma,\delta$ and $-\delta$ are in inverted involution position with respect to $\alpha$. Further, for any $w\in\operatorname{Weyl}(H_{3})$, the root $\xi$ is in (inverted) involution position with respect to $\zeta$ if and only if $\xi^{w}$ is in (inverted) involution position with respect to $\zeta^{w}$. Moreover, $\xi$ is in (inverted) involution position with respect to $\zeta$ if and only if $\zeta$ is in (inverted) involution position with respect to $\xi$.
Notation 5.13.
For the rest of this section, we denote by $\tilde{\Psi}$ the unique $H_{2}$-subsystem different from $\Psi$ which contains $\alpha$. Further, we choose an $H_{2}$-quintuple $(\alpha,\tilde{\beta},\tilde{\gamma},\tilde{\delta},\tilde{\varepsilon})$ in $\tilde{\Psi}$ starting with $\alpha$. Note that $\Psi\cap\tilde{\Psi}=\{\alpha,-\alpha\}$. Finally, we fix a positive system $\Pi$ in $H_{3}$ containing $\{\alpha,\beta,\gamma,\delta,\varepsilon\}$, whose elements will henceforth be called “positive roots”.
Before we can prove 5.15, we need the following lemma about $H_{3}$.
Lemma 5.14.
There exist unique positive roots $\eta,\vartheta\in H_{3}\cap\alpha^{\perp}$ such that
$\beta^{\sigma(\eta)}$ $=\tilde{\beta},$ $\gamma^{\sigma(\eta)}$ $=\tilde{\gamma},$ $\delta^{\sigma(\eta)}$ $=\tilde{\delta},$ $\varepsilon^{\sigma(\eta)}$ $=\tilde{\varepsilon},$
$\beta^{\sigma(\vartheta)}$ $=-\tilde{\varepsilon},$ $\gamma^{\sigma(\vartheta)}$ $=-\tilde{\delta},$ $\delta^{\sigma(\vartheta)}$ $=-\tilde{\gamma},$ $\varepsilon^{\sigma(\vartheta)}$ $=-\tilde{\beta}.$
Further, $\eta$ and $\vartheta$ are the only positive roots in $H_{3}\cap\alpha^{\perp}$.
Proof.
By 2.12 iii, there are exactly two positive roots in $H_{3}\cap\alpha^{\perp}$, so the uniqueness assertion is trivial. Let $\rho$ denote any root in $H_{3}\cap\alpha^{\perp}$. Then $\rho\notin\Psi$ and hence $\rho\notin\langle\Psi\rangle_{\mathbb{R}}$. By the defining formula for reflections, it follows that for all $\xi\in\Psi$, we have either $\xi^{\sigma(\rho)}=\xi$ or $\xi^{\sigma(\rho)}\notin\Psi$. At the same time, by 2.12 iii, $\sigma_{\rho}$ cannot fix all roots in $\Psi$. Since $\alpha^{\sigma(\rho)}=\alpha$, we conclude that $\Psi^{\sigma(\rho)}$ is the unique $H_{2}$-subsystem of $H_{3}$ containing $\alpha$ which is different from $\Psi$. That is, $\Psi^{\sigma(\rho)}=\tilde{\Psi}$. Since $\sigma_{\rho}$ is orthogonal and $\alpha^{\sigma(\rho)}=\alpha$, we must have $\beta^{\sigma(\rho)}\in\{\tilde{\beta},-\tilde{\varepsilon}\}$.
Now denote by $\eta$ and $\vartheta$ the two positive roots in $H_{3}\cap\alpha^{\perp}$. For a contradiction, suppose that $\beta^{\sigma(\eta)}=\beta^{\sigma(\vartheta)}$. Then $\sigma_{\eta}^{-1}\sigma_{\vartheta}$ is an orthogonal map of determinant $1$ which fixes the hyperplane $\langle\Psi\rangle_{\mathbb{R}}$ in $\langle H_{3}\rangle_{\mathbb{R}}$ pointwise. It follows that $\sigma_{\eta}=\sigma_{\vartheta}$, so $\eta=\vartheta$ because both roots are positive. This is a contradiction, so $\beta^{\sigma(\eta)}\neq\beta^{\sigma(\vartheta)}$. By interchanging the roles of $\eta$ and $\vartheta$ if necessary, we conclude that $\beta^{\sigma(\eta)}=\tilde{\beta}$ and $\beta^{\sigma(\vartheta)}=-\tilde{\varepsilon}$. The remaining values of $\sigma_{\eta}$ and $\sigma_{\vartheta}$ on $\Psi$ are uniquely determined by their values on $\alpha$ and $\beta$, so the assertion follows. ∎
Proposition 5.15.
Let $\xi,\zeta\in\Psi\cup\tilde{\Psi}\setminus\{\alpha,-\alpha\}$ be non-proportional and let $w_{\xi}\in M_{\xi},~{}w_{\zeta}\in M_{\zeta}$ be Weyl elements. Assume that $\xi$ and $\zeta$ are in the same position with respect to $\alpha$. Then the action of $w_{\xi}^{2}w_{\zeta}^{2}$ on $U_{\alpha}$ is trivial. Equivalently, the actions of $w_{\xi}^{2}$ and $w_{\zeta}^{2}$ on $U_{\alpha}$ are identical.
Proof.
Let $\eta,\vartheta\in H_{3}\in\alpha^{\perp}$ be as in 5.14 and choose Weyl elements $w_{\eta}\in M_{\eta}$ and $w_{\vartheta}\in M_{\vartheta}$. At first, we consider the case that $\xi$ and $\zeta$ lie in a common $H_{2}$-subsystem. Without loss of generality, we assume that both lie in $\Psi$. Further, by 3.9 (iii) and 5.9, we may assume that $\zeta\neq\xi$ and that both $\xi,\zeta$ are positive roots. Thus either $\{\xi,\zeta\}=\{\beta,\varepsilon\}$ or $\{\xi,\zeta\}=\{\gamma,\delta\}$. In any case, $\xi^{\sigma(\eta)\sigma(\vartheta)}=-\zeta$ by 5.14, so it follows from 3.9 (ii) that
$w_{\xi}^{w_{\eta}w_{\vartheta}}\in M_{\xi^{\sigma(\eta)\sigma(\vartheta)}}=M_{-\zeta}=M_{\zeta}.$
Thus by 5.9, $w_{\zeta}^{2}$ and $(w_{\xi}^{w_{\vartheta}w_{\eta}})^{2}$ act identically on $U_{\alpha}$. Here $w_{\vartheta}$ and $w_{\eta}$ act trivial on $U_{\alpha}$ because $\vartheta,-\vartheta,\eta$ and $-\eta$ are orthogonal and hence adjacent to $\alpha$. We conclude that $w_{\zeta}^{2}$ and $w_{\xi}^{2}$ act identically on $U_{\alpha}$, as desired.
Now we consider the case that $\xi$ and $\zeta$ are not contained in a common $H_{2}$-subsystem. Then $\zeta$ and $\xi^{\sigma_{\eta}}$ lie in a common $H_{2}$-subsystem and, by 5.12, are in the same position with respect to $\alpha$. Therefore, by the previous case, the actions of $w_{\zeta}^{2}$ and $(w_{\xi}^{w_{\eta}})^{2}$ on $U_{\alpha}$ are identical. Since $w_{\eta}$ acts trivially on $U_{\alpha}$, we infer that $w_{\zeta}^{2}$ and $w_{\xi}^{2}$ act identically on $U_{\alpha}$, as desired. ∎
Again, we need an elementary lemma about $H_{3}$ before we can prove 5.17.
Lemma 5.16.
There exists $\rho\in\{\tilde{\gamma},\tilde{\delta}\}$ such that both $\Psi_{\rho,\beta}$ and $\Psi_{\rho,\gamma}$ are of type $A_{2}$.
Proof.
Note that if $\rho$ has the desired properties, then $-\rho$ has the desired properties for $(\alpha,\tilde{\beta},\tilde{\gamma},\tilde{\delta},\tilde{\varepsilon})$ in 5.13 replaced by $(\alpha,-\tilde{\varepsilon},-\tilde{\delta},-\tilde{\gamma},-\tilde{\beta})$. Thus for each fixed root $\alpha$, there are only four cases to consider: Namely, there are two ways to choose $\Psi$ and two ways to choose an $H_{2}$-quintuple starting from $\alpha$ in $\Psi$. It is a straightforward computation (see 2.11) to prove the assertion in each of these cases for one fixed choice of $\alpha$. Then it follows from 2.9 that it holds for all possible choices of $\alpha$. ∎
Proposition 5.17.
Let $\xi,\zeta\in\Psi\cup\tilde{\Psi}\setminus\{\alpha,-\alpha\}$ and let $w_{\xi}\in M_{\xi},~{}w_{\zeta}\in M_{\zeta}$ be Weyl elements. Assume that $\xi$ and $\zeta$ are not in the same position with respect to $\alpha$. Then $w_{\xi}^{2}w_{\zeta}^{2}$ acts on $U_{\alpha}$ by inversion. Equivalently, the actions of $w_{\xi}^{2}$ and $w_{\zeta}^{2}$ on $U_{\alpha}$ differ by inversion.
Proof.
If $w_{\xi}^{2}w_{\zeta}^{2}$ acts on $U_{\alpha}$ by inversion, then so does $w_{\zeta}^{-2}w_{\xi}^{-2}$. Hence we may assume that $\xi$ is in involution position with respect to $\alpha$ and that $\zeta$ is in inverted involution position. By 5.15, we can thus assume that $\xi=\beta$ and $\zeta=\gamma$.
Let $x_{\alpha}\in U_{\alpha}$ be arbitrary. Using 5.8, we can write $x_{\alpha}=[x_{\beta},x_{-\gamma}]_{\alpha}$ for some $x_{\beta}\in U_{\beta},~{}x_{-\gamma}\in U_{-\gamma}$. Choose $a_{-\beta}\in U_{-\beta}$ and $b_{\beta}\in U_{\beta}$ such that $w_{\beta}=a_{-\beta}b_{\beta}a_{-\beta}$. Applying 5.7 (ii) to the $(-\beta)$-Weyl element $w_{\beta}=b_{\beta}a_{-\beta}b_{\beta}^{w_{\beta}}$ from 3.9 (iii), we see that
$x_{\alpha}^{w_{\beta}^{2}}=\big{[}b_{\beta},[a_{-\beta},x_{\alpha}]_{-\gamma}\big{]}_{\alpha}.$
Now choose $\rho\in\{\tilde{\gamma},\tilde{\delta}\}$ as in 5.16 and choose some $\rho$-Weyl element $w_{\rho}$. Then, by 5.4, $w_{\rho}^{2}$ acts on $U_{\beta}$ and $U_{\gamma}$ by inversion. Since $w_{\gamma}^{2}$ and $w_{\rho}^{2}$ act identically on $U_{\alpha}$ by 5.15, we conclude that
$x_{\alpha}^{w_{\beta}^{2}w_{\gamma}^{2}}$ $=\big{[}b_{\beta},[a_{-\beta},x_{\alpha}]_{-\gamma}\big{]}_{\alpha}^{w_{\rho}^{2}}=\big{[}b_{\beta}^{w_{\rho}^{2}},[a_{-\beta},x_{\alpha}]_{-\gamma}^{w_{\rho}^{2}}\big{]}_{\alpha}=\left[b_{\beta}^{-1},[a_{-\beta},x_{\alpha}]_{-\gamma}^{-1}\right]_{\alpha}.$
Combining 5.5 and 5.7 (iv), we infer that
$x_{\alpha}^{w_{\beta}^{2}w_{\gamma}^{2}}=\left[b_{\beta}^{-1},[a_{-\beta},x_{\alpha}]_{-\gamma}\right]_{\alpha}^{-1}=x_{\alpha}^{-1}.$
This shows that $w_{\beta}^{2}w_{\gamma}^{2}$ acts on $U_{\alpha}$ by inversion, as desired. ∎
Before we can summarise the results of this section in 5.20, we need to introduce the following notation.
Definition 5.18.
Let $\rho\in H_{3}$. The twisting involution on $U_{\rho}$ is the map $U_{\rho}\rightarrow U_{\rho},x_{\rho}\mapsto x_{\rho}^{*_{\rho}}$ defined by $x_{\rho}^{*_{\rho}}\coloneqq x_{\rho}^{w}$ where $\zeta$ is any root in involution position with respect to $\rho$, $w_{\zeta}$ is any $\zeta$-Weyl element and $w\coloneqq w_{\zeta}^{2}$. By 5.15, the twisting involution does not depend on the choices of $\zeta$ and $w_{\zeta}$. We will usually write $x_{\rho}^{*}$ in place of $x_{\rho}^{*_{\rho}}$.
Lemma 5.19.
Let $\rho,\xi\in H_{3}$, let $w_{\xi}$ be a $\xi$-Weyl element and put $\bar{\rho}\coloneqq\rho^{\sigma(\xi)}$. Then $(x_{\rho}^{*_{\rho}})^{w_{\xi}}=(x_{\rho}^{w_{\xi}})^{*_{\bar{\rho}}}$ and $(x_{\rho}^{*_{\rho}})^{-1}=(x_{\rho}^{-1})^{*_{\rho}}$ for all $x_{\rho}\in U_{\rho}$.
Proof.
Choose a root $\zeta$ in involution position with respect to $\rho$ and put $w\coloneqq w_{\zeta}^{2}$ for some $\zeta$-Weyl element $w_{\zeta}$. Then $v\coloneqq w^{w_{\xi}}=(w_{\zeta}^{w_{\xi}})^{2}$ is the square of a $\zeta^{\sigma(\xi)}$-Weyl element. Since $\zeta^{\sigma(\xi)}$ is in involution position with respect to $\rho^{\sigma(\xi)}=\bar{\rho}$, we infer that
$(x_{\rho}^{*_{\rho}})^{w_{\xi}}=(x_{\rho}^{w})^{w_{\xi}}=(x_{\rho}^{w_{\xi}})^{v}=(x_{\rho}^{w_{\xi}})^{*_{\bar{\rho}}}$
This proves the first assertion. The second assertion is clear. ∎
Proposition 5.20.
Let $\xi,\zeta\in H_{3}$, let $x_{\xi}\in U_{\xi}$, let $w_{\zeta}$ be a $\zeta$-Weyl element and put $w\coloneqq w_{\zeta}^{2}$. Then for all $x_{\xi}\in U_{\xi}$, we have
$x_{\xi}^{w}=\begin{cases}x_{\xi}&\text{if }\Psi_{\xi,\zeta}=A_{1}\text{ or }\Psi_{\xi,\zeta}=A_{1}\times A_{1},\\ x_{\xi}^{-1}&\text{if }\Psi_{\xi,\zeta}=A_{2},\\ x_{\xi}^{*}&\text{if }\zeta\text{ is in involution position with respect to }\xi,\\ (x_{\xi}^{*})^{-1}&\text{if }\zeta\text{ is in inverted involution position with respect to }\xi.\end{cases}$
Proof.
This follows from 5.45.17 and 5.18. ∎
Example 5.21.
Assume that $G=\operatorname{Ch}(D_{6},\mathcal{R})$ for a commutative ring $\mathcal{R}$ and that $(U_{\alpha})_{\alpha\in H_{3}}$ is the $H_{3}$-grading constructed in 4.10. Then a straightforward matrix computation (see 2.11) shows that for all $\alpha\in H_{3}$ and all $x,y\in\mathcal{R}$, the twisting involution is given by $\tilde{\theta}_{\alpha}^{H}(x,y)^{*}=\tilde{\theta}_{\alpha}^{H}(-x,y)$.
Remark 5.22.
Let $\xi,\zeta$ be two roots in $H_{4}$. Then either $\Psi_{\xi,\zeta}$ is of type $A_{1}$, $A_{2}$ or $A_{1}\times A_{1}$, or $\xi,\zeta$ lie in a common subsystem of type $H_{3}$. It follows from this observation that the twisting involution, defined in the same way as in 5.18, is also well-defined in $H_{4}$-graded groups, and that 5.20 holds in this case as well.
6 The parametrisation theorem
Notation 6.1.
In this section, we fix a root base $\Delta=(\rho_{1},\rho_{2},\rho_{3})$ of $H_{3}$ in standard order and an $H_{3}$-graded group $(G,(U_{\alpha})_{\alpha\in H_{3}})$. For each $\delta\in\Delta$, we fix a $\delta$-Weyl element $w_{\delta}$ and we put $w_{-\delta}\coloneqq w_{\delta}^{-1}$. For any word $\bar{\delta}=(\delta_{1},\ldots,\delta_{k})$ over Δ($\Delta\cup(-\Delta)$, we put $w_{\bar{\delta}}\coloneqq w_{\delta_{1}}\cdots w_{\delta_{k}}$ and $\sigma_{\bar{\delta}}\coloneqq\sigma(\bar{\delta})\coloneqq\sigma_{\delta_{1}}\cdots\sigma_{\delta_{k}}$.
In this section, we state the parametrisation theorem for $\Phi$-graded groups for the special case $\Phi=H_{3}$ and give a sketch of its proof (6.15). More details and a complete proof of the parametrisation theorem in full generality can be found in [Wie23, Chapter 4]. Combining this theorem with the results from the previous section, we obtain a first parametrisation of $(G,(U_{\alpha})_{\alpha\in H_{3}})$, see 6.19. This parametrisation will be refined in the following sections using the blueprint technique.
Remark 6.2.
The crucial property of $H_{3}$ that simplifies the proof of the parametrisation theorem is that any pair $\alpha$, $\beta$ of orthogonal roots is automatically adjacent. In particular, any $\beta$-Weyl element centralises $U_{\alpha}$.
Definition 6.3.
For any word $\bar{\alpha}=(\alpha_{1},\ldots,\alpha_{m})$ over Δ($\Delta\cup(-\Delta)$, we define the corresponding inverse word to be $\bar{\alpha}^{-1}\coloneqq(-\alpha_{m},\ldots,-\alpha_{1})$. For any root $\alpha$, a $\Delta$-expression of $\alpha$ is a word over Δ($\Delta\cup(-\Delta)$ of the form $(\bar{\delta}^{-1},\delta,\bar{\delta})$ where $\delta\in\Delta$ and $\bar{\delta}$ is a word over Δ($\Delta\cup(-\Delta)$ such that $\delta^{\sigma(\bar{\delta})}=\alpha$.
Remark 6.4.
If $\bar{\beta}$ is a $\Delta$-expression of a root $\beta\in H_{3}$, then $w_{\bar{\beta}}$ is a $\beta$-Weyl element by 3.9 (ii).
The purpose of the parametrisation theorem is to establish the existence of a parametrisation of $(G,(U_{\alpha})_{\alpha\in H_{3}})$. In order to make this precise, we need to introduce the notions of parity maps and parameter systems.
Notation 6.5.
Let $X$ be any set. For all $s,t\in X$ and $m\in\mathbb{N}_{0}$, we denote by $P_{m}(s,t)$ the word $(u_{1},\ldots,u_{m})$ with $u_{k}=s$ if $k$ is odd and $u_{k}=t$ if $k$ is even.
Definition 6.6.
Let $A$ be an abelian group. A $\Delta$-parity map with values in $A$ is a map $\eta\colon H_{3}\times\Delta\rightarrow A,(\alpha,\delta)\mapsto\eta_{\alpha,\delta}=\eta(\alpha,\delta)$. Given any $\Delta$-parity map $\eta$, we define $\eta(\alpha,-\delta)\coloneqq\eta(\alpha^{\sigma(\delta)},\delta)^{-1}$ for all $\alpha\in H_{3}$, $\delta\in\Delta$ and for any word $\bar{\delta}=(\delta_{1},\ldots,\delta_{m})$ over Δ($\Delta\cup(-\Delta)$, we define $\eta_{\alpha,\bar{\delta}}=\eta(\alpha,\bar{\delta})$ to be
$\prod_{i=1}^{m}\eta(\alpha^{\sigma(\delta_{1},\dots,\delta_{i-1})},\delta_{i})=\eta(\alpha,\delta_{1})\cdot\eta(\alpha^{\sigma(\delta_{1})},\delta_{2})\cdots\eta(\alpha^{\sigma(\delta_{1},\dots,\delta_{m-1})},\delta_{m})$
(with the convention that $\eta_{\alpha,\emptyset}=1_{A}$ for the empty word $\emptyset$). When $\Delta$ and $A$ are clear from the context, we will simply say that $\eta$ is a parity map. If $\eta(\alpha,P_{m}(\beta,\gamma))=\eta(\alpha,P_{m}(\gamma,\beta))$ for all $\alpha\in H_{3}$ and $\beta,\gamma\in\Delta$ where $m=m_{\alpha,\beta}$ denotes the group-theoretic order of $\sigma_{\beta}\sigma_{\gamma}$ in the Weyl group, then $\eta$ is called braid-invariant. If for all roots $\alpha,\beta\in H_{3}$ such that $\alpha$ is adjacent to $\beta$ and $-\beta$ and for all $\Delta$-expressions $\bar{\beta}$ of $\beta$, we have $\eta(\alpha,\bar{\beta})=1_{A}$, then $\eta$ is called adjacency-trivial. If for all $\alpha\in H_{3}$ and all $b\in A$, there exists a word $\bar{\delta}$ over $\Delta$ such that $\sigma_{\bar{\delta}}$ stabilises $\alpha$ and $\eta(\alpha,\bar{\delta})=b$, then $\eta$ is called complete.
Remark 6.7.
Since the Weyl group acts transitively on $H_{3}$, it suffices to verify completeness for one root $\alpha$.
Definition 6.8.
A (faithful) parameter system is a pair $(A,M)$ consisting of abelian groups $A,M$ such that $A$ acts (faithfully) on the underlying set of $M$.
Definition 6.9.
Let $(A,M)$ be a parameter system and let $\eta$ be a parity map with values in $A$. A parametrisation of $(G,(U_{\alpha})_{\alpha\in H_{3}})$ by $(A,M)$ with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$ is a family $(\theta_{\alpha}:M\rightarrow U_{\alpha})_{\alpha\in\Phi}$ of group isomorphisms such that for all $\alpha\in\Phi$, $\delta\in\Delta$ and $x\in M$, we have
$\theta_{\alpha}(x)^{w_{\delta}}=\theta_{\alpha^{\sigma(\delta)}}(\eta_{\alpha,\delta}.x).$
The family $(\theta_{\alpha})_{\alpha\in\Phi}$ will also be called a parametrisation of $G$ and the maps in this family are called root isomorphisms. We will also say that $G$ is parametrised by $(A,M)$ with respect to $\eta$, $(w_{\delta})_{\delta\in\Delta}$ and with root isomorphisms $(\theta_{\alpha})_{\alpha\in\Phi}$.
Example 6.10.
Let $A\coloneqq\{\pm 1\}^{2}$, let $\mathcal{R}$ be a commutative ring and put $M\coloneqq(\mathcal{R}^{2},+)$. Let $A$ act on $M$ via $(\varepsilon,\varepsilon^{\prime}).(x,y)\coloneqq(\varepsilon x,\varepsilon^{\prime}y)$ for all $x,y\in\mathcal{R}$ and $\varepsilon,\varepsilon^{\prime}\in\{\pm 1\}$. Then $(A,M)$ is a parameter system, called the standard parameter system for $\mathcal{R}$. It is faithful if and only if $2_{\mathcal{R}}\neq 0_{\mathcal{R}}$. We will show that every $H_{3}$-graded group is parametrised by a parameter system of this form.
The final missing ingredient in the parametrisation theorem is the notion of twisting groups.
Definition 6.11.
Let $A$ be an abelian group equipped with an action
$A\times U_{\alpha}\rightarrow U_{\alpha},(b,x_{\alpha})\mapsto b.x_{\alpha}$
on the underlying set of $U_{\alpha}$ for each $\alpha\in H_{3}$. Then $A$ is called a twisting group for $(G,(w_{\delta})_{\delta\in\Delta})$ if for all $\alpha\in H_{3}$, $\delta\in\Delta\cup(-\Delta)$ and $b\in A$, we have $(b.x_{\alpha})^{w_{\delta}}=b.(x_{\alpha}^{w_{\delta}})$ for all $x_{\alpha}\in U_{\alpha}$. We will refer to the actions of $A$ on the root groups as twisting actions (on the root groups).
Example 6.12.
Put $A\coloneqq\{\pm 1\}^{2}$. For all $\alpha\in H_{3}$ any $\alpha\in H_{3}$, we declare that $A$ acts on $U_{\alpha}$ via $(-1,-1).x_{\alpha}\coloneqq x_{\alpha}^{-1}$ and $(-1,1).x_{\alpha}\coloneqq x_{\alpha}^{*_{\alpha}}$ for all $x_{\alpha}\in U_{\alpha}$. Then by 5.19, $A$ is a twisting group for $(G,(w_{\delta})_{\delta\in\Delta})$. We will use this twisting group to construct a parametrisation of $G$. If $G=\operatorname{Ch}(D_{6},\mathcal{R})$ for some commutative ring $\mathcal{R}$ as in 4.10, then by 5.21, $(\varepsilon,\varepsilon^{\prime}).\tilde{\theta}_{\alpha}^{H}(x,y)=\tilde{\theta}_{\alpha}^{H}\big{(}(\varepsilon,\varepsilon^{\prime}).(x,y)\big{)}$ for all $x,y\in\mathcal{R}$ and all $(\varepsilon,\varepsilon^{\prime})\in A$.
Definition 6.13.
Let $\eta$ be a $\Delta$-parity map with values in a twisting group $A$ for $(G,(w_{\delta})_{\delta\in\Delta})$. We say that $G$ is square-compatible with respect to $\eta$ (and $(w_{\delta})_{\delta\in\Delta}$) if for all $\alpha\in H_{3}$ and all $\delta\in\Delta$, we have $x_{\alpha}^{w}=\eta(\alpha,(\delta,\delta)).x_{\alpha}$ for all $x_{\alpha}\in U_{\alpha}$ where $w\coloneqq w_{\delta}^{2}$.
The following assertion is the main step in the proof of the parametrisation theorem.
Proposition 6.14 ([Wie23, 4.5.6]).
Assume that $\eta$ is a complete braid-invariant $\Delta$-parity map with values in a twisting group $A$ for $(G,(w_{\delta})_{\delta\in\Delta})$ such that $G$ is square-compatible with respect to $\eta$. Let $\alpha\in H_{3}$ and let $\bar{\beta}$, $\bar{\gamma}$ be two words over Δ($\Delta\cup(-\Delta)$ such that $\alpha^{\sigma(\bar{\beta})}=\alpha^{\sigma(\bar{\gamma})}$. Put $b\coloneqq\eta_{\alpha,\bar{\beta}}\eta_{\alpha,\bar{\gamma}}^{-1}$. Then $x_{\alpha}^{w_{\bar{\beta}}}=(b.x_{\alpha})^{w_{\bar{\gamma}}}$ for all $x_{\alpha}\in U_{\alpha}$.
Sketch of the proof.
The general assertion can be deduced from the special case that $\sigma_{\bar{\beta}}$ and $\sigma_{\bar{\gamma}}$ stabilise $\alpha$ and that $\bar{\gamma}$ is the empty word. Then there exist $\rho_{1},\ldots,\rho_{n}\in H_{3}\cap\alpha^{\perp}$ with $\sigma_{\bar{\beta}}=\sigma_{\rho_{1}}\cdots\sigma_{\rho_{n}}$. Using induction, we can reduce to the case that $n=1$, or in other words, that $\sigma_{\bar{\beta}}=\sigma_{\rho}$ for a root $\rho\in\alpha^{\perp}$. Using square-compatibility, the braid-invariance of $\eta$, the braid relations for $(w_{\delta})_{\delta\in\Delta}$ and the solution of the word problem in Coxeter groups, we can replace $\bar{\beta}$ by any word $\bar{\beta}^{\prime}$ with $\sigma_{\bar{\beta}}=\sigma_{\bar{\beta}^{\prime}}$. In particular, we can replace $\bar{\beta}$ by a $\Delta$-expression $\bar{\beta}^{\prime}$ of $\rho$. Then $w_{\bar{\beta}^{\prime}}$ centralises $U_{\alpha}$ by 6.2 and $\eta(\alpha,\bar{\beta}^{\prime})=1_{A}$ because $\eta$ is adjacency-trivial. This finishes the proof. ∎
The following result is precisely the content of the parametrisation theorem [Wie23, 4.5.16] for the special case $\Phi=H_{3}$.
Theorem 6.15 ([Wie23, 4.5.16]).
Under the assumptions of 6.14, there exist an abelian group $(\mathcal{S},+)$, an action $\omega\colon A\times\mathcal{S}\rightarrow\mathcal{S},(b,r)\mapsto b.r$ of $A$ on the underlying set of $\mathcal{S}$ and a parametrisation $(\theta_{\alpha}\colon\mathcal{S}\rightarrow U_{\alpha})_{\alpha\in H_{3}}$ of $G$ by $(A,\mathcal{S})$ with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$ such that the twisting actions of $A$ on the root groups are induced by $\omega$. That is, we have $b.\theta_{\alpha}(s)=\theta_{\alpha}(b.s)$ for all $\alpha\in H_{3}$, $s\in\mathcal{S}$ and $b\in A$.
Sketch of the proof.
Fix a root $\alpha_{0}\in H_{3}$. Choose a group $\mathcal{S}$ which is isomorphic to $U_{\alpha_{0}}$ and an isomorphism $\theta_{\alpha_{0}}\colon\mathcal{S}\rightarrow U_{\alpha_{0}}$. We define an action of $A$ on $\mathcal{S}$ by $b.s\coloneqq\theta_{\alpha_{0}}^{-1}(b.\theta_{\alpha_{0}}(s))$ for all $b\in A$ and $s\in\mathcal{S}$. For any root $\beta$, we define $\theta_{\beta}\colon\mathcal{S}\rightarrow U_{\beta},s\mapsto\theta_{\alpha_{0}}(s)^{w_{\bar{\gamma}}}$ where $\bar{\gamma}$ is any word over Δ($\Delta\cup(-\Delta)$ with $\alpha_{0}^{\sigma(\bar{\gamma})}=\beta$ and $\eta(\alpha,\bar{\gamma})=1_{A}$. Such a word $\bar{\gamma}$ exists by the completeness of $\eta$, and $\theta_{\beta}$ does not depend on the choice of $\bar{\gamma}$ by 6.14. One can verify that this construction has the desired properties. ∎
In order to apply the parametrisation theorem 6.15 to $(G,(U_{\alpha})_{\alpha\in H_{3}})$, we need to find an appropriate twisting group $A$ and a $\Delta$-parity map $\eta$. The former was constructed in 6.10. The latter can be “read off” from the $H_{3}$-graded group in Section 4.
Proposition 6.16.
Consider the $H_{3}$-graded group $\mathcal{G}\coloneqq(\operatorname{Ch}(D_{6},\mathbb{C}),(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}})$ constructed in 4.10 for $\mathcal{R}\coloneqq\mathbb{C}$ and the faithful parameter system $(A,\mathbb{C}^{2})$ from 6.10. Then there exists a unique $\Delta$-parity map $\eta\colon H_{3}\times\Delta\rightarrow A$ such that $\mathcal{G}$ is parametrised by $(A,\mathbb{C}^{2})$ with respect to $\eta$, the Weyl elements $(w^{H}_{\delta}(1,1))_{\delta\in\Delta}$ from 4.14 and the root isomorphisms $(\tilde{\theta}_{\alpha}^{H})_{\alpha\in H_{3}}$ from 4.10. Further, $\eta$ is complete, braid-invariant and satisfies $\eta(\alpha,\delta)=\eta(-\alpha,\delta)$ for all $\alpha\in H_{3}$ and $\delta\in\Delta$. The values of $\eta$ are given in Figure 5.
Proof.
Put $w_{\delta}\coloneqq w^{H}_{\delta}(1,1)$ for all $\delta\in\Delta$. For all $\alpha\in H_{3}$ and $\delta\in\Delta$, a straightforward computation (see 2.11) yields a tuple $(\varepsilon,\varepsilon^{\prime})\in\{\pm 1\}^{2}$ such that, for all $x,y\in\mathbb{C}$,
$\tilde{\theta}_{\alpha}^{H}(x,y)^{w_{\delta}}=\tilde{\theta}_{\alpha^{\sigma(\delta)}}^{H}(\varepsilon x,\varepsilon^{\prime}y)\quad\big{(}=\tilde{\theta}_{\alpha^{\sigma(\delta)}}^{H}((\varepsilon,\varepsilon^{\prime}).(x,y))\big{)}.$
We now put $\eta_{\alpha,\delta}\coloneqq(\varepsilon,\varepsilon^{\prime})$. Since $1_{\mathbb{C}}\neq-1_{\mathbb{C}}$, the action of $A$ on $\mathbb{C}^{2}$ is faithful, so the signs $\varepsilon,\varepsilon^{\prime}$ are uniquely determined by $\alpha$ and $\delta$. The results of these computations are given in Figure 5, and we observe that $\eta(\alpha,\delta)=\eta(-\alpha,\delta)$ for all $\alpha\in H_{3}$ and $\delta\in\Delta$.
It remains to verify that the $\Delta$-parity map $\eta:H_{3}\times\Delta\rightarrow A$ has the desired properties. We can read off from Figure 5 that $\eta(\rho_{2},(\rho_{1},\rho_{1}))=(-1,-1)$ and $\eta(\rho_{2},(\rho_{3},\rho_{3}))=(-1,1)$, so $\eta$ is complete by 6.7. Further, braid-invariance holds by 3.15 and the faithfulness of the action of $A$ on $\mathbb{C}^{2}$. (Alternatively, braid-invariance can be derived from Figure 5 with a lot of patience.) ∎
$\alpha$ $\eta(\alpha,\rho_{1})$ $\eta(\alpha,\rho_{2})$ $\eta(\alpha,\rho_{3})$
$[0,0,1]$ $(1,1)$ $(-1,-1)$ $(-1,-1)$
$[0,\tau,1]$ $(1,1)$ $(1,-1)$ $(1,1)$
$[0,\tau,\tau]$ $(-1,-1)$ $(1,1)$ $(1,-1)$
$[\tau,\tau,1]$ $(-1,1)$ $(1,1)$ $(-1,-1)$
$[0,1,\tau]$ $(-1,1)$ $(1,-1)$ $(-1,-1)$
$[\tau,\tau,\tau]$ $(1,-1)$ $(-1,-1)$ $(-1,1)$
$[0,1,0]$ $(-1,-1)$ $(-1,-1)$ $(1,-1)$
$[1,1,\tau]$ $(1,-1)$ $(1,1)$ $(-1,1)$
$[\tau,\tau^{2},\tau]$ $(-1,1)$ $(1,1)$ $(-1,-1)$
$[1,1,0]$ $(1,1)$ $(-1,-1)$ $(1,1)$
$[1,\tau^{2},\tau]$ $(-1,-1)$ $(-1,1)$ $(1,-1)$
$[\tau,\tau^{2},\tau^{2}]$ $(1,1)$ $(-1,-1)$ $(1,1)$
$[1,0,0]$ $(-1,-1)$ $(1,1)$ $(1,1)$
$[1,\tau^{2},\tau^{2}]$ $(1,-1)$ $(1,1)$ $(-1,1)$
$[\tau,2\tau,\tau^{2}]$ $(1,1)$ $(-1,1)$ $(1,1)$
Figure 5: Values of the parity map $\eta$ in 6.16, using the notation from 4.6. The values $\eta_{\alpha,\delta}$ for negative roots $\alpha$ are given by the formula $\eta_{-\alpha,\delta}=\eta_{\alpha,\delta}$.
Remark 6.17.
Note that $\eta$ can be defined without reference to $\mathcal{G}$ by declaring that it is given by the values in Figure 5. Further, $\mathcal{G}_{\mathcal{R}}\coloneqq(\operatorname{Ch}(D_{6},\mathcal{R}),(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}})$ is parametrised by $(A,\mathcal{R}^{2})$ with respect to $\eta$ for any commutative ring $\mathcal{R}$. However, if $2_{\mathcal{R}}=0_{\mathcal{R}}$, then $(A,\mathcal{R}^{2})$ is not faithful, so there exist multiple parity maps $\eta^{\prime}$ such that $\mathcal{G}_{\mathcal{R}}$ is parametrised by $(A,\mathcal{R}^{2})$ with respect to $\eta^{\prime}$.
Proposition 6.18.
Let $\eta$ be the parity map from 6.16 and let $A\coloneqq\{\pm 1\}^{2}$ act on the root groups $(U_{\alpha})_{\alpha\in H_{3}}$ of $G$ as in 6.12. Then $G$ is square-compatible with respect to $\eta$.
Proof.
In addition to its action on the root groups, we let $A$ act on $\mathbb{C}^{2}$ as in 6.10. Since the group $\mathcal{G}$ from 6.16 is parametrised with respect to $\eta$, we have
$\tilde{\theta}_{\alpha}^{H}(x,y)^{w_{\delta}^{2}}=\tilde{\theta}_{\alpha}^{H}\big{(}\eta(\alpha,(\delta,\delta)).(x,y)\big{)}$
for all $x,y\in\mathbb{C}$. At the same time, the action of $w_{\delta}^{2}$ on $\tilde{\theta}_{\alpha}^{H}(x,y)$ is completely described by 5.20 and 5.21. Since the parameter system $(A,\mathbb{C}^{2})$ is faithful, we conclude: $\eta(\alpha,(\delta,\delta))=(1,1)$ if $\Psi_{\alpha,\delta}$ is of type $A_{1}$ or $A_{1}\times A_{1}$; $\eta(\alpha,(\delta,\delta))=(-1,-1)$ if $\Psi_{\alpha,\delta}$ is of type $A_{2}$; $\eta(\alpha,(\delta,\delta))=(-1,1)$ if $\alpha$ is in involution position with respect to $\delta$ and $\eta(\alpha,(\delta,\delta))=(1,-1)$ if $\alpha$ is in inverted involution position with respect to $\alpha$. Since 5.20 applies to $(G,(U_{\alpha})_{\alpha\in H_{3}})$ as well, the previous statement and the definition of the action of $A$ on $(U_{\alpha})_{\alpha\in H_{3}}$ imply that $G$ is square-compatible with respect to $\eta$. ∎
Conclusion 6.19.
Let $(G,(U_{\alpha})_{\alpha\in H_{3}})$ be an arbitrary $H_{3}$-graded group, let $(w_{\delta})_{\delta\in\Delta}$ be a family of Weyl elements in $G$, let $A\coloneqq\{\pm 1\}^{2}$ denote the twisting group from 6.12 and let $\eta\colon H_{3}\rightarrow A$ denote the parity map from 6.16. Then there exist an abelian group $(\mathcal{S},+)$ on which $A$ acts and a parametrisation $(\theta_{\alpha})_{\alpha\in H_{3}}$ of $G$ by $(A,\mathcal{S})$ with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$. Putting $s^{*}\coloneqq(-1,1).s$ for all $s\in\mathcal{S}$, we can summarise the property of being a parametrisation as follows:
$\theta_{\alpha}(s)^{w_{\delta}}=\begin{cases}\theta_{\alpha^{\sigma(\delta)}}(s)&\text{if }\eta_{\alpha,\delta}=(1,1),\\ \theta_{\alpha^{\sigma(\delta)}}(-s)&\text{if }\eta_{\alpha,\delta}=(-1,-1),\\ \theta_{\alpha^{\sigma(\delta)}}(s^{*})&\text{if }\eta_{\alpha,\delta}=(-1,1),\\ \theta_{\alpha^{\sigma(\delta)}}(-s^{*})&\text{if }\eta_{\alpha,\delta}=(1,-1)\end{cases}$ (6.20)
for all $\alpha\in H_{3}$, $\delta\in\Delta$ and $s\in\mathcal{S}$. Note that $(s^{*})^{*}=s$ for all $s\in\mathcal{S}$ because $(-1,1)^{2}=(1,1)$. Further, we define a multiplication on $\mathcal{S}$ by
$a\cdot b\coloneqq\theta_{\rho_{1}+\rho_{2}}^{-1}([\theta_{\rho_{1}}(a),\theta_{\rho_{2}}(b)]).$
It satisfies the distributive laws by 3.8 because $(\rho_{1},\rho_{2})$ is an $A_{2}$-pair. Finally, we define $1_{\mathcal{S}}$ to be the element of $\mathcal{S}$ which satisfies $w_{\rho_{2}}\in U_{-\rho_{2}}\theta_{\rho_{2}}(1_{\mathcal{S}})U_{-\rho_{2}}$, which is uniquely determined by 5.3 (iii). By 5.3 (iv), we have
$\theta_{\rho_{1}+\rho_{2}}(s\cdot 1_{\mathcal{S}})$ $=[\theta_{\rho_{1}}(s),\theta_{\rho_{2}}(1_{\mathcal{S}})]=\theta_{\rho_{1}}(s)^{w_{\rho_{2}}}=\theta_{\rho_{1}+\rho_{2}}(\eta_{\rho_{1},\rho_{2}}.s)=\theta_{\rho_{1}+\rho_{2}}(s)$
for all $s\in\mathcal{S}$. A similar computation (see [Wie23, 5.7.12]) shows that $1_{\mathcal{S}}\cdot r=r$ for all $r\in\mathcal{S}$ as well, so that $(\mathcal{S},+,\cdot)$ is a nonassociative ring with unit $1_{\mathcal{S}}$.
Goal 6.21.
We want to show that the multiplication on $\mathcal{S}$ from 6.19 is associative and commutative, that $\mathcal{S}$ decomposes as $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ for a pair of isomorphic ideals $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ of $\mathcal{S}$ and that $(r,s)^{*}=(-r,s)$ for all $(r,s)\in\mathcal{S}_{1}\oplus\mathcal{S}_{2}$. Further, we will show that the commutator relations in $G$ are precisely those given in Figure 4.
Remark 6.22.
All results in this section hold for $H_{4}$-graded groups as well, except that the parity map $\eta$ has to be defined with respect to the $H_{4}$-graded group from 4.23.
7 The blueprint technique
Notation 7.1.
In this section, Notation 6.1 continues to hold. Further, we define the nonassociative ring $\mathcal{S}$ and the root isomorphisms $(\theta_{\alpha})_{\alpha\in H_{3}}$ as in 6.19.
The blueprint technique was introduced in [Wie23] as a computational tool to prove identities in root graded groups for arbitrary root systems (of rank at least 3). It is inspired by the work of Ronan-Tits on the construction of buildings in [RT87]. In this section, we give a brief summary of the general technique by the example of $H_{3}$-gradings. A more detailed exposition can be found in [Wie23, 6.1]. In Section 8, we will perform the actual blueprint computations for $H_{3}$-graded groups in order to achieve the goal in 6.21.
At first, we have to introduce some notation and terminology.
Definition 7.2.
Let $f$, $g$ be two words over $\Delta$. We say that $f$ is elementary braid-homotopic to $g$ if there exist $\alpha,\beta\in\Delta$ and (possibly empty) words $f_{1},f_{2}$ over $\Delta$ such that $f=f_{1}P_{m}(\alpha,\beta)f_{2}$ and $g=f_{1}P_{m}(\beta,\alpha)f_{2}$ where $m$ denotes the group-theoretic order of $\sigma_{\alpha}\sigma_{\beta}$ in the Weyl group and $P_{m}$ is defined as in 6.5. We say that $f$ and $g$ are braid-homotopic if they lie in the equivalence relation generated by the relation “elementary braid-homotopic”. Further, we say that $f$ is reduced if it does not contain a subword of the form $(\alpha,\alpha)$ for some $\alpha\in\Delta$.
Notation 7.3.
For any word $\bar{\delta}=(\delta_{1},\ldots,\delta_{k})$ over $\Delta$, we put $U_{\bar{\delta}}\coloneqq U_{\delta_{1}}\times\cdots\times U_{\delta_{k}}$.
Definition 7.4.
For any word $\bar{\delta}$ over $\Delta$, the blueprint invariant for $\bar{\delta}$ (with respect to $(w_{\delta})_{\delta\in\Delta}$) is the map
$\gamma_{\bar{\delta}}\colon U_{\bar{\delta}}\rightarrow G,(g_{1},\ldots,g_{m})\mapsto\prod_{i=1}^{m}{w_{\delta_{i}}g_{i}}.$
For any two words $\bar{\alpha},\bar{\beta}$ over $\Delta$, a blueprint rewriting rule of type $(\bar{\alpha},\bar{\beta})$ (with respect to $(w_{\delta})_{\delta\in\Delta}$) is a map $\varphi\colon U_{\bar{\alpha}}\rightarrow U_{\bar{\beta}}$ which satisfies $\gamma_{\bar{\beta}}\circ\varphi=\gamma_{\bar{\alpha}}$.
Lemma 7.5 ([Wie23, 6.2.11]).
Concatenations of blueprint rewriting rules and inverses of bijective blueprint rewriting rules are blueprint rewriting rules.
Lemma 7.6 ([Wie23, 6.2.14]).
Let $\bar{\alpha},\bar{\alpha}^{\prime},\bar{\alpha}^{\prime\prime},\bar{\beta}$ be words over $\Delta$ and let $\varphi$ be a blueprint rewriting rule of type $(\bar{\alpha},\bar{\beta})$. Then $\mathord{\mathrm{id}}\times\varphi\times\mathord{\mathrm{id}}\colon U_{\bar{\alpha}^{\prime}}\times U_{\bar{\alpha}}\times U_{\bar{\alpha}^{\prime\prime}}\rightarrow U_{\bar{\alpha}^{\prime}}\times U_{\bar{\beta}}\times U_{\bar{\alpha}^{\prime\prime}}$ is a blueprint rewriting rule.
The heart of the blueprint technique is the following result. It follows from the fact that the blueprint invariant is injective for any reduced word, which is a consequence of 3.10.
Theorem 7.7 ([Wie23, 6.2.8]).
Let $\bar{\alpha},\bar{\beta}$ be two reduced words over $\Delta$ such that $\sigma_{\bar{\alpha}}=\sigma_{\bar{\beta}}$, let $x\in U_{\bar{\alpha}}$ and let $\varphi_{1},\varphi_{2}\colon U_{\bar{\alpha}}\rightarrow U_{\bar{\beta}}$ be two blueprint rewriting rules. Then $\varphi_{1}(x)=\varphi_{2}(x)$.
Algorithm 7.8.
Put $\theta_{i}\coloneqq\theta_{\rho_{i}}$ and $U_{i}\coloneqq U_{\rho_{i}}$ for all $i\in\{1,2,3\}$ and $U_{i_{1}\cdots i_{k}}\coloneqq U_{i_{1}}\times\cdots\times U_{i_{k}}$ for all $i_{1},\ldots,i_{k}\in\{1,2,3\}$. Further, we denote by $w_{0}$ the longest element in the Weyl group of $H_{3}$ and choose an arbitrary reduced expression of $w_{0}$, say
$f_{1}=(\rho_{3},\rho_{2},\rho_{3},\rho_{2},\rho_{3},\rho_{1},\rho_{2},\rho_{3},\rho_{2},\rho_{3},\rho_{1},\rho_{2},\rho_{3},\rho_{2},\rho_{1}).$
That is, $\sigma(f_{1})=w_{0}$. To simplify notation, we will write $f_{1}=323231232312321$. Further, we define $f_{2},\ldots,f_{63}$ to be the homotopy cycle for $w_{0}$ given by Figure 6. Observe that $f_{i}$ is elementary braid-homotopic to $f_{i+1}$ for all $i\in\{1,\ldots,62\}$ and that $f_{1}=f_{63}$.
Recall that the Weyl group of $H_{3}$ satisfies the braid relations
$\sigma(\rho_{1}\rho_{2}\rho_{1})=\sigma(\rho_{2}\rho_{1}\rho_{2}),\;\sigma(\rho_{2}\rho_{3}\rho_{2}\rho_{3}\rho_{2})=\sigma(\rho_{3}\rho_{2}\rho_{3}\rho_{2}\rho_{3}),\;\sigma(\rho_{1}\rho_{3})=\sigma(\rho_{3}\rho_{1}).$
We assume that for all distinct $i,j\in\{1,2,3\}$, we have a blueprint rewriting rule $\varphi_{ij}$ of type $(P_{m_{ij}}(\rho_{i},\rho_{j}),P_{m_{ji}}(\rho_{j},\rho_{i}))$. For example, we have blueprint rewriting rules $\varphi_{12}\colon U_{121}\rightarrow U_{212}$ and $\varphi_{21}\colon U_{212}\rightarrow U_{121}$. Using these “elementary” blueprint rewriting rules and 7.6, we can define blueprint rewriting rules $\varphi_{i}$ of type $(f_{i},f_{i+1})$ for all $i\in\{1,\ldots,62\}$. Similarly, we can define blueprint rewriting rules $\varphi_{i}^{\prime}$ of type $(f_{i+1},f_{i})$ for all $i\in\{2,\ldots,63\}$.
Now let $y_{1},y_{2},\ldots,y_{15}\in\mathcal{R}$ be arbitrary ring elements and consider the tuple
$x_{1}\coloneqq\big{(}\theta_{3}(y_{1}),\theta_{2}(y_{2}),\theta_{3}(y_{3}),\ldots,\theta_{1}(y_{15})\big{)}\in U_{f_{1}}=U_{f_{63}}.$
For all $i\in\{2,\ldots,63\}$, we can successively compute tuples $x_{i}\coloneqq\varphi_{i-1}(x_{i-1})$. Since the identity map on $U_{f_{1}}$ and $\varphi_{62}\circ\cdots\circ\varphi_{1}$ are blueprint rewriting rules, it follows from 7.7 that $x_{1}=x_{63}$. Writing
$x_{63}=\big{(}\theta_{3}(y_{1}^{(63)}),\theta_{2}(y_{2}^{(63)}),\theta_{3}(y_{3}^{(63)}),\ldots,\theta_{1}(y_{15}^{(63)})\big{)},$
we obtain 15 identities of the form $y_{i}=y_{i}^{(63)}$ which hold for arbitrary ring elements $y_{1},\ldots,y_{15}$.
In practice, we will only compute $x_{2},\ldots,x_{32}$ and, in a separate computation, tuples $x_{63}^{\prime},\ldots,x_{32}^{\prime}$ defined by $x_{63}^{\prime}\coloneqq x_{1}$ and $x_{i}^{\prime}\coloneqq\varphi_{i+1}^{\prime}(x_{i+1}^{\prime})$ for all $i\in\{32,\ldots,62\}$. By the same argument as above, we then have $x_{32}=x_{32}^{\prime}$, which yields 15 identities in $\mathcal{R}$ (which are shorter than the 15 identities in the previous paragraph). We will refer to these identities as the blueprint identities.
  1. (1)
    323231232312321
  2. (2)
    232321232312321
  3. (3)
    232312132312321
  4. (4)
    232312312312321
  5. (5)
    232312312132321
  6. (6)
    232312321232321
  7. (7)
    232312321323231
  8. (8)
    232132321323231
  9. (9)
    232132323123231
  10. (10)
    232123232123231
  11. (11)
    231213232123231
  12. (12)
    231213231213231
  13. (13)
    231231231213231
  14. (14)
    231231231231231
  15. (15)
    231231213231231
  16. (16)
    213231213231231
  17. (17)
    213232123231231
  18. (18)
    213232123231213
  19. (19)
    213232123232123
  20. (20)
    213232132323123
  21. (21)
    213232312323123
  22. (22)
    212323212323123
  23. (23)
    121323212323123
  24. (24)
    121323121323123
  25. (25)
    123123121323123
  26. (26)
    123121321323123
  27. (27)
    123212321323123
  28. (28)
    123212323123123
  29. (29)
    123212323121323
  30. (30)
    123212323212323
  31. (31)
    123213232312323
  32. (32)
    123231232312323
  33. (33)
    123231232132323
  34. (34)
    123231232123232
  35. (35)
    123231231213232
  36. (36)
    123231213213232
  37. (37)
    123232123213232
  38. (38)
    132323123213232
  39. (39)
    312323123213232
  40. (40)
    312321323213232
  41. (41)
    312321323231232
  42. (42)
    312321232321232
  43. (43)
    312312132321232
  44. (44)
    312132132321232
  45. (45)
    321232132321232
  46. (46)
    321232312321232
  47. (47)
    321232312312132
  48. (48)
    321232312132132
  49. (49)
    321232321232132
  50. (50)
    321323231232132
  51. (51)
    323123231232132
  52. (52)
    323123213232132
  53. (53)
    323123213232312
  54. (54)
    323123212323212
  55. (55)
    323123121323212
  56. (56)
    323121321323212
  57. (57)
    323212321323212
  58. (58)
    323212323123212
  59. (59)
    323212323123121
  60. (60)
    323212323121321
  61. (61)
    323212323212321
  62. (62)
    323213232312321
  63. (63)
    323231232312321
Figure 6: A homotopy cycle for the longest word in $H_{3}$. The underlines indicate the elementary braid homotopy moves which are applied at each step.
The computation in 7.8 will be refered to as the blueprint computation. In order to actually perform this computation, it remains to compute the (elementary) blueprint rewriting rules.
Notation 7.9 (compare 4.20).
For all non-proportional roots $\alpha,\gamma\in H_{3}$ and all $\beta\in\mathord{]\alpha,\gamma[}$, we define a map $\psi_{\alpha,\gamma}^{\beta}\colon\mathcal{R}^{2}\rightarrow\mathcal{R}$ by the property that
$[\theta_{\alpha}(x),\theta_{\gamma}(y)]_{\beta}=\theta_{\beta}\big{(}\psi_{\alpha,\gamma}^{\beta}(x,y))$
for all $x_{1},x_{2},y_{1},y_{2}\in\mathcal{R}$. If $\mathord{]\alpha,\gamma[}=\{\beta\}$, we will write $\psi_{\alpha,\gamma}$ in place of $\psi_{\alpha,\gamma}^{\beta}$.
Notation 7.10.
From now on, we put $\alpha\coloneqq\rho_{2}$, $\varepsilon\coloneqq\rho_{3}$ and we denote the $H_{2}$-quintuple corresponding to the $H_{2}$-pair $(\alpha,\varepsilon)$ by $(\alpha,\beta,\gamma,\delta,\varepsilon)$.
Definition 7.11.
We define the following maps:
$\varphi_{12}:~{}$ $U_{\rho_{1}}\times U_{\rho_{2}}\times U_{\rho_{1}}\rightarrow U_{\rho_{2}}\times U_{\rho_{1}}\times U_{\rho_{2}},$
$\left(\theta_{\rho_{1}}(a),\theta_{\rho_{2}}(b),\theta_{\rho_{1}}(c)\right)\mapsto\left(\theta_{\rho_{2}}(c),\theta_{\rho_{1}}(-b-ca),\theta_{\rho_{2}}(a)\right),$
$\varphi_{21}:~{}$ $U_{\rho_{2}}\times U_{\rho_{1}}\times U_{\rho_{2}}\rightarrow U_{\rho_{1}}\times U_{\rho_{2}}\times U_{\rho_{1}},$
$\left(\theta_{\rho_{2}}(a),\theta_{\rho_{1}}(b),\theta_{\rho_{2}}(c)\right)\mapsto\left(\theta_{\rho_{1}}(c),\theta_{\rho_{2}}(-b-ac),\theta_{\rho_{1}}(a)\right),$
$\varphi_{13}:~{}$ $U_{\rho_{1}}\times U_{\rho_{3}}\rightarrow U_{\rho_{3}}\times U_{\rho_{1}},$
$\left(\theta_{\rho_{1}}(a),\theta_{\rho_{3}}(b)\right)\mapsto\left(\theta_{\rho_{3}}(b),\theta_{\rho_{1}}(a)\right),$
$\varphi_{31}:~{}$ $U_{\rho_{3}}\times U_{\rho_{1}}\rightarrow U_{\rho_{1}}\times U_{\rho_{3}},$
$\left(\theta_{\rho_{3}}(a),\theta_{\rho_{1}}(b)\right)\mapsto\left(\theta_{\rho_{1}}(b),\theta_{\rho_{3}}(a)\right).$
Furthermore, we define
$\varphi_{23}:~{}$ $U_{\rho_{2}}\times U_{\rho_{3}}\times U_{\rho_{2}}\times U_{\rho_{3}}\times U_{\rho_{2}}\rightarrow U_{\rho_{3}}\times U_{\rho_{2}}\times U_{\rho_{3}}\times U_{\rho_{2}}\times U_{\rho_{3}},$
$\left(\theta_{\rho_{2}}(a),\theta_{\rho_{3}}(b),\theta_{\rho_{2}}(c),\theta_{\rho_{3}}(d),\theta_{\rho_{2}}(e)\right)$
$\quad\mapsto\left(\theta_{\rho_{3}}(e),\theta_{\rho_{2}}(B),\theta_{\rho_{3}}(C),\theta_{\rho_{2}}(D),\theta_{\rho_{3}}(a)\right)$
where
$B$ $\coloneqq-\Big{(}\psi_{\delta,\alpha}^{\beta}(-b,e)+\psi_{\varepsilon,\alpha}^{\beta}(a,e)+\psi_{\gamma,\alpha}(c,e)-d\Big{)}^{*},$
$C$ $\coloneqq-\Big{(}\psi_{\delta,\alpha}^{\gamma}(-b,e)+\psi_{\delta,\beta}\big{(}-b,\psi_{\delta,\alpha}^{\beta}(-b,e)\big{)}$
$\qquad\quad\mathord{}+\psi_{\varepsilon,\alpha}^{\gamma}(a,e)+\psi_{\varepsilon,\beta}^{\gamma}\big{(}a,\psi_{\varepsilon,\alpha}^{\beta}(a,e)+\psi_{\gamma,\alpha}(c,e)-d\big{)}$
$\qquad\quad\mathord{}+\psi_{\delta,\beta}\big{(}-b+\psi_{\varepsilon,\alpha}^{\delta}(a,e),\psi_{\varepsilon,\alpha}^{\beta}(a,e)+\psi_{\gamma,\alpha}(c,e)-d\big{)}+c\Big{)},$
$D$ $\coloneqq-\bigg{(}-b+\psi_{\varepsilon,\alpha}^{\delta}(a,e)+\psi_{\varepsilon,\beta}^{\delta}\big{(}a,\psi_{\varepsilon,\alpha}^{\beta}(a,e)+\psi_{\gamma,\alpha}(c,e)-d\big{)}$
$\qquad\quad\mathord{}+\psi_{\varepsilon,\gamma}\Big{(}a,\psi_{\varepsilon,\beta}^{\gamma}\big{(}a,\psi_{\varepsilon,\alpha}^{\beta}(a,e)+\psi_{\gamma,\alpha}(c,e)-d\big{)}+\psi_{\varepsilon,\alpha}^{\gamma}(a,e)+c\big{)}\bigg{)}^{*}.$
Similarly, we define
$\varphi_{32}:~{}$ $U_{\rho_{3}}\times U_{\rho_{2}}\times U_{\rho_{3}}\times U_{\rho_{2}}\times U_{\rho_{3}}\rightarrow U_{\rho_{2}}\times U_{\rho_{3}}\times U_{\rho_{2}}\times U_{\rho_{3}}\times U_{\rho_{2}},$
$\left(\theta_{\rho_{3}}(a),\theta_{\rho_{2}}(b),\theta_{\rho_{3}}(c),\theta_{\rho_{2}}(d),\theta_{\rho_{3}}(e)\right)$
$\quad\mapsto\left(\theta_{\rho_{2}}(e),\theta_{\rho_{3}}(B),\theta_{\rho_{2}}(C),\theta_{\rho_{3}}(D),\theta_{\rho_{2}}(a)\right)$
where
$B$ $\coloneqq-\Big{(}\psi_{\beta,\varepsilon}^{\delta}(-b^{*},e)+\psi_{\alpha,\varepsilon}^{\delta}(a,e)+\psi_{\gamma,\varepsilon}(-c,e)-d^{*}\Big{)},$
$C$ $\coloneqq\psi_{\beta,\varepsilon}^{\gamma}(-b^{*},e)+\psi_{\beta,\delta}\big{(}-b^{*},\psi_{\beta,\varepsilon}^{\delta}(-b^{*},e)\big{)}$
$\qquad\quad\mathord{}+\psi_{\beta,\delta}\Big{(}-b^{*}+\psi_{\alpha,\varepsilon}^{\beta}(a,e),\psi_{\alpha,\varepsilon}^{\delta}(a,e)+\psi_{\gamma,\varepsilon}(-c,e)-d^{*}\Big{)}$
$\qquad\quad\mathord{}+\psi_{\alpha,\delta}^{\gamma}\big{(}a,\psi_{\alpha,\varepsilon}^{\delta}(a,e)+\psi_{\gamma,\varepsilon}(-c,e)-d^{*}\big{)}+\psi_{\alpha,\varepsilon}^{\gamma}(a,e)-c,$
$D$ $\coloneqq-\bigg{(}-b^{*}+\psi_{\alpha,\varepsilon}^{\beta}(a,e)+\psi_{\alpha,\delta}^{\beta}\big{(}a,\psi_{\alpha,\varepsilon}^{\delta}(a,e)+\psi_{\gamma,\varepsilon}(-c,e)-d^{*}\big{)}$
$\qquad\quad\mathord{}+\psi_{\alpha,\gamma}\Big{(}a,\psi_{\alpha,\delta}^{\gamma}\big{(}a,\psi_{\alpha,\varepsilon}^{\delta}(a,e)+\psi_{\gamma,\varepsilon}(-c,e)-d^{*}\big{)}+\psi_{\alpha,\varepsilon}^{\gamma}(a,e)-c\big{)}\bigg{)}.$
The following computation shows how to arrive at the formulas in 7.11.
Lemma 7.12 ([Wie23, 6.3.4]).
The maps $\varphi_{12}$, $\varphi_{21}$, $\varphi_{13}$ and $\varphi_{31}$ are blueprint rewriting rules.
Proof.
Let $a,b,c,A,B,C\in\mathcal{R}$ such that
$\varphi_{12}\big{(}\theta_{\rho_{1}}(a),\theta_{\rho_{2}}(b),\theta_{\rho_{1}}(c)\big{)}=\big{(}\theta_{\rho_{2}}(A),\theta_{\rho_{1}}(B),\theta_{\rho_{2}}(C)\big{)}$
and put $f\coloneqq(\rho_{1},\rho_{2},\rho_{1})$. Then, by 7.4, Figure 5 and (6.20), we have
$\gamma_{f}(\theta_{\rho_{1}}(a),\theta_{\rho_{2}}(b),\theta_{\rho_{1}}(c))$ $=w_{\rho_{1}}w_{\rho_{2}}w_{\rho_{1}}\theta_{\rho_{1}}(a)^{w_{\rho_{2}}w_{\rho_{1}}}\theta_{\rho_{2}}(b)^{w_{\rho_{1}}}\theta_{\rho_{1}}(c)$
$=w_{\rho_{1}}w_{\rho_{2}}w_{\rho_{1}}\theta_{\rho_{1}+\rho_{2}}(a)^{w_{\rho_{1}}}\theta_{\rho_{1}+\rho_{2}}(-b)\theta_{\rho_{1}}(c)$
$=w_{\rho_{1}}w_{\rho_{2}}w_{\rho_{1}}\theta_{\rho_{2}}(a)\theta_{\rho_{1}+\rho_{2}}(-b)\theta_{\rho_{1}}(c).$
Similarly, with $g\coloneqq(\rho_{2},\rho_{1},\rho_{2})$, we compute
$\gamma_{g}(\theta_{\rho_{2}}(A),\theta_{\rho_{1}}(B),\theta_{\rho_{2}}(C))$ $=w_{\rho_{2}}w_{\rho_{1}}w_{\rho_{2}}\theta_{\rho_{2}}(A)^{w_{\rho_{1}}w_{\rho_{2}}}\theta_{\rho_{1}}(B)^{w_{\rho_{2}}}\theta_{\rho_{2}}(C)$
$=w_{\rho_{2}}w_{\rho_{1}}w_{\rho_{2}}\theta_{\rho_{1}+\rho_{2}}(-A)^{w_{\rho_{2}}}\theta_{\rho_{1}+\rho_{2}}(B)\theta_{\rho_{2}}(C)$
$=w_{\rho_{2}}w_{\rho_{1}}w_{\rho_{2}}\theta_{\rho_{1}}(A)\theta_{\rho_{1}+\rho_{2}}(B)\theta_{\rho_{2}}(C).$
Here $w_{\rho_{1}}w_{\rho_{2}}w_{\rho_{1}}=w_{\rho_{2}}w_{\rho_{1}}w_{\rho_{2}}$ by 3.15 and
$\theta_{\rho_{2}}(a)\theta_{\rho_{1}+\rho_{2}}(-b)\theta_{\rho_{1}}(c)$ $=\theta_{\rho_{1}+\rho_{2}}(-b)\theta_{\rho_{1}}(c)\theta_{\rho_{2}}(a)\left[\theta_{\rho_{1}}(c),\theta_{\rho_{2}}(a)\right]^{-1}$
$=\theta_{\rho_{1}+\rho_{2}}(-b)\theta_{\rho_{1}}(c)\theta_{\rho_{2}}(a)\theta_{\rho_{1}+\rho_{2}}(-ca)$
$=\theta_{\rho_{1}}(c)\theta_{\rho_{1}+\rho_{2}}(-b-ca)\theta_{\rho_{2}}(a).$
Since $A=c$, $B=-b-ca$ and $C=a$ by 7.11, we infer that $\varphi_{12}$ is a blueprint rewriting rule. By a similar calculation, the same holds for $\varphi_{13}$. Lastly, note that $\varphi_{21}$ and $\varphi_{31}$ are the inverses of $\varphi_{12}$ and $\varphi_{13}$, respectively. Hence they are blueprint rewriting rules by 7.5. ∎
Lemma 7.13.
The maps $\varphi_{23}$ and $\varphi_{32}$ are blueprint rewriting rules.
Proof.
The proof proceeds just as the one of 7.12. We omit the details. ∎
8 The blueprint computation for $H_{3}$-graded groups
Notation 8.1.
The notation introduced in 7.1 and 7.10 continues to hold. That is, we fix an $H_{3}$-graded group $(G,(U_{\alpha})_{\alpha\in H_{3}})$ and a choice of Weyl elements $(w_{\delta})_{\delta\in\Delta}$ for a root base $\Delta=(\rho_{1},\rho_{2},\rho_{3})$ of $H_{3}$ in standard order. We define $\mathcal{S}$ and $(\theta_{\alpha})_{\alpha\in H_{3}}$ as in 6.19 and maps $\psi_{\alpha,\gamma}^{\beta}$ as in 7.9, and we denote by $(\alpha,\beta,\gamma,\delta,\epsilon)$ the $H_{2}$-quintuple with $\alpha=\rho_{2}$ and $\epsilon=\rho_{3}$.
Using the blueprint rewriting rules from 7.11 and the homotopy cycle from Figure 6, we can perform the blueprint computation for $(G,(U_{\alpha})_{\alpha\in H_{3}})$, following the algorithm in 7.8. This yields 15 blueprint identities involving indeterminates $y_{1},\ldots,y_{15}\in\mathcal{S}$. Using these identities, we will show in this section that $\mathcal{S}$ is an associative commutative ring (see 8.3, 8.26) which decomposes as a direct sum $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ of two isomorphic ideals $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ (see 8.21, 8.24), and we will also compute explicit formulas for the commutation maps $\psi_{\xi,\zeta}^{\rho}$ from 6.19.
Remark 8.2.
Unfortunately, the majority of the 15 blueprint identities are rather long and complicated: They fill a total of 20 A4 pages. Consequently, they are not accessible for manipulation by hand, so we have decided not to include them in this paper. However, since these identities hold for all possible choices of variables $y_{1},\ldots,y_{15}\in\mathcal{S}$, we can infer more succinct evaluated blueprint identities by replacing most of the indeterminates with $0_{\mathcal{S}}$ (or, in some cases, $1_{\mathcal{S}}$). The evaluated blueprint identities can be simplified by hand, which yields identities labelled (1), …, (35). We will not present how to deduce (1), …, (35) from the evaluated blueprint identities because this is a straightforward procedure, but we will give some remarks in the proofs of 8.7, 8.12 and 8.20. The interested reader can find the original blueprint identities, the evaluated blueprint identities and the code that produced them in [BW24].
Before we begin with the derivation of the identities (1), …, (35), we make some preliminary observations.
Proposition 8.3.
The multiplication on $\mathcal{S}$ is commutative.
Proof.
This follows from the second blueprint equation, which reads $-y_{14}-y_{5}y_{15}=-y_{14}-y_{15}y_{5}$. ∎
Remark 8.4.
Let $\zeta,\xi\in H_{3}$ be non-proportional and let $\rho\in\mathord{]\zeta,\xi[}$. Then
$\psi_{\zeta,\xi}^{\rho}(x,y)=-\psi_{\xi,\zeta}^{\rho}(y,x)$
for all $x,y\in\mathcal{S}$. This holds by 3.13.
Remark 8.5.
For all $\zeta,\xi\in H_{3}$ such that $\mathord{]\zeta,\xi[}$ contains exactly one element, the map $\psi_{\zeta,\xi}$ is bi-additive by 3.8.
Remark 8.6.
Let $(\bar{\alpha},\bar{\beta},\bar{\gamma},\bar{\delta},\bar{\epsilon})$ be an $H_{2}$-quintuple and let $x,x^{\prime},y,y^{\prime}\in\mathcal{S}$. Then the following hold by 5.5 and 5.6:
  1. (i)
    $\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\beta}}(x,y+y^{\prime})=\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\beta}}(x,y)+\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\beta}}(x,y^{\prime})$.
  2. (ii)
    $\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\beta}}(x+x^{\prime},y)=\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\beta}}(x,y)+\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\beta}}(x^{\prime},y)+\psi_{\bar{\gamma},\bar{\alpha}}\big{(}\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\gamma}}(x,y),y^{\prime}\big{)}$.
  3. (iii)
    $\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\gamma}}(x,y+y^{\prime})=\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\gamma}}(x,y)+\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\gamma}}(x,y^{\prime})+\psi_{\bar{\beta},\bar{\delta}}\big{(}\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\beta}}(x,y),y^{\prime}\big{)}$.
  4. (iv)
    $\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\gamma}}(x+x^{\prime},y)=\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\gamma}}(x,y)+\psi_{\bar{\alpha},\bar{\delta}}^{\bar{\gamma}}(x^{\prime},y)$.
  5. (v)
    $\psi_{\bar{\alpha},\bar{\epsilon}}^{\bar{\beta}}(x,y+y^{\prime})=\psi_{\bar{\alpha},\bar{\epsilon}}^{\bar{\beta}}(x,y)+\psi_{\bar{\alpha},\bar{\epsilon}}^{\bar{\beta}}(x,y^{\prime})$.
Proposition 8.7.
All identities in the column “Result” of Figure 7 hold for all $y_{1},\ldots,y_{15}\in\mathcal{S}$.
Proof.
This is a straightforward computation, as explained in 8.2. For example, in order to derive (4), we put $y_{12}\coloneqq 1_{\mathcal{S}}$ and $y_{i}\coloneqq 0_{\mathcal{S}}$ for all $i\in\{1,\ldots,15\}\setminus\{4,5,12\}$ in blueprint identity 9. This yields the evaluated blueprint identity
$\psi_{\gamma,\epsilon}\big{(}-y_{4}^{*}(-1_{\mathcal{S}}^{*}),y_{5}\big{)}^{*}=-\psi_{\gamma,\epsilon}(-y_{5},y_{4})^{*}.$
Using (2), (3) and 8.5, this identity is seen to be equivalent to (4). The identities have to be derived in the given order, so that each identity may be used in the proof of all subsequent identities. ∎
No. Blueprint identity Non-zero variables Result
(1) $12$ $y_{4}$ $(-y_{4})^{*}=-y_{4}^{*}$
(2) $5$ $y_{2},y_{15}$ $y_{15}y_{2}^{*}=(y_{15}y_{2})^{*}$
(3) $9$ $y_{5},y_{8}$ $\psi_{\varepsilon,\gamma}(y_{5},y_{8})=-\psi_{\alpha,\gamma}(y_{5},y_{8})$
(4) $9$ $y_{4},y_{5}$, $y_{12}=1$ $\psi_{\alpha,\gamma}(y_{4},y_{5})=\psi_{\alpha,\gamma}(y_{5},y_{4})$
(5) $14$ $y_{1},y_{3}$ $\psi_{\alpha,\gamma}(y_{1},y_{3})^{*}=\psi_{\alpha,\gamma}(y_{1},y_{3})$
(6) $5$ $y_{1},y_{12}$ $\psi_{\alpha,\gamma}(y_{1},y_{12}^{*})=-\psi_{\alpha,\gamma}(y_{1},y_{12})$
(7) $7$ $y_{4}=1,y_{10},y_{12}$ $\psi_{\beta,\delta}(y_{10},y_{12})=\psi_{\varepsilon,\gamma}(y_{10},y_{12})$
(8) $7$ $y_{4},y_{10},y_{12}$ $\psi_{\alpha,\gamma}(y_{4}y_{10},y_{12})=\psi_{\alpha,\gamma}(y_{4}y_{12},y_{10})$
(9) $13$ $y_{1},y_{4}$ $\psi_{\varepsilon,\beta}^{\gamma}(y_{1},y_{4})=-\psi_{\alpha,\delta}^{\gamma}(y_{1},y_{4}^{*})$
(10) $3$ $y_{7},y_{15}$ $\psi_{\beta,\varepsilon}^{\delta}(y_{7},y_{15})=-\psi_{\delta,\alpha}^{\beta}(y_{7},y_{15})$
(11) $3$ $y_{10},y_{14}$ $\psi_{\beta,\varepsilon}^{\delta}(y_{10}^{*},y_{14})=\psi_{\beta,\varepsilon}^{\delta}(y_{10},-y_{14})$
(12) $12$ $y_{2},y_{5}$ $\psi_{\beta,\varepsilon}^{\delta}(y_{2},y_{5})^{*}=-\psi_{\beta,\varepsilon}^{\delta}(y_{2},-y_{5})$
(13) $7$ $y_{7},y_{10}$ $\psi_{\beta,\varepsilon}^{\delta}(y_{7},y_{10}^{*})=\psi_{\beta,\varepsilon}^{\delta}(y_{7},y_{10})$
(14) $11$ $y_{1},y_{3},y_{8}$ $\psi_{\alpha,\gamma}(\psi_{\alpha,\gamma}(y_{1},y_{3}),y_{8})=0$
(15) $11$ $y_{1},y_{4},y_{9}$ $\psi_{\alpha,\gamma}(\psi_{\alpha,\delta}^{\gamma}(y_{1},y_{4}),y_{9})=0$
(16) $6$ $y_{3},y_{5},y_{13}$ $\psi_{\alpha,\delta}^{\gamma}(\psi_{\alpha,\gamma}(y_{3},y_{5}),y_{13})=0$
(17) $7$ $y_{1},y_{4},y_{12}$ $\psi_{\alpha,\delta}^{\gamma}(\psi_{\alpha,\delta}^{\gamma}(y_{1},y_{4}),y_{12})=0$
(18) $8$ $y_{4},y_{6},y_{15}$ $\psi_{\alpha,\gamma}(y_{6},\psi_{\alpha,\varepsilon}^{\gamma}(y_{15},y_{4}))=y_{4}\psi_{\beta,\varepsilon}^{\delta}(y_{6},y_{15})$
(19) $6$ $y_{7},y_{10},y_{14}$ $\psi_{\alpha,\gamma}(\psi_{\beta,\varepsilon}^{\delta}(y_{10},y_{14}),y_{7})=0$
(20) $9$ $y_{2},y_{5},y_{14}=1$ $\psi_{\beta,\varepsilon}^{\delta}(y_{5},y_{2})=-\psi_{\varepsilon,\beta}^{\gamma}(y_{5},-y_{2})$
(21) $3$ $y_{1},y_{14}$ $\psi_{\varepsilon,\alpha}^{\beta}(y_{1},y_{14})=\psi_{\alpha,\varepsilon}^{\delta}(y_{1},-y_{14})$
Figure 7: Evaluation of the blueprint identities, part 1. See 8.7.
Remark 8.8.
Let $x,y\in\mathcal{S}$. Since $\mathcal{S}$ is commutative, it follows from (2) that
$y^{*}$ $=(y1)^{*}=y1^{*}=1^{*}y\quad\text{and}\quad xy=((xy)^{*})^{*}=(xy^{*})^{*}=x^{*}y^{*}.$
Notation 8.9.
From now on, we put $f\coloneqq\psi_{\alpha,\gamma}$, $g\coloneqq\psi_{\alpha,\delta}^{\beta}$, $h_{1}\coloneqq\psi_{\alpha,\varepsilon}^{\beta}$ and $h_{2}\coloneqq\psi_{\alpha,\varepsilon}^{\gamma}$.
The following lemma summarises the identities in Figure 7. In particular, we see that all commutation maps $\psi_{\xi,\zeta}^{\rho}$ for distinct $\xi,\zeta\in\{\alpha,\beta,\gamma,\delta,\varepsilon\}$ and $\rho\in\mathord{]\xi,\zeta[}$ can be expressed in terms of the four maps $f$, $g$, $h_{1}$ and $h_{2}$.
Lemma 8.10.
The following hold (for all $x,y,z\in\mathcal{S}$), where $s$ denotes the map $(x,y)\mapsto(y,x)$:
  1. (i)
    $(xy)^{*}=xy^{*}$.
  2. (ii)
    $f=\psi_{\alpha,\gamma}=\psi_{\gamma,\epsilon}=-\psi_{\beta,\delta}=-\psi_{\gamma,\alpha}=-\psi_{\epsilon,\gamma}=\psi_{\delta,\beta}$.
  3. (iii)
    $f(x,y)=f(y,x)=f(x,y)^{*}$ and $f(x^{*},y)=f(x,y^{*})=-f(x,y)$.
  4. (iv)
    $f(x,y)$ depends only on the image of $xy$ in the largest associative quotient of $\mathcal{S}$.
  5. (v)
    $g=\psi_{\alpha,\delta}^{\beta}=-\psi_{\varepsilon,\beta}^{\delta}=-\psi_{\beta,\varepsilon}^{\gamma}=\psi_{\delta,\alpha}^{\gamma}=-\psi_{\delta,\alpha}^{\beta}\circ s=\psi_{\beta,\varepsilon}^{\delta}\circ s=\psi_{\varepsilon,\beta}^{\gamma}\circ s=-\psi_{\alpha,\delta}^{\gamma}\circ s$.
  6. (vi)
    $f$ and $g$ are biadditive and $h_{1}$ is additive in the second component.
  7. (vii)
    $g(x^{*},y)=g(x,y)=g(x,y)^{*}$ and $g(x,y^{*})=-g(x,y)$.
  8. (viii)
    $h_{1}(x,y)=\psi_{\alpha,\varepsilon}^{\beta}(x,y)=-\psi_{\alpha,\varepsilon}^{\delta}(y,-x)=-\psi_{\varepsilon,\alpha}^{\beta}(y,x)=\psi_{\varepsilon,\alpha}^{\delta}(-x,y)$.
  9. (ix)
    $f\big{(}x,h_{2}(y,z)\big{)}=f\big{(}h_{2}(y,z),x\big{)}=zg(y,x)$.
  10. (x)
    $f\big{(}f(x,y),z\big{)}=f\big{(}g(x,y),z\big{)}=g\big{(}x,f(y,z)\big{)}=g\big{(}x,g(y,z)\big{)}=0$.
Proof.
Assertion i is (2). Assertion iii follows from (4), (5) and (6) and iv follows from (4) and (8). Assertion ii follows from (3) and (7) for the first three equalities and from (4) and 8.4 for the remaining equalities. By (10), (20) and (9), we have
$g(x,y)=\psi_{\alpha,\delta}^{\beta}(x,y)=-\psi_{\epsilon,\beta}^{\delta}(x,y)=\psi_{\beta,\epsilon}^{\gamma}(-x,y)=-\psi_{\delta,\alpha}^{\gamma}(-x^{*},y).$ (8.11)
In particular, the images of the maps $\psi_{\alpha,\delta}^{\beta}$, $\psi_{\epsilon,\beta}^{\delta}$, $\psi_{\beta,\epsilon}^{\gamma}$, $\psi_{\delta,\alpha}^{\gamma}$ are identical, which allows us to apply (15) and (19) in a more general context. Assertion vi follows from 8.5, 8.6, (15), (19) and the previous observation. Assertion vii follows from (11), (13),  (8.11) and vi. Assertion v follows from (8.11), vi and vii. Finally, assertion viii is (21), assertion ix is (18) and assertion x is the culmination of (14) to (17). ∎
Proposition 8.12.
All identities in the column “Result” of Figure 8 hold for all $y_{1},\ldots,y_{15}\in\mathcal{S}$.
Proof.
We proceed as in 8.7. For most identities, the first step is to apply 8.10 x in order to remove zero summands. For (25), we use (24). For (30), we use (22) and (23). For (28), we use (24), (23) and (21). For (31), we use (21) and the substitutions $y_{5}\mapsto-y_{5}$ and $y_{15}\mapsto-y_{15}$. For (32), we use that the images of $f$ and $g$ are fixed by *. ∎
No. Blueprint identity Non-zero variables Result
(22) $3$ $y_{4},y_{15}$ $h_{1}(-y_{15},y_{4})=h_{1}(y_{15},y_{4}^{*})$
(23) $12$ $y_{1},y_{5}$ $h_{1}(-y_{5},y_{1})=h_{1}(y_{5},y_{1})^{*}$
(24) $11$ $y_{1},y_{5},y_{8}$ $f\big{(}h_{1}(y_{1},y_{5}),y_{8}\big{)}=f(y_{5},y_{8})y_{1}$
(25) $4$ $y_{4},y_{10},y_{15}$ $f(y_{4},y_{10})y_{15}=-g(y_{15},y_{4}y_{10})$
(26) $5$ $y_{1},y_{4},y_{15}$ $g\big{(}y_{1},h_{1}(y_{15},y_{4})\big{)}=g(y_{1},y_{4})y_{15}$
(27) $3$ $y_{5},y_{10},y_{15}$ $g(y_{5}y_{15},y_{10})=g(y_{15},y_{10})y_{5}$
(28) $13$ $y_{1},y_{5}$ $h_{2}(y_{1},y_{5})=h_{2}(y_{5},y_{1})$
[…]
(30) $3$ $y_{3},y_{5},y_{15}$ $f\big{(}h_{1}(y_{15},y_{3}),y_{15}y_{5}\big{)}=h_{1}\big{(}y_{15},f(y_{3},y_{5})\big{)}$
(31) $3$ $y_{1},y_{5},y_{15}$ $h_{1}\big{(}y_{15},h_{1}(y_{5},y_{1})\big{)}=h_{1}(y_{15}y_{5},y_{1})$
(32) $4$ $y_{1},y_{7},y_{15}$ $g\big{(}g(y_{15},y_{7}),y_{1}\big{)}=-g(y_{15},y_{1}y_{7})$
Figure 8: Evaluation of the blueprint identities, part 2. See 8.12.
Before we can prove the remaining identities (33), (34), (35), we have to collect some consequences of (22), …, (32).
Proposition 8.13.
The map $-g(1,\mathord{\>\cdot\>})\colon\mathcal{S}\rightarrow\mathcal{S}$ preserves the addition and multiplication of $\mathcal{S}$. In other words, $g(1,x+y)=g(1,x)+g(1,y)$ and $g(1,xy)=-g(1,x)g(1,y)$ for all $x,y\in\mathcal{S}$.
Proof.
The map is additive by 8.10 vi and we have
$g(1,xy)=-g\big{(}g(1,y),x\big{)}=-g\big{(}g(1,y)1,x\big{)}=-g(1,x)g(1,y)$
for all $x,y\in\mathcal{S}$ by (32) and (27). ∎
Lemma 8.14.
The following hold:
  1. (i)
    $[x,y,g(1,1)]=[x,g(1,1),y]=[g(1,1),x,y]=0$ for all $x,y\in\mathcal{S}$ where $[x,y,z]\coloneqq(xy)z-x(yz)$.
  2. (ii)
    $g(x,y)g(1,1)=-g(x,y)=g(1,1)g(x,y)$ for all $x,y\in\mathcal{S}$.
Proof.
Let $x,y\in\mathcal{S}$. By (27) and the commutativity of $\mathcal{S}$, we have
$x\big{(}g(1,1)y\big{)}$ $=xg(y,1)=g(xy,1)=g(x,1)y=\big{(}xg(1,1)\big{)}y,$
so that $[x,g(1,1),y]=0$. It follows in a similar way that $[x,y,g(1,1)]=[g(1,1),x,y]=0$ as well. Further, by two applications of (27) and one application of (32),
$g(x,y)g(1,1)$ $=g\big{(}g(1,1)x,y\big{)}=g\big{(}g(x,1),y\big{)}=-g(x,y).$
This finishes the proof. ∎
Proposition 8.15.
The map $-g(\mathord{\>\cdot\>},1)\colon\mathcal{S}\rightarrow\mathcal{S}$ preserves the addition and multiplication of $\mathcal{S}$. In other words, $g(x+y,1)=g(x,1)+g(y,1)$ and $g(xy,1)=-g(x,1)g(y,1)$ for all $x,y\in\mathcal{S}$.
Proof.
Again, the map is additive by 8.10 vi. Further,
$g(xy,1)$ $=g(y,1)x=-\big{(}g(y,1)g(1,1)\big{)}x=-g(y,1)\big{(}g(1,1)x\big{)}=-g(y,1)g(x,1)$
for all $x,y\in\mathcal{S}$ by (27) and 8.14. ∎
Notation 8.16.
We put $\mathcal{S}_{1}\coloneqq\ker(g(\mathord{\>\cdot\>},1))=\{x\in\mathcal{S}\mid g(x,1)=0\}$ and $\mathcal{S}_{2}\coloneqq\ker(g(1,\mathord{\>\cdot\>}))=\{x\in\mathcal{S}\mid g(1,x)=0\}$. Further, we define maps
$\operatorname{pr}_{1}\colon\mathcal{S}\rightarrow\mathcal{S},x\mapsto-h_{1}\big{(}-f(x,1),1\big{)}\quad\text{and}\quad\operatorname{pr}_{2}\colon\mathcal{S}\rightarrow\mathcal{S},x\mapsto-g(x,1).$
Remark 8.17.
By 8.13 and 8.15, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are ideals of $\mathcal{S}$. Further, for all $x\in\mathcal{S}$ and $y\in\mathcal{S}_{2}$, we have $g(x,y)=g(1,y)x=0x=0$ by (27).
Remark 8.18.
It follows from (26) that $h_{1}(y_{15},y_{4})\in\mathcal{S}_{2}$ for all $y_{4}\in\mathcal{S}_{2}$ and $y_{15}\in\mathcal{S}$.
Remark 8.19.
By (25), we have $f(1,x)=-g(1,x)$ for all $x\in\mathcal{S}$. In particular, $\mathcal{S}_{2}=\ker(f(1,\mathord{\>\cdot\>}))=\ker(f(\mathord{\>\cdot\>},1))$.
Proposition 8.20.
All identities in the column “Result” of Figure 9 hold for all $y_{1},\ldots,y_{15}\in\mathcal{S}$.
Proof.
We proceed as in 8.7 and 8.12. For (33), we use in particular 8.10 ix, 8.17 and 8.18. For (34), we use (24), (32) and 8.19. For (35), we use 8.10 iv and 8.10 x. ∎
No. Blueprint identity Non-zero variables Result Remarks
(33) $6$ $y_{5},y_{10},y_{15}$ $g\big{(}h_{1}(y_{5},y_{10}),y_{15}\big{)}=0$ if $y_{5}\in\mathcal{S}_{2}$
(34) $11$ $y_{1},y_{5},y_{12}=1$ $g\big{(}h_{2}(y_{1},y_{5}),1\big{)}$ $=g\big{(}\big{(}f(y_{1},y_{5})y_{5}\big{)}y_{1},1\big{)}$
(35) $6$ $y_{1},y_{4},y_{14}$ $g\big{(}h_{1}(y_{14},y_{1}),y_{4}\big{)}$ $=h_{1}(-y_{14},g(y_{1},y_{4}))$
Figure 9: Evaluation of the blueprint identities, part 3. See 8.20.
It remains to establish that $\mathcal{S}$ decomposes as $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$, that it is associative and to compute explicit formulas for the commutation maps and for the involution *.
Proposition 8.21.
We have $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$, that is, the ring $\mathcal{S}$ is the direct sum of its ideals $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. Further, the projection maps on $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are given by $\operatorname{pr}_{1}$ and $\operatorname{pr}_{2}$, respectively.
Proof.
Let $x\in\mathcal{S}$. By 8.10 x, $f(x,1)$ lies in $\mathcal{S}_{2}$, so (33) yields that
$g\big{(}\operatorname{pr}_{1}(x),1\big{)}=-g\big{(}h_{1}(-f(x,1),1),1\big{)}=0.$
Hence $\operatorname{pr}_{1}(x)$ lies in $\mathcal{S}_{1}$. Further, $\operatorname{pr}_{2}(x)=-g(x,1)$ lies in $\mathcal{S}_{2}$ by 8.10 x.
Identity (29) with $y_{12}=y_{5}=1$ and $y_{1}=x$ says that
$x=-h_{1}\big{(}-f(x,1),1\big{)}+g\big{(}1,h_{2}(x,1)\big{)}=\operatorname{pr}_{1}(x)+g\big{(}1,h_{2}(x,1)\big{)}.$
Since $g\big{(}1,h_{2}(x,1)\big{)}=-f\big{(}1,h_{2}(x,1)\big{)}=-g(x,1)=\operatorname{pr}_{2}(x)$ by 8.19 and 8.10 ix, we conclude that $x=\operatorname{pr}_{1}(x)+\operatorname{pr}_{2}(x)$. In particular, $\mathcal{S}=\mathcal{S}_{1}+\mathcal{S}_{2}$.
Finally, let $x\in\mathcal{S}_{1}\cap\mathcal{S}_{2}$. Then $\operatorname{pr}_{2}(x)=-g(x,1)=0$ by the definition of $\mathcal{S}_{1}$. Further,
$\operatorname{pr}_{1}(x)$ $=-h_{1}\big{(}-f(x,1),1\big{)}=-h_{1}\big{(}g(1,x),1\big{)}=-h_{1}(0,1)=0$
by 8.19. Thus $x=\operatorname{pr}_{1}(x)+\operatorname{pr}_{2}(x)=0$. This finishes the proof. ∎
Remark 8.22.
It follows from 8.21 that $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ are nonassociative rings with identity elements $\operatorname{pr}_{1}(1_{\mathcal{S}})=-h_{1}(-f(1,1),1)$ and $\operatorname{pr}_{2}(1_{\mathcal{S}})=-g(1,1)=f(1,1)$, respectively, and that $\operatorname{pr}_{1}\colon\mathcal{S}\rightarrow\mathcal{S}_{1}$, $\operatorname{pr}_{2}\colon\mathcal{S}\rightarrow\mathcal{S}_{2}$ are homomorphisms.
Notation 8.23.
We define $\varphi\colon\mathcal{S}_{1}\rightarrow\mathcal{S}_{2},x\mapsto f(1,x)=-g(1,x)$.
Proposition 8.24.
The map $\varphi$ is a well-defined isomorphism of nonassociative rings with inverse $\varphi^{-1}\colon\mathcal{S}_{2}\rightarrow\mathcal{S}_{1},x\mapsto h_{1}(x,1)$.
Proof.
The image of $\varphi$ lies in $\mathcal{S}_{2}$ by 8.10 x and it preserves the addition and multiplication by 8.13 and 8.22. Further, it follows from (24) that
$\varphi(1_{\mathcal{S}_{1}})$ $=f\big{(}1,-h_{1}\big{(}-f(1,1),1\big{)}\big{)}=f(1,1)f(1,1)=1_{\mathcal{S}_{2}}1_{\mathcal{S}_{2}}=1_{\mathcal{S}_{2}},$
so $\varphi$ is a homomorphism. The kernel of $\varphi$ is $\mathcal{S}_{1}\cap\mathcal{S}_{2}$, so $\varphi$ is injective. Further,
$\varphi\big{(}h_{1}(x,1)\big{)}$ $=f\big{(}h_{1}(x,1),1\big{)}=f(1,1)x=1_{\mathcal{S}_{2}}x=x$
for all $x\in\mathcal{S}_{2}$ by (24). Hence $\varphi$ is an isomorphism with the desired inverse. ∎
Remark 8.25.
It is not true that the map $\mathcal{S}_{2}\rightarrow\mathcal{S}_{1},x\mapsto f(1,x)$ is an isomorphism. In fact, it is the zero map by 8.19.
Proposition 8.26.
The multiplication on $\mathcal{S}$ is associative.
Proof.
For all $x,y,z\in\mathcal{S}_{1}$, we have
$\varphi\big{(}x(yz)\big{)}$ $=f\big{(}1,x(yz)\big{)}=f\big{(}1,(xy)z\big{)}=\varphi\big{(}(xy)z\big{)}$
by 8.10 iv. Since $\varphi\colon\mathcal{S}_{1}\rightarrow\mathcal{S}_{2}$ is an isomorphism by 8.24, it follows that $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are associative. Hence $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is associative as well. ∎
Proposition 8.27.
For all $x\in\mathcal{S}$, we have $x^{*}=-\operatorname{pr}_{1}(x)+\operatorname{pr}_{2}(x)$.
Proof.
Let $x\in\mathcal{S}_{1}$. By 8.10 iii, we have $\varphi(x^{*})=f(1,x^{*})=-f(1,x)=\varphi(-x)$ and $\varphi(x)^{*}=f(1,x)^{*}=f(1,x)=\varphi(x)$. Since $\varphi\colon\mathcal{S}_{1}\rightarrow\mathcal{S}_{2}$ is an isomorphism and * is additive, the assertion follows. ∎
Notation 8.28.
For all $x,y\in\mathcal{S}_{1}$, we put $[x,y]\coloneqq x+\varphi(y)\in\mathcal{S}$. This defines an isomorphism between the (abstract) direct sum $\mathcal{S}_{1}\oplus\mathcal{S}_{1}$ and $\mathcal{S}$ with the property that $\varphi([x,0])=[0,x]$ and $\varphi^{-1}([0,x])=[x,0]$ for all $x\in\mathcal{S}_{1}$.
Proposition 8.29.
Let $x_{1},x_{2},y_{1},y_{2}\in\mathcal{S}_{1}$. Then the following hold:
  1. (i)
    $f([x_{1},x_{2}],[y_{1},y_{2}])=[0,x_{1}y_{1}]$.
  2. (ii)
    $g([x_{1},x_{2}],[y_{1},y_{2}])=[0,-y_{1}x_{2}]$.
  3. (iii)
    $h_{1}([x_{1},x_{2}],[y_{1},y_{2}])=[x_{2}y_{1},x_{1}x_{2}y_{2}]$.
  4. (iv)
    $h_{2}([x_{1},x_{2}],[y_{1},y_{2}])=[-x_{2}y_{2},x_{1}x_{2}y_{1}y_{2}]$.
Proof.
Put $x\coloneqq[x_{1},x_{2}],y\coloneqq[y_{1},y_{2}]\in\mathcal{S}$. By 8.10 x and 8.10 iv, we have
$f([x_{1},x_{2}],[y_{1},y_{2}])$ $=f\big{(}x_{1}+f(x_{2},1),y_{1}+f(y_{2},1)\big{)}=f(x_{1},y_{1})=f(1,x_{1}y_{1})$
$=\varphi(x_{1}y_{1})=[0,x_{1}y_{1}],$
which proves i. It follows from (32) and (25) that
$g(x,y)$ $=-g\big{(}g(x,1),y\cdot 1\big{)}=f(y,1)g(x,1).$
Here $g(x,1)=-\operatorname{pr}_{2}(x)=[0,-x_{2}]$ and, by i, $f(y,1)=[0,y_{1}]$. Hence ii holds. Further, by an application of (24),
$\operatorname{pr}_{1}\big{(}h_{1}(x,y)\big{)}$ $=-h_{1}\big{(}-f(h_{1}(x,y),1),1)=-h_{1}\big{(}-f(y,1)x,1\big{)}.$
Here $-f(y,1)x=-[0,y_{1}][x_{1},x_{2}]=[0,-y_{1}x_{2}]$. Since $h_{1}(\mathord{\>\cdot\>},1)=\varphi^{-1}$ on $\mathcal{S}_{2}$ by 8.24, we infer that
$\operatorname{pr}_{1}\big{(}h_{1}(x,y)\big{)}$ $=-\varphi^{-1}([0,-y_{1}x_{2}])=[y_{1}x_{2},0]=y_{1}x_{2}.$
Similarly, by (35) and (30),
$\operatorname{pr}_{2}\big{(}h_{1}(x,y)\big{)}$ $=-g\big{(}h_{1}(x,y),1\big{)}=-h_{1}\big{(}-x,g(y,1)\big{)}=h_{1}\big{(}-x,[0,y_{2}]\big{)}$
$=h_{1}\big{(}-x,f(1,[y_{2},0])\big{)}=-f\big{(}h_{1}(-x,1),x[y_{2},0]\big{)}.$
Since $\operatorname{pr}_{1}(h_{1}(-x,1))=-x_{2}$ by the previous part of the proof, we infer that
$\operatorname{pr}_{2}\big{(}h_{1}(x,y)\big{)}$ $=-[0,-x_{1}x_{2}y_{2}]=\varphi(x_{1}x_{2}y_{2}).$
Assertion iii follows. Finally, by 8.10 ix,
$\operatorname{pr}_{1}\big{(}h_{2}(x,y)\big{)}$ $=-h_{1}\big{(}-f\big{(}h_{2}(x,y),1\big{)},1\big{)}=-h_{1}\big{(}-yg(x,1),1\big{)}$
$=-h_{1}([0,x_{2}y_{2}],[1,1])=[-x_{2}y_{2},0]$
and, by (34),
$\operatorname{pr}_{2}\big{(}h_{2}(x,y)\big{)}$ $=-g\big{(}h_{2}(x,y),1\big{)}=-g\big{(}f(x,y)yx,1\big{)}$
$=-g([0,x_{1}x_{2}y_{1}y_{2}][1,1])=[0,x_{1}y_{1}x_{2}y_{2}].$
This finishes the proof. ∎
Remark 8.30.
The formulas in 8.29 show that the maps $f$, $g$, $h_{1}$, $h_{2}$ have the same form as the maps $\tilde{\psi}_{\alpha,\gamma}$, $\tilde{\psi}_{\alpha,\delta}^{\beta}$, $\tilde{\psi}_{\alpha,\varepsilon}^{\beta}$, $\tilde{\psi}_{\alpha,\varepsilon}^{\gamma}$ in Figure 4. Using 8.10, we infer that for all $\xi,\zeta\in\{\alpha,\beta,\gamma,\delta,\varepsilon\}$ and all $\rho\in\mathord{]\xi,\zeta[}$, the commutation map $\psi_{\xi,\zeta}^{\rho}$ in $G$ has the same form as the map $\tilde{\psi}_{\xi,\zeta}^{\rho}$ in Figure 4. A similar assertion holds for $\xi,\zeta\in\{\rho_{1},\rho_{1}+\rho_{2},\rho_{2}\}$ and for $\xi,\zeta\in\{\rho_{1},\rho_{3}\}$. Further, we have shown in  5.21 and 8.27 that the involution * has the same form as in the $H_{3}$-grading of $\operatorname{Ch}(D_{6},\mathcal{S}_{1})$ constructed in 4.10.
We can now state the main result of this paper, which will be reformulated in the language of foldings in 8.40. Note the distinction that we make between “parametrisations” and “coordinatisations”.
Theorem 8.31.
Let $G$ be a group with an $H_{3}$-grading $(U_{\xi})_{\xi\in H_{3}}$ and fix a family $(w_{\rho})_{\rho\in\Delta}$ of Weyl elements in $G$. Denote by $\eta\colon H_{3}\times\Delta\rightarrow\{\pm 1\}^{2}$ the parity map from 6.16 and by $(\alpha,\beta,\gamma,\delta,\varepsilon)$ the $H_{2}$-quintuple corresponding to the $H_{2}$-pair $(\rho_{2},\rho_{3})=(\alpha,\varepsilon)$ in the fixed root base $\Delta$ of $H_{3}$. Then there exists a commutative ring $\mathcal{R}$ such that $G$ is coordinated by $\mathcal{R}\times\mathcal{R}$ with respect to $\eta$ and $(w_{\rho})_{\rho\in\Delta}$. By this we mean that there exists a parametrisation $(\theta_{\xi}\colon\mathcal{R}\times\mathcal{R}\rightarrow U_{\xi})_{\xi\in H_{3}}$ (in the sense of 6.9) of $(G,(U_{\xi})_{\xi\in H_{3}})$ by the standard parameter system ({±$(\{\pm 1\}^{2},\mathcal{R}\times\mathcal{R})$ for $\mathcal{R}$ from 6.10 with respect to $\eta$ and $(w_{\rho})_{\rho\in\Delta}$ such that the same commutator relations as in Figure 4 are satisfied.
Proof.
By 8.3 and 8.26, $\mathcal{R}\coloneqq\mathcal{S}_{1}$ is a commutative ring. By 8.1, 8.28 and 8.27, $(G,(U_{\alpha})_{\alpha\in H_{3}})$ is parametrised by the standard parameter system for $\mathcal{R}$ with respect to $\eta$ and $(w_{\rho})_{\rho\in\Delta}$. The desired commutator relations in the $H_{2}$-quintuple $(\alpha,\beta,\gamma,\delta,\varepsilon)$ hold by 8.30 and the commutator relation in the $A_{2}$-pair $(\rho_{1},\rho_{2})$ holds by the definition of the multiplication on $\mathcal{S}$ in 6.19. ∎
Remark 8.32.
Let everything be as in 8.31 and let $(\xi,\zeta)$ be a pair of non-proportional roots in $H_{3}$. Then there exists $w\in\operatorname{Weyl}(H_{3})$ such that $\xi^{w},\zeta^{w}$ lie in the positive part of the rank-2 subsystem spanned by two roots in $\Delta$. Thus we can conjugate one of the formulas in Figure 4 by an appropriate product of Weyl elements to obtain a commutator formula for $[\theta_{\xi}(r,s),\theta_{\zeta}(r^{\prime},s^{\prime})]$ where $r,s,r^{\prime},s^{\prime}\in\mathcal{R}$. We will see an example in 8.33.
It is a corollary of 8.31 that the $H_{3}$-grading $(U_{\alpha})_{\alpha\in H_{3}}$ of $G$ (and hence any $H_{3}$-grading of any group) is the folding of a $D_{6}$-grading in the sense of 4.10 (and 4.11). The remaining part of this paper deals with the technical details of this deduction. The first intermediate goal is 8.36.
Lemma 8.33.
For all $x,y\in\mathcal{S}$, we have
$[\theta_{-\delta}(x),\theta_{\gamma}(y)]$ $=\theta_{-\varepsilon}\big{(}h_{1}(x^{*},-y)^{*}\big{)}\theta_{\alpha}\big{(}-h_{2}(x^{*},-y)^{*}\big{)}\theta_{\beta}\big{(}-h_{1}(y,x^{*})\big{)}.$
Proof.
For all $x,y\in\mathcal{S}$, we know from 8.10 that
$[\theta_{\alpha}(x),\theta_{\varepsilon}(y)]=\theta_{\beta}\big{(}h_{1}(x,y)\big{)}\theta_{\gamma}\big{(}h_{2}(x,y)\big{)}\theta_{\delta}\big{(}-h_{1}(-y,x)\big{)}.$
Using the values of the parity map $\eta$ from Figure 5 and (6.20), we can conjugate this equation by $w_{\alpha}w_{\varepsilon}$. This yields
$[\theta_{-\delta}(x^{*}),\theta_{\gamma}(-y)]=\theta_{-\varepsilon}\big{(}h_{1}(x,y)^{*}\big{)}\theta_{\alpha}\big{(}-h_{2}(x,y)^{*}\big{)}\theta_{\beta}\big{(}-h_{1}(-y,x)\big{)}$
for all $x,y\in\mathcal{S}$. The assertion follows. ∎
Lemma 8.34.
Let $(\bar{\alpha},\bar{\beta},\bar{\gamma},\bar{\delta},\bar{\varepsilon})$ be an arbitrary $H_{2}$-quintuple. Then $\theta_{\bar{\alpha}}(\mathcal{S}_{2})$ and $\theta_{\bar{\delta}}(\mathcal{S}_{2})$ commute.
Proof.
We have $g(\mathcal{S}_{2},\mathcal{S}_{2})=\{0\}$ by 8.29 ii. Thus by 8.10 v, the assertion holds for $(\bar{\alpha},\bar{\beta},\bar{\gamma},\bar{\delta},\bar{\varepsilon})=(\alpha,\beta,\gamma,\delta,\varepsilon)$. The general assertion follows by conjugation with Weyl elements. ∎
Remark 8.35.
For all $x,y,z$ in any group, the Hall-Witt identity says that
$\big{[}[x,y],z^{x}]\big{[}[z,x],y^{z}]\big{[}[y,z],x^{y}]=1.$
Proposition 8.36.
For all $\alpha\in H_{3}$, we have $[\theta_{\alpha}(\mathcal{S}_{1}),\theta_{-\alpha}(\mathcal{S}_{2})]=\{1\}$.
Proof.
Let $x,y\in\mathcal{S}_{1}$. By 8.35, we have $ABC=1_{G}$ where
$A$ $=\big{[}[\theta_{-\delta}([0,1]),\theta_{\gamma}([0,x])],\theta_{-\alpha}([0,y])^{\theta_{-\delta}([0,1])}],$
$B$ $=\big{[}[\theta_{-\alpha}([0,y]),\theta_{-\delta}([0,1])],\theta_{\gamma}([0,x])^{\theta_{-\alpha}([0,y])}],$
$C$ $=\big{[}[\theta_{\gamma}([0,x]),\theta_{-\alpha}([0,y])],\theta_{-\delta}([0,1])^{\theta_{\gamma}([0,x])}].$
It follows from 8.34 that $B=1=C$ and that $\theta_{-\alpha}([0,y])^{\theta_{-\delta}([0,1])}=\theta_{-\alpha}([0,y])$. Further, since $h_{1}(\mathcal{S}_{2},\mathcal{S}_{2})=\{0\}$ by 8.29 iii, we infer from 8.33 that
$[\theta_{-\delta}([0,1]),\theta_{\gamma}([0,x])]$ $=\theta_{\alpha}\big{(}-h_{2}([0,1],[0,-x])^{*}\big{)}=\theta_{\alpha}([x,0]).$
Altogether,
$1_{G}$ $=ABC=A=[\theta_{\alpha}([x,0]),\theta_{-\alpha}([0,y])].$
This finishes the proof. ∎
We now define the $D_{6}$-grading of $G$ whose folding is $(U_{\alpha})_{\alpha\in H_{3}}$.
Notation 8.37.
For the rest of this section, we use the notation of 4.5 and 4.6. In particular, we write $\Delta_{D}=(\delta_{1},\ldots,\delta_{6})$ for the standard root base of $D_{6}$, $\Delta_{H}=\Delta=(\rho_{1},\rho_{2},\rho_{3})$ for the (corresponding) standard root base of $H_{3}$ and $\phi\circ\pi$ for the chain of maps $D_{6}\rightarrow GH_{3}\rightarrow H_{3}$. For all $\alpha\in H_{3}$, we define $V_{\kappa^{\text{sh}}(\alpha)}\coloneqq\theta_{\alpha}(\mathcal{S}_{1})$ and $V_{\kappa^{\text{lo}}(\alpha)}\coloneqq\theta_{\alpha}(\mathcal{S}_{2})$ where $\kappa^{\text{sh}}(\alpha),\kappa^{\text{lo}}(\alpha)\in D_{6}$ are as in 4.6. In this way, we have defined a root group $V_{\beta}$ for all $\beta\in D_{6}$.
Lemma 8.38.
Let $(i,j)\in\{(1,6),(2,4),(3,5)\}$ and put $\rho\coloneqq\phi\circ\pi(\delta_{i})=\phi\circ\pi(\delta_{j})\in\Delta\subset H_{3}$. Then for all $\xi\in D_{6}$, we have $V_{\xi}^{w_{\rho}}=V_{\xi^{\sigma(\delta_{i})\sigma(\delta_{j})}}$.
Proof.
Let $\zeta\in H_{3}$ and put $\xi\coloneqq\kappa^{\text{sh}}(\zeta)\in D_{6}$. Then
$V_{\xi}^{w_{\rho}}$ $=\theta_{\zeta}(\mathcal{S}_{1})^{w_{\rho}}=\theta_{\zeta^{\sigma(\rho)}}(\eta_{\zeta,\rho}.\mathcal{S}_{1})=\theta_{\zeta^{\sigma(\rho)}}(\mathcal{S}_{1})=V_{\kappa^{\text{sh}}(\zeta^{\sigma(\rho)})}$
$=V_{\kappa^{\text{sh}}(\zeta)^{\sigma(\delta_{i})\sigma(\delta_{j})}}=V_{\xi^{\sigma(\delta_{i})\sigma(\delta_{j})}}.$
by 4.13. The assertion for $\xi\coloneqq\kappa^{\text{lo}}(\zeta)$ can be proven similarly. ∎
Lemma 8.39.
Let $\xi_{1},\xi_{2}\in D_{6}$ and put $\zeta_{1}\coloneqq\phi\circ\pi(\xi_{1}),\zeta_{2}\coloneqq\phi\circ\pi(\xi_{2})\in H_{3}$. Assume that $\zeta_{1},\zeta_{2}\in\{\alpha,\beta,\gamma,\delta,\varepsilon\}$ or $\zeta_{1},\zeta_{2}\in\{\rho_{1},\rho_{1}+\rho_{2},\rho_{2}\}$ or $\zeta_{1},\zeta_{2}\in\{\rho_{1},\rho_{3}\}$. Then $[V_{\xi_{1}},V_{\xi_{2}}]\leq V_{\xi_{1}+\xi_{2}}$ if $\xi_{1}+\xi_{2}$ is a root (in $D_{6}$) and $[V_{\xi_{1}},V_{\xi_{2}}]=\{1\}$ otherwise.
Proof.
The $D_{6}$-root groups $V_{\xi_{1}},V_{\xi_{2}}$ are contained in the $H_{3}$-root groups $U_{\zeta_{1}}$, $U_{\zeta_{2}}$, respectively. For the latter, we have a commutator formula $[U_{\zeta_{1}},U_{\zeta_{2}}]\subseteq U_{\mathord{]\zeta_{1},\zeta_{2}[}}$. By 8.30, the commutator maps $\psi_{\zeta_{1},\zeta_{2}}^{\rho}$ in this commutator formula are of the same form as in the folding $(\operatorname{Ch}(D_{6},\mathbb{C}),(\tilde{U}_{\rho}^{D})_{\rho\in D_{6}})$. The assertion follows. ∎
Theorem 8.40.
The family $(V_{\xi})_{\xi\in D_{6}}$ is a $D_{6}$-grading of $G$, and its folding with respect to the diagram in Figure 1(b) is precisely $(U_{\zeta})_{\zeta\in H_{3}}$. In particular, every $H_{3}$-graded group is the folding of a $D_{6}$-graded group.
Proof.
We begin by verifying the $D_{6}$-commutator relations. Let $\xi_{1},\xi_{2}\in D_{6}$ be non-proportional and put $\zeta_{1}\coloneqq\pi(\xi_{1})$, $\zeta_{2}\coloneqq\pi(\xi_{2})\in GH_{3}$. If $\zeta_{1}=e\zeta_{2}$ for some $e\in\{\pm 1\}$, then $\xi_{1}=e\xi_{2}$ because $\pi\colon D_{6}\rightarrow GH_{3}$ is bijective and linear, which contradicts the assumption that $\xi_{1},\xi_{2}$ are non-proportional. If $\zeta_{1},\zeta_{2}$ are proportional but not of the same length, then $[V_{\xi_{1}},V_{\xi_{2}}]=\{1_{G}\}$ by 8.36. Hence we can assume that $\zeta_{1},\zeta_{2}$ are non-proportional.
Choose a root base $\Delta_{H}^{\prime}$ of the rank-2 subsystem of $H_{3}$ spanned by $\bar{\zeta}_{1}\coloneqq\phi(\zeta_{1})$ and $\bar{\zeta}_{2}\coloneqq\phi(\zeta_{2})$ such that $\bar{\zeta}_{1},\bar{\zeta}_{2}$ are $\Delta_{H}^{\prime}$-positive and put $r_{1}\coloneqq\sigma(\delta_{1})\sigma(\delta_{6})$, $r_{2}\coloneqq\sigma(\delta_{2})\sigma(\delta_{4})$, $r_{3}\coloneqq\sigma(\delta_{3})\sigma(\delta_{5})$. Recall from 4.5 and 4.13 that $r_{1},r_{2},r_{3}$ are the simple reflections in $\operatorname{Weyl}(H_{3})$ corresponding to $\Delta_{H}=(\rho_{1},\rho_{2},\rho_{3})$ and that they also act on $D_{6}$. The transitivity properties of the Weyl group yield $i_{1},\ldots,i_{k}\in\{1,2,3\}$ such that $\Delta_{H}$ contains $(\Delta_{H}^{\prime})^{r}$ where $r\coloneqq r_{i_{1}}\cdots r_{i_{k}}$. For all $j\in\{1,2\}$, we further put $\bar{\zeta}_{j}^{\prime}\coloneqq\bar{\zeta}_{j}^{r}$ and $\xi_{j}^{\prime}\coloneqq\xi_{j}^{r}$, so that
$\phi\pi(\xi_{j}^{\prime})=\phi(\pi(\xi_{j})^{r})=\phi(\zeta_{j}^{r})=\phi(\zeta_{j})^{r}=\bar{\zeta}_{j}^{\prime}$
by 4.13. Then $\bar{\zeta}_{1}^{\prime}$ and $\bar{\zeta}_{2}^{\prime}$ are $\Delta_{H}$-positive, so it follows from 8.39 that $V_{\xi_{1}^{\prime}}$ and $V_{\xi_{2}^{\prime}}$ satisfy the desired commutator relations. By conjugation with $(w_{\rho_{i_{1}}}\cdots w_{\rho_{i_{k}}})^{-1}$ and an application of 8.38, it follows that $V_{\xi_{1}}$ and $V_{\xi_{2}}$ satisfy the desired commutator relations as well. Hence $(V_{\xi})_{\xi\in D_{6}}$ satisfies $D_{6}$-commutator relations.
Even more, the previous paragraphs show that $(V_{\xi})_{\xi\in D_{6}}$ satisfies the same commutator formulas as the Chevalley group $\operatorname{Ch}(D_{6},\mathcal{S}_{1})$. This implies that
$w_{\kappa^{\text{sh}}(\zeta)}$ $\coloneqq\theta_{-\zeta}(-1,0)\theta_{\zeta}(1,0)\theta_{-\zeta}(-1,0)\quad\text{and}$
$w_{\kappa^{\text{lo}}(\zeta)}$ $\coloneqq\theta_{-\zeta}(0,-1)\theta_{\zeta}(0,1)\theta_{-\zeta}(0,-1)$
are Weyl elements with respect to $(V_{\xi})_{\xi\in D_{6}}$ for all $\zeta\in H_{3}$ (see [Wie23, 5.6.6] for details). Further, since $\phi\pi(\Pi_{D})=\Pi_{H}$ by 4.8, Axiom 3.4 (iv) is satisfied. Hence $(V_{\xi})_{\xi\in D_{6}}$ is indeed a $D_{6}$-grading of $G$. Finally, since $U_{\zeta}=\langle V_{\kappa^{\text{sh}}(\zeta)},V_{\kappa^{\text{lo}}(\zeta)}\rangle$ for all $\zeta\in H_{3}$, the grading $(U_{\zeta})_{\zeta\in H_{3}}$ is the folding of $(V_{\xi})_{\xi\in D_{6}}$. ∎
Remark 8.41.
The arguments and blueprint computations in this section also apply to $H_{4}$-graded groups by considering an $H_{3}$-graded subgroup. We conclude that 8.31 holds for $H_{4}$-graded groups as well, except that we have to use the parity map from 6.22 and that we have another commutator relation for the additional $A_{2}$-pair in $H_{4}$ (which is also given by the ring multiplication). Similarly, we see that every $H_{4}$-graded group is the folding of an $E_{8}$-graded group via the diagram in Figure 1(c).